### v1

# CroCoDai: A Stablecoin for Cross-Chain Commerce

Abstract
Decentralized Finance (DeFi), in which digital assets are exchanged without trusted intermediaries, has grown rapidly in value in recent years. The global DeFi ecosystem is fragmented into multiple blockchains, fueling the demand for cross-chain commerce. Existing approaches for cross-chain transactions, e.g., bridges and cross-chain deals, achieve atomicity by locking assets in escrow. However, locking up assets increases the financial risks for the participants, especially due to price fluctuations and the long latency of cross-chain transactions. Stablecoins, which are pegged to a non-volatile asset such as the US dollar, help mitigate the risk associated with price fluctuations. However, existing stablecoin designs are tied to individual blockchain platforms, and trusted parties or complex protocols are needed to exchange stablecoin tokens between blockchains.
Our goal is to design a practical stablecoin for cross-chain commerce. Realizing this goal requires addressing two challenges. The first challenge is to support a large and growing number of blockchains efficiently. The second challenge is to be resilient to price fluctuations and blockchain platform failures. We present CroCoDai to address these challenges. The key insight is to extend the existing Dai stablecoin system to a stablecoin that is backed with collateral tokens from multiple blockchains. CroCoDai leverages the system’s governance layer (which could be another blockchain) to exchange tokens between different blockchains – this enables secure cross-chain commerce and addresses the first challenge. Furthermore, the relatively low correlation between the price of assets on different blockchains makes CroCoDai more resilient to financial shocks than a single-chain stablecoin. We demonstrate this advantage through an empirical analysis that uses real token price data from a period that includes the collapse of the FTX exchange, which caused high token price volatility. We present three prototype implementations of our stablecoin system, and show that it incurs small execution overhead.
keywords:
Stablecoins, blockchains, interoperability, financial risk analysis
category:
[…]
1 Introduction
Decentralized Finance (DeFi) promises to revolutionize finance by removing the need for trusted middlemen to process transactions. In DeFi, transactions commonly involve a change of ownership of digital tokens that exist on a distributed ledger such as a blockchain – e.g., one user sending bitcoins to another [45]. Many blockchain platforms (e.g., Ethereum [8]) also support smart contracts, which enable automated versions of financial concepts such as auctions, loans, swaps, decentralized exchanges, and ownership shares. The total value of the tokens held in DeFi smart contracts peaked at over \$150 billion USD in late 2021 [53, 15], with at least \$100 billion held in decentralized exchanges alone [55], and has since fallen to around \$45-50 billion as of June 2023 [15].
The DeFi ecosystem is currently fragmented over a multitude of blockchains which each offer their own choices in terms of efficiency, security, and functionality. For example, Cardano [28, 31], Solana [57], Avalanche [49], and Algorand [24] each had a valuation of more than a billion USD in early 2023. These blockchains are not designed to interoperate: i.e., there is no native mechanisms for transferring tokens between chains. However, the recent emergence of blockchain bridges [19] and dedicated interoperability platforms such as Cosmos [7] and Polkadot [33] demonstrates an increasing demand for cross-chain interoperability. A major challenge in cross-chain trades is atomicity: the transactions in a trade must be committed on all blockchains, or not at all. For example, in a cross-chain auction, the first party to transfer tokens (e.g., the auctioneer) cannot be sure that the other party will uphold their part of the trade in the absence of a cross-chain protocol. This is of particular concern in DeFi, as individual blockchain transactions are irreversible by design.
Existing solutions that provide cross-chain atomicity, including relays and hashed time locks [27, 2, 50], essentially lock the assets involved in a trade and only release (or return) them once the transactions in all of the blockchains have been committed (or aborted). While ensuring atomicity, this approach increases the financial risks for the users in the following ways: 1. the tokens held in escrow may lose value due to price fluctuations, 2. one party may abort the transaction, leading to an opportunity cost for the other parties, and 3. the DeFi platforms that facilitate the trade may be faulty or malicious . These risks are interrelated: if there is a significant change in the value of the tokens in escrow (1.), then this may cause a party to abandon a trade (2.), creating an opportunity cost to the other users. The risks identified above are a limitation of all cross-chain protocols that hold tokens in escrow – the longer the tokens are held, the longer they are exposed to DeFi’s notorious price volatility [1] and vulnerability to attacks [60]. Although the abandonment of trades can be disincentivized through the use of premiums [56], price fluctuations can be mitigated in principle if the tokens held in escrow are stablecoins [32, 9, 44] whose value is pegged to a less volatile real-world asset (e.g., the US dollar). However, a flawed stablecoin design can itself put users’ tokens at risk, as witnessed by the collapse of the Terra/Luna stablecoin in May 2022 [6], in which stablecoins worth 60 billion USD lost nearly all of their value.
Our goal is to design a practical stablecoin as a central point of exchange for cross-chain commerce. Realizing this goal requires that two main challenges are addressed. The first challenge is to support a large and growing number of blockchains efficiently, including new and relatively low-use platforms. For established platforms, Automated Market Makers (AMMs, [17]) such as Uniswap can be used to exchange tokens for stablecoins within minutes, and the stablecoins can be transferred between blockchains using a wallet on a cryptocurrency exchange such as Coinbase or Binance. However, the latter step may take several hours before the withdrawal on the receiving chain is processed (e.g., 2-3 hours on Crypto.com [12]), is subject to the exchange’s transaction fees, and the exchange is a single point of failure. Furthermore, AMMs depend on depositors to deposit collateral for each supported pair of cryptocurrencies, and this may not be available for smaller platforms – similarly, centralized exchanges may not always support transfers on smaller platforms. Just as an AMM requires depositors willing to create liquidity pairs, Xclaim [58] requires the availability of a custodian who is willing to put up collateral on the receiving chain, which can be a challenge for small blockchains. Finally, secure bridges such as zkBridge [54] either require a direct bridge to transfer stablecoins between two small blockchains, which may again be unavailable, or require that they are sent via a larger blockchain in a two-step process, which incurs additional time and transaction fee costs. The second challenge in designing a cross-chain stablecoin is to ensure that it is resilient to black swan events (e.g., 51% attacks) that affect individual blockchain platforms, and to price fluctuations that affect multiple blockchains simultaneously. This is particularly relevant if the stablecoin’s peg is maintained through tokens held as collateral on different blockchains. If the total value of the collateral drops below the intended value of the issued stablecoins, then the stablecoin may lose its peg, forcing holders to sell their tokens at a loss – we will refer to such an event as system failure.
In this work, we present CroCoDai, a stablecoin for cross-chain commerce that addresses the above challenges. The key insight is to back CroCoDai with collateral tokens from a multitude of blockchains, using the same principles as the existing Dai stablecoin system (DSS) [38] but without being restricted to the Ethereum ecosystem. To that end, CroCoDai manages communication between the different blockchains through a governance layer which can be implemented through a quorum of trusted nodes, or on top of an existing blockchain. By using the governance layer to create/destroy the same amount of stablecoins on different chains, CroCoDai supports cross-chain transfers by design. Crucially, CroCoDai allows tokens on smaller blockchains to be backed by collateral on larger chains, thus providing stablecoin support to emerging platforms and addressing the first challenge. Unlike Polkadot or Cosmos, CroCoDai achieves its goal without requiring intrusive changes to existing blockchains. To demonstrate how well CroCoDai addresses the second challenge, we present a technical analysis that shows that attacks on a limited number of blockchain platforms can be tolerated as long as the system is sufficiently overcollateralized. We also show that attacks against a limited number of price oracles (which determine the value of the collateral) can be tolerated – this is not a challenge specific to multi-chain settings, so this security property holds for the single-chain DSS as well. Next, we provide evidence for the greater resilience of a multi-chain stablecoin to price fluctuations through an empirical analysis based on the collected data from several periods, including the collapse of the FTX exchange which led to severe depreciation of the Solana cryptocurrency. Specifically, we compare (1) a portfolio that approximates the current state of the DSS’ collateral to a similar portfolio with tokens from multiple chains, and (2) a portfolio that is optimized for low failure risk using the tokens supported by the DSS to one with tokens from multiple chains. Finally, we provide a reference implementation for CroCoDai that uses a dedicated relay chain, and use experiments to show that the computational overhead of CroCoDai is small. The code/data used in this work have been temporarily redacted to preserve anonymity, but they will be open-sourced at a later date to facilitate reproducibility and reusability.
Contributions
Our work makes the following contributions:
  • We present CroCoDai, a cross-chain stablecoin – we present two design variants: one uses a quorum of nodes as a trust anchor; the other uses an existing blockchain (Section 4).
  • We demonstrate CroCoDai’s security through a technical analysis that demonstrates that CroCoDai can tolerate a limited number of compromised blockchains or price oracles (Section 5.1), and an empirical analysis in which we compare single-chain portfolios to portfolios that use collateral from multiple chains, and find that the latter has lower failure probabilities. Specifically, the failure probability is roughly a third lower for optimized pure-crypto portfolios (Sections 5.2 and 5.3).
  • We present three implementations of CroCoDai subvariants that use the quorum-based approach and compare their gas costs and processing times, to demonstrate CroCoDai’s practicality and to make its generality/performance trade-offs explicit (Section 6).
The remainder of the paper is structured as follows. Section 2 provides background on DeFi and related work. Section 3 presents the system model. Section 4 presents the design of CroCoDai. Section 5 contains the analysis of CroCoDai’s resilience against black swan events and price fluctuations. Section 6 details the three subvariant implementations and corresponding experimental results. Section 7 discusses future work and concludes the paper.
2 Background
2.1 General Background
Blockchains & Tokens. DeFi uses cryptographically secured, decentralized data structures to replace trusted intermediaries. Blockchains are by far the most prominent example of such data structures. A blockchain is an ordered list of transactions grouped into blocks. The principal function of transactions in DeFi is to transfer tokens, which represent ownership of assets. There are two common types of tokens: fungible tokens – also commonly called cryptocurrencies – and non-fungible tokens (NFTs) [13]. Fungible tokens of the same type are identical, whereas each NFT represents a unique asset.
In the first cryptocurrency platform, Bitcoin [45], each transaction transfers some amount of the native cryptocurrency (bitcoin, or BTC) between parties. In Ethereum [8], transactions were extended from transferring the native cryptocurrency (ETH) to the creation or execution of smart contracts, i.e., on-chain software programs. Smart contracts can store data, hold tokens, and transfer tokens according to the contract’s internal logic, e.g., to implement an auction. A major use case for smart contracts is to define new cryptocurrencies that are different from the blockchain’s native cryptocurrency. In Ethereum, this is particularly facilitated by the ERC20 token standard [51] which defines a common set of smart contract functions required by any new cryptocurrency. Smart contracts can also support NFTs, e.g., via OpenSea or Mintable.
Stablecoins. A stablecoin is a cryptocurrency whose value is pegged to an underlying fiat currency, e.g., the US dollar. [^1] Cryptocurrencies whose value is pegged to a real asset such as gold, e.g., the PAXG token, are typically not considered stablecoins, e.g., they are not labelled as such by CoinMarketCap. There are multiple types of stablecoins [44, 32], of which the most common ones are:
  1. 1.
    Fiat-backed stablecoins, which are backed by fiat currency held largely in cash or cash equivalents. The essential guarantee that underpins the price stability of a fiat-backed stablecoin is that in case of system shutdown, each stablecoin can be used to redeem a corresponding amount of the pegged fiat currency.
  2. 2.
    Crypto-collateral-backed stablecoins (or crypto-backed stablecoins for brevity), which are backed by other cryptocurrencies that are stored in one or more smart contracts. In this case, the guarantee is that in case of system collapse, each stablecoin can be redeemed for a basket of tokens whose value equals one unit of the pegged currency.
  3. 3.
    Algorithmic stablecoins, which do not use reserves or collateral and which attempt to maintain a peg by adjusting the supply of a paired token [44].
A multitude of fiat-backed stablecoins exist, the most prominent (as of June 2023) being Tether (USDT), USD Coin (USDC), and Binance USD (BUSD). The largest crypto-backed stablecoin is the Dai stablecoin (DAI), which is maintained by the MakerDAO foundation and which we discuss in more detail in Section 2.2. During our observation period (as discussed in Section 5), most fiat- and crypto-backed stablecoins did an overall sound job in maintaining their peg with the US dollar: the observed price ranges of DAI, USDT, USDC, and BUSD were [0.9921, 1.0062], [0.9834, 1.0005], [0.9991, 1.0073] and [0.9988, 1.0126] respectively. The analysis in [37] suggests that fiat-backed stablecoins have a trade-off between price stability and risk of system failure in the sense that better arbitrage improves the former at the cost of increasing the latter.
Fiat-backed stablecoins introduce a trust assumption, namely on the agents who handle and store the fiat currency. In Tether, this trust assumption has been widely questioned – accusations include Tether using fiat collateral to buy high-risk assets [20] and creating unbacked stablecoins to prop up the price of bitcoins [25]. Furthermore, algorithmic coins have nothing to underpin them in case of a death spiral [44, 6], as witnessed by the loss of over 60 billion USD when the Terra/Luna stablecoin system failed. As such, in the rest of this paper we will focus on crypto-backed stablecoins such as Dai.
2.2 The Dai Stablecoin System (DSS)
In this section, we describe the core features of the existing single-chain DSS. Although there are documents and code repositories that describe the DSS in its various stages of evolution, e.g., [38, 39, 40, 44, 3, 10], we focus on the GitHub code [40] and the description of the multi-collateral DSS [39].
Dai Tokens. In the DSS, Dai tokens are minted by users who deposit cryptocurrency tokens as collateral. In the first iteration of the Dai stablecoin (the single-collateral Dai or “Sai”), only ETH tokens could be deposited for this purpose. The procedure for a user, Alice, who wants to mint new Sai tokens is as follows. First, Alice must send a transaction to a Dai vault to create a Collateralized Debt Position (CDP) [38]. Next, Alice must send ETH tokens as collateral to the contract to fund the CDP. If the CDP is funded, then Alice can use it to create Dai stablecoins and withdraw them from the CDP to her wallet, incurring debt. At any later time, she can send Dai tokens back to the CDP to repay her debt. The above can be generalized to a multi-collateral (but single-chain) DSS by associating each CDP with a token type. Debt increases exponentially over time, and the rate at which it increases is a system parameter called the stability fee.
Liquidations. At all times, Dai CDPs must be overcollateralized, i.e., the ratio of the total value of the Alice’s collateral to her debt must exceed some threshold called the liquidation ratio [3, 38]. By default, the liquidation ratio in Dai is set to $1.5$ – i.e., Alice can withdraw Dai stablecoins worth at most $\frac{2}{3}$ of the total value of the collateral in their CDP [44]. Similarly, Alice can return Dai and withdraw collateral from her CDP as long as the above condition holds. If a CDP has no associated debt, then it can be closed and all remaining collateral can be withdrawn. If, at any time, a CDP’s ratio of its collateral value to its debt is insufficient, then Alice’s CDP is underwater and can be liquidated. CDPs can go underwater either because the value of the collateral has dropped due to price fluctuations, or because the stability fee has caused the debt to become too high. To avoid liquidation, Alice must either deposit additional collateral, or raise Dai to clear the position when it is at risk – if this fails, a collateral auction can be started by other users. During a collateral auction, other users can bid to buy (portions of) Alice’s collateral using Dai stablecoins. Once enough Dai has been raised to repay the CDP’s total debt, the remaining collateral (if any) is sent back to Alice, minus a penalty.
Price Feeds. To determine the value of deposited collateral, MakerDAO uses information from price feeds supplied by oracles [35]. The number of price feeds used by the DSS has increased from 14 between 2017 and November 2019 [26] to 20 currently [43]. Smart contracts that maintain CDPs – vaults – periodically update token prices by setting each price equal to the median of a set of recent prices for the associated token.
Dai Savings Rate. Due to the stability fee, the total amount of debt in the protocol may grow to exceed the total number of stablecoins in circulation. For example, if Alice’s debt of 100 Dai has grown to 105 Dai, then clearing her debt and closing the CDP would (in principle) leave 5 Dai in the vault. These surplus Dai are part of the Dai Savings Rate (DSR) mechanism, which allows users who deposit stablecoins in a DSR contract to periodically withdraw surplus Dai tokens from associated vaults. The DSR helps MakerDAO to keep the value of Dai soft-pegged to the dollar by reducing or decreasing the interest rate: if Dai becomes less valuable than the dollar, then the interest rate is increased to incentivize users to buy Dai to deposit in the DSR contract (and decreased for higher Dai values).
Governance. All major decisions about the functioning of the protocol are taken by users who hold MKR, which is MakerDAO’s governance token. Such decisions include 1) verifying and adding smart contracts that implement any of the functionalities above, 2) changing the Dai savings rate, and 3) occasionally changing other system-level parameters. The following are the main system-level parameters (see page 15 of [39]): 1. the debt ceiling for each token type, i.e., the maximum amount of total debt that can be created using a given collateral type, 2. the stability fee, 3. the liquidation ratio, 4. the liquidation penalty, 5. the duration of collateral auctions, 6. the duration of auction bids, and 7. the minimum auction step size (which prevents bids that are only marginally above the previous bid, which would waste transaction fees).
Out-of-Scope Concepts. Some of the concepts in the DSS are beyond the scope of this paper. In particular, these are 1) debt auctions, which occur when a collateral auction fails to raise enough Dai to cover the CDP’s debt, 2) surplus auctions, which occur when the amount of Dai left in vaults due to auctions and stability fees exceeds some threshold, and 3) (emergency) shutdowns, which can be triggered after a catastrophic event (e.g., system failure) and during which the total remaining collateral is distributed to Dai holders.
2.3 Existing Cross-Chain Approaches
Cross-chain deals [27] provide a modeling framework to describe sequences of cross-chain token transfers. A cross-chain deal consists of five phases: clearing, escrow, transfer, validation, and commitment. The goal of cross-chain deals is to guarantee atomicity, i.e., that either all transfers occur or none do. In [27], two frameworks for cross-chain deals are proven to satisfy technical safety (all honest users end up with an acceptable payoff) and liveness (no assets are escrowed forever, and if all users are honest then all transfers eventually happen): 1. hashed time lock contracts, 2. dedicated blockchains for commit messages . In cross-chain deals, the cost of a participant abandoning the protocol before completion (sore-loser attacks [56]) can be made explicit – this form of misbehavior can be disincentivized through a premium [18] paid by the parties that abandon the protocol before termination.
Cosmos [7] and Polkadot [33] are blockchain platforms that were specifically developed with blockchain interoperability in mind, but which do not trivially interoperate with existing blockchains outside their ecosystems. Furthermore, Xclaim [58] and zkBridge [54] claim to provide secure cross-chain transfers, but have limitations when transferring to/from smaller blockchain platforms as discussed in the introduction.
3 System Model
3.1 Entities
Our system model has the following types of entities. In the following, we use the notation $[n]$ for $\{1,\ldots,n\}$. A list of symbols can be found in Table 4 in Appendix A.
  • Token: a type of cryptocurrency or an NFT (although typically the former due to the better availability of price information for cryptocurrencies). We denote the number of different tokens in the system by $M\geq 2$.
  • Governance Token: this is a type of cryptocurrency that gives holders the right to vote on decisions that affect the functioning of the crypto-backed stablecoin. The governance token cannot be used as collateral for the stablecoin.
  • Blockchain (or chain): an immutable, distributed ledger that tracks the ownership of its tokens. We assume that each chain either supports smart contracts or that a wrapped version of its native tokens exist on a chain that does (e.g., WBTC [53]). We denote the number of supported chains by $B\geq 2$, and the chain of token $m\in[M]$ by $b_{m}\in[B]$.
  • Consensus Node: is responsible for maintaining the blockchain, e.g., through the creation and/or validation of new blocks. The nature of the consensus nodes depends on the blockchain’s consensus protocol – examples include variations of proof-of-work, proof-of-stake, or Byzantine Fault Tolerance (BFT) protocols. Each consensus node has some amount of consensus power – e.g., processing power in proof-of-work, staked tokens in proof-of-stake, or a private key in a BFT protocol.
  • User: owns blockchain tokens and seeks to exchange them for tokens on another chain. We denote the number of users in our system by $U\geq 1$.
  • Smart contract: a software program on a blockchain that may hold tokens which can be sent to other contracts or wallets depending on the program’s internal logic.
  • Vault: a smart contract or address that holds tokens as collateral, and to/from which tokens can be sent/received leading to the possible creation/destruction of stablecoins managed by another smart contract.
  • Price Oracle: a type of user who provides price information to the system [35], e.g., Chainlink [5]. We denote the number of price oracles in our system by $O\geq 1$. Each oracle $o\in[O]$ can provide price information for any selection of tokens. Let $p_{m\,o\,t}$ be the price of token $m\in[M]$ reported by oracle $o\in[O]$ at time $t\geq 0$. A price feed is a sequence of prices for a given token $m$ reported by an oracle $o$.
3.2 Token Price Model
[…]
To model the evolution of token prices over time, we use Geometric Brownian Motion (GBM). This is a common choice in the financial literature on asset prices [47] and related work on stablecoins, e.g., by Salehi et al. [48], particularly due to the GBM’s role in the Black-Scholes formula for option prices. A stochastic process $S_{t}$ that follows GBM is defined using a stochastic differential equation:
$dS_{t}=\mu S_{t}dt+\sigma S_{t}dW_{t},$ (1)
where $\mu$ determines the drift of $S_{t}$ (i.e., its expected change per unit of time), $\sigma$ represents the volatility of $S_{t}$ (i.e., the variance of the change), and $W_{t}$ is the Wiener process that represents standard Brownian motion [47]. In this model, each token price process $p_{m\,t}$, $t\geq 0$, can be represented using a process defined as in Equation 1. Essentially, this means that the logarithm of the relative difference of token prices, $\log\left(p_{m\,t+t^{\prime}}/p_{m\,t}\right)$, is normally distributed for all $t^{\prime}>0$ with mean and variance dependent on $t^{\prime}$. Unlike [48], we consider a setting with multiple assets; i.e., we have a stochastic process $S_{m\,t}$ for each token $m$ with its own mean $\mu_{m}$ and variance $\sigma^{2}_{m}$. In this setting, we must also consider the covariance between the token prices [47], as cryptocurrency prices are often highly correlated – we will discuss this in more detail in Section 5. More details about how the standard GBM can be used to simulate asset prices can be found in Appendix B.
The standard GBM as decribed above is based on the normal distribution, i.e., $\log(S_{t+t^{\prime}}/S_{t})$ has a normal distribution. However, it is known from the academic literature that the probability distribution of stock price fluctuations is usually more heavy-tailed and that Student’s $t$-distribution is often a better fit [29, 34]. This is also visible in the data for price changes in our dataset – see Figure 5 in Appendix C. As such, we will focus on the $t$-distribution in our analysis. Although the $t$-distribution does not allow for an analytical solution for the asset price change after each individual time slot (e.g., as per (5) in Appendix B), we can use Monte Carlo simulation to analyze its behavior, as we discuss in more detail in Section 5.
3.3 DSS Execution Model
Let $\mathcal{J}_{t}$ denote the set of all CDPs at time $t$, and $c_{j\,t}$ the total number of collateral tokens held in CDP $j\in\mathcal{J}_{t}$ at time $t$. Let $D_{j\,t}$ be the outstanding debt resulting from Dai withdrawals/deposits from/to CDP $j\in\mathcal{J}_{t}$. Each CDP is associated with a specific token type – let $y_{j}$ be the token type of CDP $j$. Debt is subject to fees, i.e., $D_{j\,t+1}=(D_{j\,t})^{1+\delta}$, where $\delta$ represents the stability fee. We denote the number of stablecoins held by user $u\in[U]$ at time $t$ by $h_{u\,t}$. We note that, in principle, $\sum_{u\in[U]}h_{u}\leq\sum_{j\in\mathcal{J}}D_{j\,t}$, as stablecoins are created and transferred to a user account either through withdrawal from a CDP (in which case the created stablecoins and debt are equal) or from a vault through the Dai savings rate (in which case the created stablecoins cannot exceed the excess debt). However, we will consider attacks in Section 3.4 that can cause this assumption to be violated.
We formalize the liquidation criterion as follows: for each CDP $j\in\mathcal{J}_{t}$ and $t\geq 0$, it must hold that $p_{y_{j}\,t}c_{j\,t}/y_{j\,t}>\gamma,$ where $\gamma$ is the liquidation ratio – if this condition does not hold, then CDP $j$ can be liquidated through a collateral auction. We assume that under standard conditions (i.e., the absence of an attack or sudden price collapse), collateral auctions are concluded successfully, i.e., a sufficient amount of stablecoins is raised to close the position. No user can create debt if the result would violate the liquidation criterion. We also note that price fluctuations do not affect the liquidation criterion, nor does the Dai savings rate as the proportional effect on all CDP debts is the same.
We denote the debt ceiling for each token type $m\in[M]$ by a fraction $\zeta_{m}\in[0,1]$, i.e., the total debt in CDPs of type $m$ relative to the system’s total debt cannot exceed $\zeta_{m}$. (Debt ceilings are specified in absolute terms in DSS contracts, but we assume that the difference is negligible for short periods, e.g., for escrow.) The debt ceilings are used to enforce $\sum_{\begin{subarray}{c}j\in\mathcal{J}_{t}\\ y_{j}=m\end{subarray}}D_{j\,t}<\zeta_{m}\sum_{j\in\mathcal{J}_{t}}D_{j\,t}$ for all $m\in[M]$. Let $D_{t}^{*}=\sum_{j\in\mathcal{J}_{t}}D_{j\,t}$ and $C^{*}_{t}=\sum_{j\in\mathcal{J}_{t}}p_{y_{j}\,t}c_{j\,t}$ denote the total debt and the total collateral value across the system at time $t$, respectively.
3.4 Threat Model
We make the following assumptions about the system entities. We assume that a majority of the governance token holders are honest. For each underlying blockchain, we assume that it is either compromised or not. If it is not compromised, we assume that it guarantees safety and liveness for all of its transactions. However, both safety and liveness can be violated in a compromised blockchain – in particular, transactions in which collateral is sent to/from a CDP can be overturned. For the price oracles, we assume that a majority is online and honest at all times, but that dishonest oracles can report arbitrary prices.
3.5 System Requirements
The central requirement that we focus on in this paper is the following:
  1. (*)
    Full Backing. At any time $t\geq 0$, the ratio of the value of the collateral tokens to the total number of stablecoins in the system exceeds a safety threshold $\theta$, i.e.,
    $C^{*}_{t}/D^{*}_{t}>\theta.$ (2)
We assume that $\theta>1$, i.e., system failure occurs when the ratio of the collateral value to the debt gets sufficiently close enough to $1$. The motivation behind this requirement is as follows: the fundamental guarantee that underpins crypto-backed stablecoins is that in the event of a system shutdown, there is enough collateral in the vaults to refund each user who holds $h$ stablecoins with a basket of tokens whose value equals $h$ units of the pegged asset. If price fluctuations or an attack would cause this requirement to be violated, then we can reasonably expect the stablecoin to lose its peg. This would be detrimental to the user experience as their tokens held in escrow would depreciate, causing financial losses.
Due to random price fluctuations, we cannot guarantee with absolute certainty that (2) will hold in all circumstances. However, we use our analysis to show that for reasonable choices of $\theta$ and the liquidation ratio $\gamma$, the probability that (2) is violated after a medium-length period (i.e., one day) is very small for historical data. We do this both for portfolios that approximate the current state of the DSS (as of late April 2023), and portfolios that are optimized for minimal risk given a selection of possible single-chain or cross-chain tokens. In Section 5, we also investigate under which conditions (2) continues to hold in the event of a black swan event such as a sudden token price collapse or a 51% attack on a blockchain.
Furthermore, we require the following properties under the threat and token price model.
  1. 1.
    Safety. Users do not lose tokens during stablecoin transfers.
  2. 2.
    Liveness. Stablecoin transfers eventually happen.
  3. 3.
    Safety of Collateral (informal). The probability that a minority of corrupted oracles cause a CDP that is not underwater to be inadvertently liquidated is small.
More formally, for the last property we consider a minority of corrupted price oracles whose goal is to convince the protocol to accept a price $p^{*}$ instead of a true base price $p$ – safety of collateral then holds if their probability of success vanishes exponentially as $|p-p^{*}|$ increases (see also Theorem 5.3 in Section 5.1).
As we will see in Section 5, the former two properties largely follow from results in the existing scientific literature. The latter is a result that we prove for the multi-chain setting in Section 5, but which is also applicable in the single-chain setting.
4 CroCoDai: A Cross-Chain Stablecoin
Figure 1: Overview of the single-chain DSS.
In this section, we present CroCoDai, which extends the functionality of the single-chain DSS (depicted in Figure 1) to a multi-chain setting (depicted in Figure 2). Our main design goal is to have stablecoin tokens that are backed by collateral on multiple blockchains, and which can be moved securely and efficiently between the different chains. This is non-trivial as transactions between different blockchains are not atomic by default, so our design incorporates a secure cross-chain protocol. In the following, we will refer to blockchains that store tokens as coin chains (as per [27]). The communication between the different blockchains is managed by a relay chain, which may be a coin chain or a separate chain. The main difference between the single-chain DSS and CroCoDai is the inclusion of relay nodes – depicted with double circles in Figure 2 – which is a non-intrusive change as the other DSS contracts remain unchanged.
In the following, we first discuss in detail which components as discussed in Section 2.2 remain local to each chain and which require interaction with the relay chain. We then discuss the essential components of the smart contracts on the coin chains and the relay chain. Finally, we discuss two ways to implement the relay chain: one that relies on a quorum of trusted nodes, and one that uses an existing blockchain. We will present three implementations of the former in Section 6.
4.1 Overview
Local Operations. In CroCoDai, each coin chain keeps track of the balance of stablecoins in accounts on that chain. The stablecoins can be exchanged locally between accounts. Furthermore, each coin chain has its own vaults that allow users to create CDPs using tokens on that blockchain. Since CDPs are local, liquidation auctions and savings rate mechanisms are also processed locally. Finally, price feeds are still predominantly local, although price feeds from the relay chain can add an additional source of price information.
Cross-Chain Operations. The relay chain has four main functions in CroCoDai. The first is to enable cross-chain transactions. In a cross-chain transaction, stablecoins are burned on one coin chain and an equivalent number is minted on another chain. The second function is to add/remove addresses that are authorized to call security-critical functions such as mint or burn on another contract – we call such addresses wards as per DSS terminology. Periodically, smart contracts may be added or deprecated – e.g., 1. to add support for an emerging blockchain platform, 2. because a critical bug has been found in an existing smart contract, 3. because a chain has undergone a hard fork that interferes with the functioning of the contract, or 4. to implement a more efficient version of a component (e.g., upgrading the collateral auction mechanism from Cat.sol to Dog.sol in the DSS). The third function is to update the system parameters discussed in Section 2.2. The final function is to provide an additional source of token price information to vaults through relay chain oracles – this may particularly benefit smaller blockchains that lack a mature DeFi ecosystem.
The mechanism for cross-chain transfers is as follows: a node that holds stablecoins on a coin chain must create a transaction that stores the tokens to be transferred in escrow – furthermore, the transaction takes as input two strings to represent the target blockchain and the address on the target blockchain. This is then added to a list where it awaits commitment. The nodes in the relay network monitor the coin chains, and once they detect a cross-chain transfer request, they will sign a commit message. When the coin chain receives the commit message, it will commit the transaction and burn the tokens held in escrow. A commit message is also sent to the target blockchain, upon which it will mint an equivalent number of stablecoins. If the transaction is aborted for any reason, then the stablecoins in escrow are returned to the sender.
Figure 2: Overviews of CroCoDai, a multi-chain stablecoin system based on the DSS. The relays (double circle) are components that are not shared with the single-chain DSS.
4.2 Coin Chain Components
Each coin chain has the following components. Coin: implements a contract that is similar to Dai.sol in the DSS, and processes transfers of stablecoins between users on the coin chain. There is only one coin contract per chain. Vault: implements contracts that are similar to Vat.sol and Join.sol in the DSS, which allow users to create CDPs and withdraw/deposit collateral and stablecoins. There can be multiple vaults per chain. Auction: implements auction contracts that liquidate the collateral of underwater CDPs, with any appropriate auction logic (not necessarily limited to Cat.sol or Dog.sol) in the DSS. There can be multiple auction contracts per chain. Savings: implements a savings rate contracts that are similar to Pot.sol and Jug.sol in the DSS. Bottom-Level Oracle: implements a price feed using a contract similar to Spot.sol in the DSS, combined with a Medianizer. Bottom-Level Relay: validates messages from the relay chain. It processes 1. adding wards to each of the aforementioned smart contracts, 2. incoming cross-chain transfers, 3. commitments of outgoing cross-chain transfers, and 4. updates of system parameters.
4.3 Relay Chain Components
The relay chain has the following components. Governance: allows users to transfer governance tokens. Users who hold governance tokens can vote on system-level actions such as adding wards, processing cross-chain transfers, and updating system-level parameters. Top-Level Oracle: The relay chain may implement system-level price feeds to complement the price feeds on the coin chains. Top-Level Relay: To process cross-chain transfers, the relay network must be able to read the state of the coin chains, in particular the list of cross-chain transfer requests.
4.4 Relay Chain Designs
In this section, we discuss two possible implementation designs of the relay chain. The first relies on a quorum of trusted nodes with known public keys – each trusted node runs an off-chain (full or light) blockchain client for each of the underlying chains to check the validity of cross-chain transfers, and the nodes maintain a distributed ledger for governance decisions. The second uses an existing blockchain, and uses smart contracts that implement the consensus logic of the underlying chains on-chain to verify cross-chain operations.
Off-Chain Blockchain Clients. In this approach, the stablecoin system is maintained by $n$ relay nodes with known public keys. The relay nodes use their (light) clients to detect and validate cross-chain transactions, and generate signatures using their private/public key pair to commit cross-chain transfers and approve governance decisions. In this setting, the only functionality of the bottom-chain relay contracts is to verify digital signatures. If different chains use different built-in signature schemes, or the same scheme with a different parametrization, then the relay nodes may need to generate multiple signatures for a single cross-chain transfer (this is safe if the threat model of Section 3.4 holds when each signature is generated). We discuss three different implementations of this approach in Section 6.
To ensure security, we note that it is vital that the relay nodes are controlled by independent entities: e.g., in the attack against the Ronin bridge, 5 out of 9 validators were successfully attacked [30], of which 4 belonged to the same entity [36].
On-Chain Consensus Logic. In this approach, the consensus logic of the coin chains is implemented in the top-level relay contracts, and the consensus logic of the relay chain is implemented in the bottom-level relay contracts. Essentially, the functionality of a light client – i.e., validating block headers and proving the state of a contract using the header – is implemented inside the smart contracts. This is the approach taken in, e.g., Xclaim [58]. For example, in Bitcoin’s Nakamoto Consensus [45] this means the ability to compute header hashes and compare them to the proof-of-work difficulty, whereas in Ethereum’s proof-of-stake it means validating that nodes who control a sufficient amount of stake have approved a block. Techniques to increase the efficiency of light clients have been proposed in the literature for both proof-of-work (e.g., Bitcoin [14]) and proof-of-stake blockchains (e.g., Algorand [42] and Mina [4]). Another variation of this approach is to encode the consensus logic of the blockchains in a zero-knowledge proof, e.g., using the zkBridge protocol [54]. Finally, the data on the different coin chains can be integrated using TrustBoost [52].
Although this design avoids a trust assumption on a quorum of relay nodes, current iterations of such designs have a large computational overhead, typically requiring several kilobytes of data sent to relay contracts and more than 10 seconds to generate signatures [42, 54]. As such, we focus on the off-chain design in Section 6.
5 Analysis
In this section, we analyze whether our system satisfies the requirements of Section 3.5. In Section 5.1, we focus on the technical properties of safety and liveness, full backing under the threat model, and safety of collateral. We then focus on full backing under the price fluctuation model: we present our approach in Section 5.2, and the evaluation using empirical data in Section 5.3. The proofs of the theorems in this section can be found in Appendix D.
Our analysis shows that CroCoDai is resilient to serious failures of multiple coin chains, as long as the over-collateralization of the CDPs, $\gamma$, is high compared to the total debt ceiling of the affected chains (which we will denote by $\zeta^{\prime}$). We emphasize that CroCoDai is therefore fundamentally more resilient than a single-chain stablecoin such as the DSS, which cannot tolerate the failure of its single coin chain. Our analysis also shows that a broader selection of collateral tokens leads to a lower probability of system failure than a single-chain system such as the DSS. Although the findings of our price analysis strongly depend on the choice of data, we emphasize that we included a period of notoriously high volatility (i.e., the FTX collapse) and that the risks would appear smaller if we used a less volatile period.
5.1 Safety Analysis
Technical Safety and Liveness. For local operations, the consensus protocol of the coin chain on which the local transfer occurs ensures the safety and liveness of operations on this chain – see, e.g., [22]. This includes local stablecoin transfers, operations on collateral, auctions, and interest withdrawals. For cross-chain operations, safety is violated if tokens in a bottom-level relay are burnt without a corresponding amount being minted on the receiving chain, and liveness is violated if tokens are held in escrow forever. However, the relay chain satisfies safety and liveness in our threat model, i.e., the transfer is eventually committed and coins are burned and minted on the outgoing and incoming chains, respectively. An alternative way to reach this conclusion is to view cross-chain transfers in CroCoDai as simple cross-chain deals [27] with the sender as the only party who votes, and the relay chain as a certified blockchain – depending on the design (as discussed in Section 4.4), either $12f+$ 1 nodes or the smart contracts on the relay chain sign the commit vote.
To ensure safety and liveness, we emphasize that it is essential that the relay network is not compromised. A compromised relay network can burn coins without creating a corresponding amount on the receiving chain, violating user safety. Likewise, if the relay network is unable to process requests because more than $f$ relay nodes are offline or malicious, then the system does not satisfy liveness as cross-chain transfer requests may never happen. We also note that if the relay network is compromised, then full backing can be violated by creating an incoming transfer request without a matching outgoing transfer request, thus creating tokens without collateral. Although the security of the relay chain is therefore essential to CroCoDai, we investigate the impact of compromised coin chains on full backing below.
Sudden Crash of Tokens. Occasionally, the price of a cryptocurrency token may suddenly collapse, e.g., because of the discovery of a serious bug in one of its smart contracts, or because of off-chain events such as a bankruptcy or a legal inquiry into the organization responsible for its maintenance. As an example: the bankruptcy of Celsius caused the value of the CEL token to collapse, whereas the inquiry into FTX did the same for the FTT token and Serum (SRM). In this case, the collateral auctions are likely to fail to generate enough revenue and the system’s over-collateralization is under threat. In the following, we consider a worst-case scenario: the price of some tokens falling to zero within a single time slot, i.e., between a given time slot $t^{\prime}$ and time slot $t^{\prime\prime}$ – for simplicity, we assume $t^{\prime}=0$ and $t^{\prime\prime}=1$. We show that CroCoDai remains fully backed at time $t^{\prime\prime}$ if it is sufficiently over-collateralized and if the total debt created using tokens on the affected chains is limited.
Formally, let $M^{\prime}\subset[M]$ denote the set of affected tokens, and $\zeta^{\prime}=\sum_{m\in M^{\prime}}\zeta_{m}$ the total debt ceiling of these tokens. Let
$C^{-}_{t}=\sum_{\begin{subarray}{c}j\in\mathcal{J}_{t}\\ y_{j}\in M^{\prime}\end{subarray}}p_{y_{j}\,t}c_{j\,t},\;\;\;\;\;C^{+}_{t}=\sum_{\begin{subarray}{c}j\in\mathcal{J}_{t}\\ y_{j}\notin M^{\prime}\end{subarray}}p_{y_{j}\,t}c_{j\,t}$
i.e., $C^{-}_{t}$ represents the total value of the collateral tokens affected by the collapse, and $C^{+}_{t}$ the total value of the unaffected tokens. Similarly, let
$D^{-}_{t}=\sum_{\begin{subarray}{c}j\in\mathcal{J}_{t}\\ y_{j}\in M^{\prime}\end{subarray}}D_{j\,t},\;\;\;\;\;D^{+}_{t}=\sum_{\begin{subarray}{c}j\in\mathcal{J}_{t}\\ y_{j}\notin M^{\prime}\end{subarray}}D_{j\,t}$
so that $D^{-}_{t}$ represents the debt associated with affected CDPs and $D^{+}_{t}$ the debt of unaffected CDPs at time $t$. We consider an attack that occurs at time $1$, so that $C^{-}_{1}=0$, $C^{+}_{1}=C^{+}_{0}$, $D^{-}_{1}=D^{-}_{0}$, and $D^{+}_{1}=D^{+}_{0}$. We can then prove the following theorem (see Appendix D):
Theorem 5.1.
[…]
Individual Compromised Coin Chain. If a majority of consensus power on a coin chain $b^{\prime}\in[B]$ is compromised then the consensus nodes could perform a 51% attack in the following way: send some amount of stablecoins to another chain, wait for the transaction to be committed by the relay network, then fork the compromised coin chain and undo the outgoing transfer request and the burning of the blockchain tokens. Similarly, they could undo the funding of a CDP. As a consequence, stablecoins are created on another chain without backing collateral. In the worst case, this attack would be performed simultaneously on all of the token types supported by the compromised chain $b^{\prime}$. Let $M^{\prime}=\{m\in[M]:b_{m}=b^{\prime}\}$ be the set of affected tokens. We again consider a worst-case attack that has been completed at $t=1$. Let $h^{\prime}$ be the total amount of stablecoins created in the attack, so that $D^{-}_{1}=D^{-}_{0}+h^{\prime}.$ We can then prove the following theorem (see Appendix D):
Theorem 5.2.
[…]
Compromised Price Feeds. Finally, we consider an attack in which a dishonest minority of price oracles tries to corrupt the price maintained by a vault. This could cause CDPs to be liquidated even if they are not underwater, thus violating safety of collateral.
Let $p$ be the base price for token $m\in[M]$ at time $t\geq 0$ (we omit $m$ and $t$ from the notation in this section as they are not relevant to the analysis). Let $p_{o}$ be the price returned by oracle $o\in[O]$, so that $p_{o}=p+\Delta_{o}$ as discussed in Section 3.2. In the following, we assume some subset $O^{\prime}\subset[O]$ is controlled by malicious parties who aim to convince a vault in CroCoDai that of a price $p^{\prime}$ that is different from the base price $p$. Let $p^{*}$ be the actual median price among all price feeds, including those that have been distorted. We can then prove the following theorem (see Appendix D):
Theorem 5.3.
If $|O^{\prime}|<M/2$, then the probability that the difference between the base price $p$ and the median price $p^{*}$ exceeds a constant $c>0$ vanishes exponentially as $c$ increases.
5.2 Financial Analysis: Approach
In this section, we investigate the probability that CroCoDai remains fully backed even in the presence of standard price fluctuations. In particular, we will estimate this probability for several choices of supported collateral tokens: 1) a single token type, 2) tokens that exist on the Ethereum blockchain, and 3) tokens that exist on a variety of independent blockchains. The second probability corresponds to the safety of the single-chain DSS, whereas the third option corresponds to the safety of CroCoDai. The purpose of this section is to demonstrate that the third option has considerably lower risk than the second, using multi-chain collateral baskets that are similar to those observed for the DSS in practice (see Table 8 in Appendix E). We discuss our approach in Section 5.2 and present our empirical results in Section 5.3.
As is common in financial analysis, we will refer to a basket of collateral tokens as a portfolio, and to the relative increase in value of a token type over one time slot as a return. We will also discuss a method based on convex optimization that allows protocol engineers to set the system parameters such that the risk of system failure is minimized.
Problem Statement. Let $\mathcal{J}_{m\,t}=\{j\in\mathcal{J}_{m}:y_{j}=m\}$ be the set of CDPs at time $t$ whose token type is $m$, and $\alpha_{m\,t}=\sum_{j\in\mathcal{J}_{m\,t}}c_{j\,t}$ be the total number of collateral tokens of type $m$ at time $t$. In the following, we assume for simplicity that $\alpha_{m\,t}\equiv\alpha_{m}$, i.e., that the composition of the portfolio does not change during the observation period. This is a reasonable assumption as interactions with CDPs do not occur frequently. In our model, failure of the stablecoin system occurs if, for any $t$ we have
$\sum_{m\in[M]}\alpha_{m}p_{m\,t}\leq\theta D^{*},$ (3)
[…]
In our simulation, we represent the portfolio as relative proportions of different assets. We define a vector $\vec{v}$ that denotes the value of individual assets in the portfolio as $\textbf{v}=(v_{1},v_{2},\ldots,v_{M})^{T},$ where $v_{i}$ is value of asset $i$. Similarly, we define a vector of random variables for log returns of individual assets $\textbf{X}=(X_{1},X_{2},\ldots,X_{M})^{\intercal},$ where $X_{m}$ is the random variable for log returns of asset $m$. We then simulate the price of each asset as an individual GBM with Student’s $t$-distribution for the return after each time slot, as discussed in Section 3.2. To accurately simulate price fluctuations, we must consider the correlation between the prices – to do so, we use Cholesky decomposition to compute a matrix $\mathbf{L}$ that satisfies $\textnormal{Cov}(\mathbf{X})=\mathbf{L}\mathbf{L}^{T}$. A matrix of random samples $\mathbf{R}$ can then be generated as $\mathbf{R}=\mathbf{LT}$, where the matrix $\mathbf{T}$ contains a $t$-distributed random sample for each asset type and simulated time slot. In Appendix B we discuss alternative simulation methods based purely on historical data and the standard GBM.
Choosing the System Parameters. In CroCoDai, the values $\alpha_{m}$, $m\in[M]$, which represent the fraction of each asset $m$ in the total portfolio, are not set by the protocol engineers – instead, they emerge naturally as users create CDPs and withdraw stablecoins using the currencies of their choosing. However, protocol engineers can influence the range of potential portfolios by adjusting the debt ceiling $\zeta_{m}$ for each token $m$.
When optimizing the portfolio of cryptocurrencies for a stablecoin, the protocol engineers’ main goal should be to minimize the portfolio’s volatility. In principle, a positive mean drift in the returns is beneficial, as that would imply that the over-collateralization grows naturally over time. However, the mean return over our observation period was close to zero, and for risk analysis, with cross-chain commerce as our driving use case, we focus mainly on short-term price changes. In this context, the mean return is not a meaningful optimization criterion, so we focus on minimizing the volatility instead.
Assuming a normal or $t$-distribution of returns, we can calculate the covariance matrix $\text{Cov}(\mathbf{X})$ as discussed previously. We can then use convex optimization to find a portfolio that minimizes total volatility. We can write this optimization problem as
$\begin{array}[]{rcl}\displaystyle\min_{\textbf{v}}&\textbf{v}^{T}\text{Cov}(\mathbf{X})\textbf{v},&\\[5.69046pt] \text{s.t.}&\displaystyle\sum_{i=1}^{M}v_{i}=1,&\\[5.69046pt] &\displaystyle 0\leq v_{i}\leq\lambda_{i}&\forall i\in[M],\end{array}$ (4)
where we consider $M$ different cryptocurrencies and $v_{i}$ denotes the fraction of the portfolio invested in asset $i\in[M]$. The bounds $0\leq\lambda_{i}\leq 1$ enforce maximum contributions for the different individual asset types, to avoid portfolios with a small number of low-volatility tokens (such portfolios would be vulnerable to black-swan events as discussed in Section 5.1). Given a solution $v_{1},\ldots,v_{M}$ to (4), we can enforce the optimal portfolio up to some margin by setting $\zeta_{m}=(1+\beta)v_{m}$ where $\beta>0$ is the flexibility margin: a higher margin encourages more users to create CDPs, but also allows for larger deviations from the optimal portfolio.
5.3 Financial Analysis: Evaluation
In this section, we analyze the risk of system failure in both single-chain and multi-chain settings using the approach described above. We first describe the data for our model, then the portfolio compositions, and finally the results of our analysis.
Description of Data. We have collected data using the CoinMarketCap API over six different periods that span a total of 1264 hours of data, separated into 5-minute slots. The duration of these periods can be found in Table 9 in Appendix F. The final period includes the collapse of the FTX exchange on November 11, 2022. In each time slot, we collected price information of 100 tokens, namely those that were the 100 largest in terms of market capitalization during the first time slot. A complete list of the included token symbols, and the observed correlation between them, can be found in Figure 6 in Appendix F.
DSS-Based Portfolios. For our analysis, we have considered single-asset portfolios consisting of Bitcoin and several of the highest-value smart contract platforms, namely: Ethereum, Cardano, Solana, Polkadot, Tron, Avalanche, Algorand, and EOS – their cryptocurrency symbols are: BTC, ETH, ADA, SOL, DOT, TRX, AVAX, ALGO, and EOS. The results are displayed in Table 1 (single-asset). We use the following multi-asset portfolios:
  1. 1.
    D-All, which approximates the collateral portfolio of the DSS as of early 2023 (see Table 8 in Appendix E). (We replace GUSD, which was not in our dataset, with another fiat-backed stablecoin (USDT), and Maker’s real-world assets with gold-backed PAXG.)
  2. 2.
    D-ERC, which approximates the DSS portfolio after removing all fiat-backed assets.
  3. 3.
    D-Mix1 and D-Mix2, which use the same tokens as D-ERC but which increase the proportion of the minor ERC-20 tokens.
  4. 4.
    C-Mix1 and C-Mix2, which replace the ERC-20 tokens in D-Mix1 and D-Max2 with tokens from other popular blockchain platforms: ADA, DOT, TRX, and SOL.
The exact compositions of these portfolios can be found in Table 10 in Appendix G.
In our experiments, the expected over-collateralization to avert system failure is $\theta=1.1$. For each table entry, we calculate the probability of failure within 1 day, i.e., 288 5-minute time slots, using 100 000 simulation runs. In each run, we simulate the evolution of token prices by randomly drawing, in each consecutive time slot, the price change for each token type in the portfolio. We then compute the risk of failure as the fraction of runs that resulted in failure among the total number of runs. The results are displayed in Table 1Appendix B includes similar tables with different simulation methods and confidence interval widths.
Table 1: Chance of failure for single asset portfolios and initial over-collateralization.
single-asset multi-asset $\gamma^{\prime}$     BTC     ETH     ADA     SOL     DOT     TRX     AVAX     ALGO     EOS     D-All     D-ERC     D-Mix1     D-Mix2     C-Mix1     C-Mix2 1.2 0.021 0.047 0.078 0.228 0.063 0.018 0.126 0.166 0.097 0.001 0.051 0.040 0.050 0.030 0.029 1.3 0.003 0.008 0.014 0.093 0.012 0.003 0.029 0.051 0.016 0.000 0.007 0.005 0.006 0.003 0.003 1.4 0.001 0.003 0.005 0.059 0.004 0.001 0.009 0.019 0.004 0.000 0.002 0.002 0.002 0.001 0.001 1.5 0.001 0.001 0.002 0.041 0.002 0.001 0.004 0.007 0.002 0.000 0.001 0.001 0.001 0.001 0.001
From Table 1, we observe a considerable difference between the failure probabilities of the single-token portfolios. The most high-risk portfolio was Solana (SOL), which was severely affected by the collapse of FTX, since it was one of FTX’s most prominent backers [11]. Even for an initial collateralization $\gamma^{\prime}=1.5$, a pure SOL portfolio would have a failure probability of more than $4\%$ after a day, given the data from our observation period. In contrast, BTC and TRX were the most stable cryptocurrencies over this period.
Figure 3: Correlation between the different token types.
Furthermore, we observe from Table 1 that the DSS’s approximate portfolio has a very low probability of failure due to $80\%$ of its collateral being stablecoins and other pegged assets. In fact, for $\gamma^{\prime}=1.2$, its observed failure probability (0.001 for D-All) was roughly 40-50 times lower than for the pure crypto-backed portfolios (0.051, 0.040, and 0.050 respectively for D-Erc, D-Mix1, and D-Mix2). For the pure crypto-backed portfolios, we find that D-ERC has slightly lower failure probabilities than a pure ETH portfolio, which can be explained by the higher share of low-volatility WBTC in the portfolio. Similarly, C-Mix1 and C-Mix2 portfolios have lower risk than the D-ERC due to the inclusion of more WBTC. We do not observe much benefit from the inclusion of more different asset types due to the high correlation between the included tokens. However, when comparing the C-Mix1 and C-Mix2 portfolios we do note that C-Mix2 has a lower failure probability despite the inclusion of less low-volatility WBTC. The reason is the inclusion of TRX, which is not only similarly low-risk as WBTC, but which also has a relatively low correlation with the other asset types – this can be seen in Figure 2(a). We also note that the relative differences become starker when $\gamma^{\prime}$ increases – e.g., for $\gamma^{\prime}=1.2$, the risk of D-Mix1 is roughly a third higher than C-Mix2 (4.0% vs. 2.9%), whereas for $\gamma^{\prime}=1.3$ the risk is two thirds higher (0.5% vs. 0.3%).
Optimized Portfolios. The portfolios in the previous section approximated the existing DSS portfolio, which is aligned with the willingness of real-world users to deposit collateral of the various types. Although we did observe a slightly lower risk probability for the best multi-chain portfolio (C-Mix2) than for the best single-chain portfolio (D-Mix1), the underlying fundamental issue is how they compare if we were to use the convex optimization method based on (4) to determine low-volatility portfolios.
A first question in understanding this is whether to include stablecoins in the portfolio. If we use convex optimization across the entire dataset of 100 token types, we obtain the portfolio from Table 12 in Appendix G. With the exception of BTT (the BitTorrent token), this portfolio only includes USD- or gold-pegged assets. A crypto-backed stablecoin system that only uses fiat-backed stablecoins as collateral has trivially low price volatility, yet would share most of the drawbacks of a fiat-backed stablecoin as discussed in Section 2 (i.e., higher vulnerability to black swan events that involve off-chain entities). As such, we will also consider a stablecoin that purely uses unpegged cryptocurrency assets as collateral. This represents a worst-case scenario in terms of volatility risk: in practice, an appropriate mixture of stablecoins and unpegged cryptocurrencies can be chosen. If we restrict the portfolios to the same tokens as the portfolios of Table 10, we obtain the optimal portfolios D-Opt and C-Opt of Table 11 (both in Appendix G). We also include the portfolio A-Opt which may contain any of the Top 100 tokens except those already pegged to a real-world asset. Since we observe that the optimal portfolios tend to restrict themselves to a handful of tokens (e.g., only WBTC and GNO/TRX for D-Opt and C-Opt), we also set a threshold of $\lambda=0.2$ for D-Opt and C-Opt and $0.1$ for A-Opt to avoid high exposure to events on a single blockchain.
Table 2: Chance of failure for given portfolio and initial over-collateralization.
D-Opt C-Opt A-Opt risk vs. D-Opt
$\gamma^{\prime}$ $\lambda=0.2$ $\lambda=1$ $\lambda=0.2$ $\lambda=1$ $\lambda=0.1$ $\lambda=0.2$ $\lambda=1$ C-Opt A-Opt
1.2 0.03419 0.01423 0.02103 0.00934 0.00895 0.00624 0.00617 -38.49% -81.75%
1.3 0.00397 0.00133 0.00232 0.00131 0.00076 0.00055 0.00081 -41.56% -86.14%
1.4 0.00115 0.00058 0.00087 0.00056 0.00030 0.00020 0.00024 -24.34% -82.61%
1.5 0.00057 0.00027 0.00037 0.00024 0.00014 0.00008 0.00012 -35.08% -85.97%
The results of our optimal portfolio simulations are displayed in Table 2. The final (bold) column displays the decrease in terms of failure risk estimates for C-Opt and A-Opt versus D-Opt for $\lambda=0.2$. In general, we find that the optimized portfolios have slightly lower failure probabilities than the DSS-based ones from the previous section (e.g., for $\gamma^{\prime}=1.1$ and $\lambda=0.2$, we obtain 0.034 for D-Opt vs. 0.040 for D-Mix1, and 0.021 for C-Opt vs. 0.029 for C-Mix2). The C-Opt portfolios again outperform D-Opt portfolios by roughly a third, confirming our expectation that a multi-chain stablecoin is less vulnerable to price fluctuations than a single-chain one. Furthermore, we find even lower failure probabilities for A-Opt, although this portfolio contains many tokens on blockchains that are not as high-use as the tokens in the other portfolios – e.g., BSV or KLAY vs. BTC and ETH in other portfolios. Less-used tokens may be more likely to experience black swan events as there are fewer incentives to scrutinize smart contract code or the identity of the consensus nodes. In any case, we find that even for pure crypto-backed portfolios (C-Opt) and sufficient over-collateralization ($\gamma^{\prime}=1.4$, which is similar to $\gamma=1.5$ in the DSS and $\theta=1.1$) CroCoDai’s estimated system failure (i.e., de-peg) risk after a day is less than one in a thousand, which may be acceptable for low-to-medium-value commerce (e.g., a ticket auction).
6 Implementation
In this section, we discuss the technical details of three prototype implementations of our stablecoin system. We compare them in terms of the on-chain processing cost (gas) and the off-chain computation time. Both are essential to system performance and thus usability: if the on-chain cost is too high then each transfer will incur large fees, whereas if the computation time is too high then there is a large delay for cross-chain transactions. We focus on the implementation of the relay nodes, because all other components depicted in Figure 2 (the coin, vault, auction, savings, and oracle contracts) can be re-used from the DSS without modification – we only add the relay contracts as wards for the coin contract. As such, we cannot provide a straightforward comparison with the DSS, as the performance of CroCoDai contracts is either exactly the same or there is no DSS equivalent. However, we will discuss the comparison with alternatives from the scientific literature, e.g., zkBridge.
Each of our three designs uses a quorum of trusted nodes, but has a different tradeoff in terms of scalability and generality. Each implementation consists of two smart contracts written in Ethereum’s Solidity language: Relay.sol and CCTransfer.sol, although only Relay.sol differs between implementations. The Relay and CCTransfer contracts interact with an ERC-20 compliant contract for token transfers, Coin.sol, which re-uses the code from the Dai.sol contract in the DSS [41]. Our experiments were run on a Lenovo IdeaPad with an Intel(R) Core(TM) i5-1035G1 CPU (1.00GHz) and 8GB RAM (although the gas costs are machine-independent). The scripts that performed off-chain computations were written in Python and used the Web3 package for most cryptographic operations, while the underlying blockchain was simulated using Truffle’s Ganache tool [21]. For communication about cross-chain transfers, the relay nodes can use a private blockchain, e.g., a private Go-Ethereum blockchain, or a bespoke cross-chain communication framework such as [46].
6.1 Smart Contracts
CCTransfer. This contract stores cross-chain transfer requests as a hashmap of Request instances, which are implemented as a Solidity struct. Each CCTransfer contract is linked to a specific Coin and Relay contract pair, and must be made a Coin contract ward to allow it to mint or burn stablecoins upon the commitment of cross-chain transfers.
Relay. In the first design, a hashmap of relay node addresses in Solidity’s address format is stored in the contract. For each cross-chain transfer, at least $12f+$ 1 relay nodes must call the Relay node’s approve function using their stored address. This design exploits the EVM’s built-in signature verification, and generalizes trivially to other blockchain platforms. However, the downside is that each relay node must have an account and spend gas on each supported blockchain for every cross-chain transfer. In the second design, a hashmap of relay node public keys is stored in the contract, and a batch of $12f+$ 1 32-byte signatures is passed in the form of a byte array as a function parameter for each transaction. These signatures are then validated one by one by repeatedly calling the built-in ecrecover function. Although this is not as generic as the previous approach (e.g., each chain with its different cryptographic primitives would require its own off-chain signature generation script) it is more efficient and does not require each individual relay node to maintain an account. In the third design, we use a $t$-out-of-$n$ threshold signature scheme (where in our case $t=2f+1$) to produce a single 32-byte signature for the relay node quorum, e.g., using the procedure of [23]. The resulting signature is a standard ECDSA signature that can be verified using ecrecover. For our implementation we have used the multi-party ECDSA library by ZenGo X [59].
Table 3: Performance comparison of the three implementation designs.
gas cost $T_{\textnormal{proc}}$ (s)
design $n=4$ $n=10$ $n=16$ $n=4$ $n=10$ $n=16$
(1) $n$/$n$ 171791 511473 977350
(2) $n$/$1$ 214373 430690 647118 0.025 0.071 0.091
(3) $1$/$1$ 53112 53117 53110 1.101 7.780 34.513
6.2 Experimental Results
Table 3 compares the different designs in terms of the gas cost and the processing time at the relay nodes, $T_{\textnormal{proc}}$. We observe that the first design is the worst in terms of performance, which is the cost of its generality and ease of implementation. The second design is more efficient in terms of gas cost for large relay node quora, and only has little processing time at the relay nodes. The third design is the most efficient in terms of gas cost, but at the cost of considerable processing time at the relay nodes, especially for large quora. For the first ($n/n$) design, we could not distinguish $T_{\textnormal{proc}}$ from the transaction processing time in Ganache as both happen in a single step in Python’s Web3 package – however, these are assumed to be similar to signing in the $n/1$ design, and are indicated with ‘—’ in the table.
By comparison, zkBridge [54] reports a gas cost of $\approx$230k gas for on-chain verification, but with a processing cost of $\approx$18 seconds to generate a proof with 32 high-end Amazon Web Services machines (c5.24xlarge). This lower performance is to be expected, as the additional complexity of zkBridge removes the need for a trusted quorum of relay nodes.
7 Conclusions and Future Work
In this work we have presented CroCoDai, a stablecoin system for cross-chain commerce. CroCoDai facilitates cross-chain transfers by design, and is more resilient than a single-chain system to black swan events and price fluctuations. Our empirical analysis shows that with sufficient overcollateralization, the probability of system failure can be made small even under a model whose price data includes the highly volatile period after the collapse of the FTX exchange. We have presented three prototype implementations that rely on a quorum of relay nodes for cross-chain transfers and governance decisions.
For future work, a first interesting direction is to implement a version of CroCoDai that performs all relay operations on-chain, as per the second design of Section 4.4. Promising directions to achieve such a design would be to use trustless bridges as proposed in zkBridge, or succinct proofs of state such as those proposed for Algorand or Mina. This would remove the need for a trusted quorum of relay nodes, while maintaining CroCoDai’s advantages over the single-chain DSS, i.e., higher resilience to black swan events and price fluctuations. A second interesting direction is to further investigate alternative collateral liquidation mechanisms – for example, Salehi et al. [48] propose to mitigate liquidations from the DSS or remove them completely. For example, collateral of the same token type in a vault could be pooled together and liquidated only when the entire pool is underwater – however, this would remove an incentive from users to overcollateralize. A third and final interesting research direction is system recovery after a 51% attack on a coin chain. As we discussed in Section 5, such an attack could create stablecoins without matching collateral on a different chain. This is similar to the situation after a failed collateral auction in the DSS, so it may be possible to resolve it similarly – i.e., with the governance token holders absorbing the losses.
Appendix A List of Symbols
A list of symbols used in this paper can be found in Table 4.
Table 4: List of symbols.
symbol description
$p_{m\,t}$ base price of token $m$ at time $t$
$p_{m\,o\,t}$ price of token $m$ at time $t$, as reported by oracle $o$
$\mu_{m}$ mean drift of the stochastic process $p_{m\,t}$
$\sigma^{2}_{m}$ variance of the stochastic process $p_{m\,t}$
$\mu_{m\,o}$ mean of the random variable $p_{m\,o\,t}-p_{m\,t}$
$\sigma^{2}_{m\,o}$ variance of the random variable $p_{m\,o\,t}-p_{m\,t}$
$y_{j}$ collateral token type of CDP $j$
$c_{j\,t}$ number of collateral tokens held in CDP $j$ at time $t$
$D_{j\,t}$ debt of CDP $j$ at time $t$
$\zeta_{m}$ debt ceiling of token type $m$
$\zeta^{\prime}$ total debt ceiling of the failed/compromised asset types
$\gamma$ liquidation ratio
$\lambda$ maximum contribution of each individual asset type during convex optimization
$\theta$ system failure threshold
$\mathcal{J}_{t}$ set of all CDPs at time $t$
$M$ number of tokens
$O$ number of price feeds
$U$ number of users
Appendix B Alternative Simulation Methods
In the following, we discuss two alternative methods for estimating failure risks through simulation: using pure historical data (without an underlying price model), and with a standard GBM as discussed in Section 3.2.
For the first method, we assume for each time slot that our portfolio had a given overcollateralization at the start of the slot and check if system failure occurs within a period of interest (e.g., one day). We repeat this process for all time steps in our dataset. The downside of this approach is that our dataset captures a period with many sharp price drops, which results in a somewhat pessimistic outlook on the failure risk – on the other hand, it captures realistically the limitations of stablecoins that rely on crypto-assets.
For the second method, we generate correlated, normally distributed samples that have the same covariance matrix as the original prices. To achieve this, let $\mathbf{L}$ be computed as $\textnormal{Cov}(\mathbf{X})=\mathbf{L}\mathbf{L}^{T}$ using Cholesky decomposition. Random samples can then be generated as $\mathbf{LW}=\mathbf{R}$, where the matrix $\mathbf{W}$ contains random values sampled from the standard normal distribution. We can then calculate the returns of the assets in portfolio as
$S_{t}=S_{0}\cdot\left\{\left(\mu-\frac{\sigma^{2}}{2}\right)+\sigma W_{t}\right\}.$ (5)
We emphasize that the method used in Section 5.2 is the same as the second method, with the exception being that we use the multivariate $t$-distribution to represent the returns instead of the multivariate normal distribution.
[…]
Table 5: Estimated portfolio failure probabilities - historical method.
single-asset multi-asset $\gamma^{\prime}$     BTC     ETH     ADA     SOL     DOT     TRX     AVAX     ALGO     EOS     D-All     D-ERC     D-Mix1     D-Mix2     C-Mix1     C-Mix2 1.2 0.052 $\pm$0.0057 0.072 $\pm$0.0067 0.060 $\pm$0.0062 0.198 $\pm$0.0103 0.048 $\pm$0.0056 0.040 $\pm$0.0051 0.134 $\pm$0.0088 0.135 $\pm$0.0089 0.089 $\pm$0.0074 $0$$\pm$0.0000 0.067 $\pm$0.0065 0.060 $\pm$0.0062 0.060 $\pm$0.0062 0.058 $\pm$0.0061 0.052 $\pm$0.0058 1.3 0.002 $\pm$0.0012 0.024 $\pm$0.0040 0.003 $\pm$0.0013 0.054 $\pm$0.0059 0.006 $\pm$0.0020 0.002 $\pm$0.0012 0.020 $\pm$0.0036 0.037 $\pm$0.0049 0.022 $\pm$0.0038 $0$$\pm$0.0000 0.024 $\pm$0.0039 0.022 $\pm$0.0038 0.024 $\pm$0.0039 0.019 $\pm$0.0035 0.006 $\pm$0.0020 1.4 $0$$\pm$0.0000 0.004 $\pm$0.0016 $0$$\pm$0.0000 0.037 $\pm$0.0049 $0$$\pm$0.0000 $0$$\pm$0.0000 0.003 $\pm$0.0015 0.019 $\pm$0.0035 $0$$\pm$0.0000 $0$$\pm$0.0000 0.001 $\pm$0.0010 0.001 $\pm$0.0010 0.001 $\pm$0.0010 $0$$\pm$0.0000 $0$$\pm$0.0000 1.5 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 0.030 $\pm$0.0044 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 0.003 $\pm$0.0015 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000
Table 6: Estimated portfolio failure probabilities - standard GBM.
single-asset multi-asset $\gamma^{\prime}$     BTC     ETH     ADA     SOL     DOT     TRX     AVAX     ALGO     EOS     D-All     D-ERC     D-Mix1     D-Mix2     C-Mix1     C-Mix2 1.2 0.024 $\pm$0.0005 0.046 $\pm$0.0007 0.043 $\pm$0.0006 0.147 $\pm$0.0011 0.036 $\pm$0.0006 0.024 $\pm$0.0005 0.073 $\pm$0.0008 0.124 $\pm$0.0010 0.069 $\pm$0.0008 $0$$\pm$0.0000 0.045 $\pm$0.0007 0.037 $\pm$0.0006 0.042 $\pm$0.0006 0.034 $\pm$0.0006 0.031 $\pm$0.0005 1.3 0.001 $\pm$0.0001 0.005 $\pm$0.0002 0.003 $\pm$0.0002 0.070 $\pm$0.0008 0.002 $\pm$0.0001 0.001 $\pm$0.0001 0.010 $\pm$0.0003 0.025 $\pm$0.0005 0.007 $\pm$0.0003 $0$$\pm$0.0000 0.004 $\pm$0.0002 0.002 $\pm$0.0001 0.003 $\pm$0.0002 0.002 $\pm$0.0001 0.001 $\pm$0.0001 1.4 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 0.038 $\pm$0.0006 $0$$\pm$0.0000 $0$$\pm$0.0000 0.001 $\pm$0.0001 0.004 $\pm$0.0002 $0$$\pm$0.0001 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 1.5 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 0.019 $\pm$0.0004 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0001 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000
Table 7: Estimated portfolio failure probabilities - Student’s $t$ GBM.
single-asset multi-asset $\gamma^{\prime}$     BTC     ETH     ADA     SOL     DOT     TRX     AVAX     ALGO     EOS     D-All     D-ERC     D-Mix1     D-Mix2     C-Mix1     C-Mix2 1.2 0.022 $\pm$0.0005 0.047 $\pm$0.0007 0.077 $\pm$0.0008 0.227 $\pm$0.0013 0.062 $\pm$0.0008 0.019 $\pm$0.0004 0.126 $\pm$0.0010 0.164 $\pm$0.0012 0.099 $\pm$0.0009 0.001 $\pm$0.0001 0.050 $\pm$0.0007 0.040 $\pm$0.0006 0.048 $\pm$0.0007 0.028 $\pm$0.0005 0.029 $\pm$0.0005 1.3 0.003 $\pm$0.0002 0.008 $\pm$0.0003 0.014 $\pm$0.0004 0.092 $\pm$0.0009 0.012 $\pm$0.0003 0.003 $\pm$0.0002 0.029 $\pm$0.0005 0.051 $\pm$0.0007 0.015 $\pm$0.0004 $0$$\pm$0.0000 0.007 $\pm$0.0003 0.005 $\pm$0.0002 0.007 $\pm$0.0003 0.003 $\pm$0.0002 0.003 $\pm$0.0002 1.4 0.001 $\pm$0.0001 0.003 $\pm$0.0002 0.005 $\pm$0.0002 0.059 $\pm$0.0007 0.004 $\pm$0.0002 0.001 $\pm$0.0001 0.009 $\pm$0.0003 0.019 $\pm$0.0004 0.005 $\pm$0.0002 $0$$\pm$0.0000 0.002 $\pm$0.0001 0.001 $\pm$0.0001 0.002 $\pm$0.0001 0.001 $\pm$0.0001 0.001 $\pm$0.0001 1.5 0.001 $\pm$0.0001 0.002 $\pm$0.0001 0.002 $\pm$0.0001 0.042 $\pm$0.0006 0.002 $\pm$0.0001 0.001 $\pm$0.0001 0.004 $\pm$0.0002 0.008 $\pm$0.0003 0.002 $\pm$0.0002 $0$$\pm$0.0000 0.001 $\pm$0.0001 0.001 $\pm$0.0001 0.001 $\pm$0.0001 0.001 $\pm$0.0001 0.001 $\pm$0.0001
We visualize the difference between the standard and $t$-distribution GBM for 10 separate simulation runs in Figure 4(a) and Figure 4(b), respectively. The $t$-distribution GBM runs are visibly more likely to experience sharp jumps than the standard GBM runs.
Figure 4: For 10 separate asset price simulation runs, the total value of the portfolio 50% BTC and 50% ETH) over a 1 day period, using the standard GBM (a) and the GBM with the $t$-distribution (b). The dashed line represents the system failure threshold for $\gamma$ = 1.2 and $\theta$ = 1.1.
Appendix C Distribution Fitting Examples
The GBM is usually assumed to follow a normal distribution. In previous simulations we used normal distributions to simulate the returns. As mentioned by many papers, distribution of stock returns is usually more heavy-tailed and Student’s t-distribution is better fit (see, e.g., [34, 29]). When comparing residual sum of squares for normal and t-distribution on returns of ETH, we get a residual sum of squares (RSS) that is approximately equal to 31900 and 14000 for the normal and $t$- distribution, respectively. Moreover, the fitted $t$-distribution has a lower RSS for all other assets as well. In Figure 5, we have displayed a comparison of the normal and $t$-distribution fitted on Ethereum and Solona returns. As such, we prefer to use the $t$-distribution for the risk analysis in this paper.
Figure 5: Comparison between t-distribution and normal distribution fitted on historical returns.
Appendix D Proofs of Theorems 5.1 – 5.3
Before we prove the main theorems, we need the following technical lemmas. We need the first to prove that if all CDPs are over-collateralized, then so is system as a whole.
Lemma D.1.
If $a_{i}/b_{i}>c$ for $i=1,\ldots,n$, then $\frac{\sum_{i=1}^{n}a_{i}}{\sum_{i=1}^{n}b_{i}}>c$.
Proof D.2.
Note that $\frac{\sum_{i=1}^{n}a_{i}}{\sum_{i=1}^{n}b_{i}}=\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}\frac{b_{i}}{\sum_{j=1}^{n}b_{j}}$. Since $\sum_{i=1}^{n}\frac{b_{i}}{\sum_{j=1}^{n}b_{j}}=1$, $\frac{\sum_{i=1}^{n}a_{i}}{\sum_{i=1}^{n}b_{i}}$ is a convex combination of $a_{i}/b_{i},\ldots,a_{n}/b_{n}$. As such, it cannot be smaller than $\min_{i=1,\ldots,n}a_{i}/b_{i}$, which is greater than $c$ by the statement of the lemma.
The second lemma is a commonly-known result, but we include the proof here for completeness.
Lemma D.3.
If $X$ has a normal distribution with mean 00 and variance $1$, then
$R(x)=\mathbb{P}(X>x)\leq\frac{e^{-x^{2}/2}}{x\sqrt{2\pi}}.$
Proof D.4.
Since for $y\geq x\geq 0$ we have that $1\leq\frac{y}{x}$, we have that
[…]
With these lemmas, we are ready to prove the three theorems from Section 5.
Proof D.5 (Proof of Theorem 5.1).
If $p_{y_{j}\,t}c_{j\,t}/y_{j\,t}>\gamma$ holds for all $j\in\mathcal{J}$, then by Lemma D.1 it must hold that
$\frac{C^{+}_{0}}{D^{+}_{0}}>\gamma.$ (6)
Furthermore, we know from the definition of the debt ceiling that $D^{-}_{0}<\zeta^{\prime}(D^{-}_{0}+D^{+}_{0})$, where $\zeta^{\prime}$ (as defined in Section 5) is the total debt ceiling of the affected tokens. The above can be rewritten to
$D^{-}_{0}<\zeta^{\prime}/(1-\zeta^{\prime})D^{+}_{0}.$ (7)
We therefore find that
$\frac{C^{*}_{1}}{D^{*}_{1}}=\frac{C^{+}_{0}}{D^{-}_{0}+D^{+}_{0}}>\frac{C^{+}_{0}}{\left(1+\frac{\zeta^{\prime}}{1-\zeta^{\prime}}\right)D^{+}_{0}}>\gamma(1-\zeta^{\prime})$
where the first inequality holds because of (7) and the second because of (LABEL:eq:unaffected_over-collateralized).
Proof D.6 (Proof of Theorem 5.2).
Because of the debt ceilings on the tokens, the total debt incurred for each token type $m$ must still satisfy $\sum_{\begin{subarray}{c}j\in\mathcal{J}_{t}\\ y_{j}=m\end{subarray}}D_{j\,t}<\zeta_{m}\sum_{j\in\mathcal{J}_{t}}D_{j\,t}.$ – as such, even after the attack, we must have $D^{-}_{1}=D^{-}_{0}+h^{\prime}<\zeta^{\prime}(D^{+}_{0}+D^{-}_{0}+h^{\prime})$ or
$D^{-}_{0}+h^{\prime}<\zeta^{\prime}/(1-\zeta^{\prime})D^{+}_{0}$ (8)
where $D^{-}_{0}$, $D^{+}_{0}$, and $\zeta^{-}$ are as defined in the previous section. We have that
$\frac{C^{*}_{1}}{D^{*}_{1}}=\frac{C^{-}_{0}+C^{+}_{0}}{D^{-}_{0}+D^{+}_{0}+h}>\frac{C^{-}_{0}+C^{+}_{0}}{\left(1+\frac{\zeta^{\prime}}{1-\zeta^{\prime}}\right)D^{+}_{0}}>\gamma(1-\zeta^{\prime})$
where the first inequality holds because of (8) and the second because $C^{-}_{0}>0$ and because of (6). As such, we again find that the debt ceilings ensure that CroCoDai remains fully backed if it is sufficiently over-collateralized.
Proof D.7 (Proof of Theorem 5.3).
In the following, we first assume that $O$ is odd and that some subset $O^{\prime}\subset[O]$ is controlled by malicious parties who aim to convince a vault in CroCoDai that the price is at most than some $p^{*}>p$. In this case, the median equals the $\frac{1}{2}(M+1)+|O^{\prime}|$ th order statistic from a sample of $M-|O|$ normally distributed random variables with mean $p$ and standard deviation $\sigma_{o}^{2}$. This order statistic is greater than the maximum of the $M-O^{\prime}$ random variables, which we denote by $Y$. Let $\sigma_{*}=\min_{o\in[O]}\sigma_{o}$. $Y$’s cumulative distribution function equals
$\begin{split}\mathbb{P}(Y\geq p^{*})&=\mathbb{P}\left(\bigwedge_{o\in[O]\backslash M^{\prime}}\Delta_{o}>p^{*}-p\right)=\prod_{o\notin O^{\prime}}R\left(\frac{p^{*}-p}{\sigma_{o}}\right)\\ &\leq R^{M-|O|}\left(\frac{p^{*}-p}{\sigma_{*}}\right)\leq\frac{e^{-(M-|O^{\prime}|)(p^{*}-p)^{2}/2}}{(p^{*}-p)\sigma\sqrt{2\pi}},\end{split}$
where the first inequality holds because $R$ is a monotonically decreasing function, and the second inequality follows from Lemma D.3. As such, the probability $\mathbb{P}(Y\geq p^{*})$ vanishes exponentially as $p^{*}-p$ increases. The proof for $p^{*}<p$ is very similar and uses the symmetry of the normal distribution.
Appendix E DSS Parametrization
The parameters of the various different vault types in the DSS can be found here. A sample of parameter values as of 25 April 2023 (Ethereum block 17121982) can be found in Table 8. For each of the supported token type $j$, we indicate the token name, total value of the collateral of CDPs of this type $\sum_{j\in\mathcal{J}}p_{j\,t}c_{j\,t}$, total debt of the collateral of CDPs of this type $\sum_{j\in\mathcal{J}}D_{j\,t}$, collateralization ratio, and debt ceiling $\zeta_{j}$. Although the exact values vary throughout the day, this table is meant to illustrate the typical composition of the collateral type in the DSS.
The following token types were supported in the DSS, but were later deprecated: AAVE, BAL, BAT, COMP, KNC, LRC, MANA, RENBTC, TUSD, UNI, USDT, and ZRX. The following liquidity pool tokens were deprecated: AAVE/ETH, LINK/ETH, UNI/ETH, wBTC/Dai, and wBTC/ETH.
Table 8: Information about the various collateral types used by the DSS (as of Ethereum block 17121982).
Name Token/Asset Platform debt coll. value overcoll. debt ceil. stab. fee liq. rate contrib.
Standard ERC-20 Tokens
ETH-A Ether 233580929 421276897 180% 15000000000 1.50% 145% 4.69%
ETH-B 45480335 67797760 149% 250000000 3.00% 130% 0.91%
ETH-C 306467435 407905059 133% 2000000000 0.75% 170% 6.16%
WSTETH-A Wrapped Ether 105440104 123821333 117% 500000000 1.50% 160% 2.12%
WSTETH-B 198020222 212522461 107% 500000000 0.75% 185% 3.98%
RETH-A 3503590 6503590 186% 20000000 0.50% 170% 0.07%
WBTC-A Wrapped Bitcoin 35684532 55869143 157% 500000000 1.75% 145% 0.72%
WBTC-B 22535846 32535847 144% 250000000 3.25% 130% 0.45%
WBTC-C 40106301 63565858 158% 500000000 1.00% 175% 0.81%
LINK-A ChainLink 711324 3708567 521% 5000000 2.50% 165% 0.01%
YFI-A Yearn 148553 1941526 1307% 4000000 1.50% 165% 0.00%
MATIC-A Polygon 194080 5193556 2676% 15000000 3.00% 175% 0.00%
GNO-A Gnosis 3020632 5000000 166% 5000000 2.50% 350% 0.06%
  Total (ERC-20 tokens): 994893882 1407641598 20.00%
LP Tokens
UNIV2USDCETH-A USDC/ETH 2139902 0 0% 1.50% 120% 0.04%
UNIV2DAIUSDC-A Dai/USDC 23983390 0 0% 0.02% 102% 0.48%
GUNIV3DAIUSDC1-A Dai/USDC 59841202 0 0% 0.02% 102% 1.20%
GUNIV3DAIUSDC2-A Dai/USDC 104141233 0 0% 0.06% 102% 2.09%
CRVV1ETHSTETH-A ETH/wETH 56643346 66542992 117% 100000000 1.50% 155% 1.14%
  Total (Liquidity Pool ERC-20 tokens): 246749074 66542992 4.96%
Real-World Assets
RWA001-A 6s Capital 14348036 15000000 105% 3.00% 0.29%
RWA002-A Centrifuge 6429638 20000000 311% 3.50% 0.13%
RWA003-A 1824487 2000000 110% 6.00% 0.04%
RWA004-A 1468187 7000000 477% 7.00% 0.03%
RWA005-A 6076346 15000000 247% 4.50% 0.12%
RWA006-A AEA 0 0 2.00% 0.00%
RWA007-A Monetalis 499950318 1250000000 250% 0.00% 10.05%
RWA008-A SG Forge 0 30000000 0.05% 0.00%
RWA009-A H. V. Bank 100000000 100000000 100% 0.00% 2.01%
RWA010-A Centrifuge 0 20000000 4.00% 0.00%
RWA011-A 0 30000000 4.00% 0.00%
RWA012-A 13619456 30000000 220% 4.00% 0.27%
RWA013-A 69525268 70000000 101% 4.00% 1.40%
  Total (Real-World Assets): 713241736 1589000000 14.34%
PSMs
PSM-USDC-A USD Coin 2118503205 2608700392 123% 10000000000 42.58%
PSM-USDP-A Pax Dollar 499997224 500000000 100% 500000000 10.05%
PSM-GUSD-A Gemini Dollar 387538598 437538598 113% 500000000 7.79%
  Total (Peg Stability Modules): 3006039027 3546238990 60.42%
Appendix F Price Data Table
Table 9 contains a specification of the time periods duirng which we collected price data for the empirical analysis of Section 5. Figure 6 visualizes the correlations between the 100 tokens in our dataset.
Table 9: Data periods (times are in Greenwich Mean Time).
start end number of slots
Period 1 26-08-2022 02:16 30-08-2022 00:26 1131
Period 2 07-09-2022 03:15 15-09-2022 08:45 2371
Period 3 30-09-2022 07:37 02-10-2022 05:27 551
Period 4 07-10-2022 04:31 15-10-2023 22:46 2010
Period 5 17-10-2022 02:51 29-10-2022 16:26 3620
Period 6 04-11-2022 03:50 23-11-2022 05:15 5490
Figure 6: The correlations between all 100 tokens in our dataset.
Appendix G Portfolio Compositions
Table 10 displays the composition of the DSS-based simulation portfolios. Table 11 displays the composition of the optimized portfolios. Table 12 displays the optimal portfolio can be constructed using the 100 tokens in our dataset, if other stablecoins can be included.
Table 10: DSS-based portfolio compositions.
portfolio name token distribution
D-All 47% USDC, 19% ETH, 14% PAXG
10% USDP, 8% USDT, 2% WBTC
D-ERC 90.2% ETH, 9.4% WBTC, 0.2% GNO,
0.1% LINK, 0.05% MATIC, 0.05% YFI
D-Mix1 50% ETH, 30% WBTC, 5% GNO,
5% LINK, 5% MATIC, 5% YFI
D-Mix2 20% ETH, 20% WBTC, 15% GNO,
15% LINK, 15% MATIC, 15% YFI
C-Mix1 50% ETH, 30% WBTC, 5% ADA,
5% DOT, 5% TRX, 5% AVAX
C-Mix2 20% ETH, 20% WBTC, 15% ADA,
15% DOT, 15% TRX, 15% AVAX
Table 11: Optimal portfolio compositions.
name optimal token distribution
D-Opt WBTC 74.015%, GNO 25.985%
D-Opt WBTC 20%, ETH 20%
($\lambda=0.2$) GNO 20%, LINK 20%, YFI 20%
C-Opt WBTC 62.204%, TRX 37.796%
C-Opt WBTC 20%, ETH 16.656%, ADA 14.493%
($\lambda=0.2$) DOT 20%, TRX 20%, EOS 8.851%
A-Opt BSV 8.728%, BTG 0.796%, BTT 36.821%,
DCR 4.357%, GNO 5.221%, HT 0.0762%
KCS 2.726%, KLAY 5.705%, LEO 17.568%
OKB 3.223%, TRX 13.017%, TWT 0.483%, XLM 0.593%
BSV 10%, BTG 1.354%, BTT 10%,
A-Opt CAKE 0.644%, DCR 7.278%, GNO 10%,
($\lambda=0.1$) HT 1.742%, KCS 5.819%, KLAY 9.434%,
LEO 10%, NEXO 0.311%, OKB 5.604%,
TRX 10%, TWT 8.61%, WBTC 10%, XLM 6.951%
Table 12: Optimal simulation portfolio if other stablecoins are included.
BTT 0.34% BUSD 0.1504% PAXG 8.72%
USDC 35.90% USDN 0.12% USDT 39.89%

### v4

@@
 CroCoDai: A Stablecoin for Cross-Chain Commerce
-Abstract
+Abstract.
 Decentralized Finance (DeFi), in which digital assets are exchanged without trusted intermediaries, has grown rapidly in value in recent years.
-The global DeFi ecosystem is fragmented into multiple blockchains, fueling the demand for cross-chain commerce.
-Existing approaches for cross-chain transactions, e.g., bridges and cross-chain deals, achieve atomicity by locking assets in escrow.
-However, locking up assets increases the financial risks for the participants, especially due to price fluctuations and the long latency of cross-chain transactions.
-Stablecoins, which are pegged to a non-volatile asset such as the US dollar, help mitigate the risk associated with price fluctuations.
-However, existing stablecoin designs are tied to individual blockchain platforms, and trusted parties or complex protocols are needed to exchange stablecoin tokens between blockchains.
-Our goal is to design a practical stablecoin for cross-chain commerce.
-Realizing this goal requires addressing two challenges.
-The first challenge is to support a large and growing number of blockchains efficiently.
-The second challenge is to be resilient to price fluctuations and blockchain platform failures.
+Stablecoins, which are pegged to a non-volatile asset such as the US dollar, are a prominent feature of DeFi as they mitigate the risk associated with price fluctuations.
+However, existing stablecoin systems are tied to individual blockchain platforms, and trusted parties or complex protocols are needed to exchange stablecoin tokens between blockchains.
+Our goal is to design a practical stablecoin system for cross-chain commerce, and to do so we must overcome two main challenges.
+The first is to support a large and growing number of blockchains efficiently.
+The second is for the stablecoin to be resilient to blockchain platform failures and to price fluctuations that affect its collateral.
 We present CroCoDai to address these challenges.
-The key insight is to extend the existing Dai stablecoin system to a stablecoin that is backed with collateral tokens from multiple blockchains.
-CroCoDai leverages the system’s governance layer (which could be another blockchain) to exchange tokens between different blockchains – this enables secure cross-chain commerce and addresses the first challenge.
-Furthermore, the relatively low correlation between the price of assets on different blockchains makes CroCoDai more resilient to financial shocks than a single-chain stablecoin.
-We demonstrate this advantage through an empirical analysis that uses real token price data from a period that includes the collapse of the FTX exchange, which caused high token price volatility.
-We present three prototype implementations of our stablecoin system, and show that it incurs small execution overhead.
-keywords:
-Stablecoins, blockchains, interoperability, financial risk analysis
-category:
+We demonstrate CroCoDai’s efficiency by comparing the performance of a prototype implementation to related baselines, and its resilience through an empirical analysis of historical token price data.
+1. Introduction
+Decentralized Finance (DeFi) promises to revolutionize finance by removing the need for trusted middlemen to process transactions.
+In DeFi, transactions commonly involve a change of ownership of digital tokens that exist on a distributed ledger such as a blockchain [^1] In the remainder of this paper, we will use the terms ‘blockchain’ or ‘chain’ to refer to any type of distributed ledger, including ledgers that are technically not chains (e.g., Avalanche’s directed acyclic graphs (rocket2018snowflake, )).
+– e.g., one user sending bitcoins to another (btc_origin, ).
+Many blockchain platforms (e.g., Ethereum (buterin2014next, )) also support smart contracts, which enable automated versions of financial concepts such as auctions, loans, swaps, decentralized exchanges, and ownership shares.
+The total value of the tokens held in DeFi smart contracts peaked at over \$150 billion (USD) in late 2021 (werner2021sok, ; defillama, ), with at least \$100 billion held in decentralized exchanges alone (xu2023sok, ), however it had fallen to around \$50 billion as of December 2023 (defillama, ).
+(a) Fully Siloed
+(b) Wrapped + DAI
+(c) CroCoDai
+Figure 1. Overview of different cross-chain stablecoin systems. V, C, and G respectively indicate the presence of vault, stablecoin, and governance contracts.
+The DeFi ecosystem is currently fragmented over a multitude of different blockchain platforms which each offer their own choices in terms of efficiency, security, functionality, and structure of DeFi contracts.
+For example, Cardano (hoskinson2017cardano, ; kiayias2017ouroboros, ), Solana (yakovenko2018solana, ), Avalanche (rocket2019scalable, ), and Algorand (gilad2017algorand, ) each had a valuation of more than a billion USD in early 2023.
+These blockchains are not designed to interoperate: i.e., there are no native mechanisms for transferring tokens between chains.
+However, the emergence of blockchain bridges (bridges, ; celer, ) and dedicated interoperability platforms such as Cosmos (Polkadot20, ) and Polkadot (cosmos20, ) demonstrates an increasing demand for cross-chain interoperability.
+A major challenge in cross-chain trades is atomicity: the transactions in a trade must be committed on all blockchains, or not at all.
+Existing protocols for cross-chain atomicity, including relays and hashed time locks (herlihy2021cross, ; belchior2021survey, ; thyagarajan2022universal, ), essentially lock the assets involved in a trade and only release (or return) them once the transactions have been committed (or aborted) on all blockchains.
+While ensuring atomicity, this approach increases the financial risks for the users in the following ways: (1) the tokens held in escrow may lose value due to price fluctuations, (2) one party may abort the transaction, leading to an opportunity cost for the other parties, and (3) the DeFi platforms that facilitate the trade may be faulty or malicious .
+These risks are interrelated: if there is a significant change in the value of the tokens in escrow (1), then this may cause a party to abandon a trade (2), creating an opportunity cost to the other users.
+Price volatility can be mitigated in principle if the tokens held in escrow are stablecoins (klages2020stablecoins, ; cao2021designing, ; moin2020sok, ) whose value is pegged to a less volatile real-world asset (e.g., the US dollar).
+However, a flawed stablecoin design can itself put users’ tokens at risk, as witnessed by the collapse of the Terra/Luna stablecoin in May 2022 (briola2023anatomy, ), in which stablecoins worth \$currency-dollar\\$\$60 billion lost nearly all of their value.
+Our goal is to design a practical stablecoin as a central point of exchange for cross-chain commerce.
+We focus on stablecoin designs that are backed with crypto collateral – e.g., Dai (daiwhitepaper, ) on the Ethereum blockchain – to minimize the dependence on trusted parties outside the DeFi ecosystem (e.g., banks and asset management companies).
+To realize our goal, we must address two main challenges.
+The first challenge is to support a large and growing number of blockchains efficiently, including new and relatively low-use platforms.
+These low-use blockchains typically have lower transaction fees, but also less liquidity and existing DeFi support.
+The second challenge is to ensure that the stablecoin is resilient to black swan events (e.g., 51% attacks or critical smart contract bugs) that affect individual blockchain platforms, and to price fluctuations that affect multiple blockchains simultaneously.
+If the total value of the collateral drops below the intended value of the issued stablecoins – which we will refer to as system failure – then this may trigger the DeFi equivalent of a bank run, forcing stablecoin holders to sell their tokens at a loss.
+Our key insight is that a wide range of collateral can address both challenges.
+In particular, it allows users from smaller chains to create robust stablecoins without interacting with a more expensive blockchain or relying on a trusted third party.
+Furthermore, it reduces the exposure of the stablecoin system as a whole to sudden crashes and price fluctuations.
+In this work, we present CroCoDai, a stablecoin for cross-chain commerce that addresses the above challenges.
+Compared to its closest baseline approaches, CroCoDai has greater efficiency, i.e., lower transaction costs, and better or similar resilience to black swan events and price fluctuations.
+A first baseline is for each blockchain to support its own crypto-backed stablecoin and required components: vaults to store collateral (V), coin contracts to process stablecoin transfers (C), and governance contracts for system-level decisions (G) – this is depicted in Figure 1a.
+In this baseline, stablecoins can be transferred between chains through bridges that are unique to each chain pair.
+Such bridges either lock tokens as collateral on one chain and create wrapped (wbtc, ) tokens on the other, or burn tokens (karantias2020proof, ) on one chain and mint them on the other.
+[^2] The Circle-created USDC bridge between Ethereum and Avalanche (usdcbridge, ) is an example of the latter.
+However, the lack of liquidity on smaller blockchains makes it difficult to find collateral locally, and each individual stablecoin is highly vulnerable to black swan events and price fluctuations on its platform.
+A second baseline is to combine an existing crypto-backed stablecoin on a large (i.e., high-use) blockchain, e.g., Ethereum’s Dai, with two bridges for each smaller chain: one to deposit wrapped versions of the smaller chains’ native tokens on the large chain for collateral, and another to create wrapped versions of the stablecoin on the smaller chains.
+In this setting, each smaller chain has a vault contract (V) for collateral and an independent coin contract (C) for the wrapped stablecoins – see Figure 1b.
+However, this approach requires unnecessary transactions on the large blockchain to create stablecoins or move them between chains – such transactions are typically more expensive than on the smaller chains.
+By contrast, CroCoDai uses integrated vault and coin contracts (V and C) on each smaller chain, and leverages the governance layer on a dedicated relay chain for cross-chain transactions and system-level decisions, as depicted in Figure 1c.
+In this paper, we make the following contributions:
+  1. (1)
+    We present CroCoDai, a stablecoin for cross-chain commerce (Section 3), and explore two main design variants: one uses a quorum of trusted nodes for the relay chain, and the other uses an existing blockchain (Section 3.3).
+While CroCoDai re-uses parts of the existing Dai stablecoin system, interactions with vaults are strictly local to the smaller chains, and it supports cross-chain price oracles to improve price discovery on smaller blockchains.
+  2. (2)
 […]
-1 Introduction
-Decentralized Finance (DeFi) promises to revolutionize finance by removing the need for trusted middlemen to process transactions.
-In DeFi, transactions commonly involve a change of ownership of digital tokens that exist on a distributed ledger such as a blockchain – e.g., one user sending bitcoins to another [45].
-Many blockchain platforms (e.g., Ethereum [8]) also support smart contracts, which enable automated versions of financial concepts such as auctions, loans, swaps, decentralized exchanges, and ownership shares.
-The total value of the tokens held in DeFi smart contracts peaked at over \$150 billion USD in late 2021 [53, 15], with at least \$100 billion held in decentralized exchanges alone [55], and has since fallen to around \$45-50 billion as of June 2023 [15].
-The DeFi ecosystem is currently fragmented over a multitude of blockchains which each offer their own choices in terms of efficiency, security, and functionality.
-For example, Cardano [28, 31], Solana [57], Avalanche [49], and Algorand [24] each had a valuation of more than a billion USD in early 2023.
-These blockchains are not designed to interoperate: i.e., there is no native mechanisms for transferring tokens between chains.
-However, the recent emergence of blockchain bridges [19] and dedicated interoperability platforms such as Cosmos [7] and Polkadot [33] demonstrates an increasing demand for cross-chain interoperability.
-A major challenge in cross-chain trades is atomicity: the transactions in a trade must be committed on all blockchains, or not at all.
-For example, in a cross-chain auction, the first party to transfer tokens (e.g., the auctioneer) cannot be sure that the other party will uphold their part of the trade in the absence of a cross-chain protocol.
-This is of particular concern in DeFi, as individual blockchain transactions are irreversible by design.
-Existing solutions that provide cross-chain atomicity, including relays and hashed time locks [27, 2, 50], essentially lock the assets involved in a trade and only release (or return) them once the transactions in all of the blockchains have been committed (or aborted).
-While ensuring atomicity, this approach increases the financial risks for the users in the following ways: 1.
-the tokens held in escrow may lose value due to price fluctuations, 2.
-one party may abort the transaction, leading to an opportunity cost for the other parties, and 3.
-the DeFi platforms that facilitate the trade may be faulty or malicious .
-These risks are interrelated: if there is a significant change in the value of the tokens in escrow (1.), then this may cause a party to abandon a trade (2.), creating an opportunity cost to the other users.
-The risks identified above are a limitation of all cross-chain protocols that hold tokens in escrow – the longer the tokens are held, the longer they are exposed to DeFi’s notorious price volatility [1] and vulnerability to attacks [60].
-Although the abandonment of trades can be disincentivized through the use of premiums [56], price fluctuations can be mitigated in principle if the tokens held in escrow are stablecoins [32, 9, 44] whose value is pegged to a less volatile real-world asset (e.g., the US dollar).
-However, a flawed stablecoin design can itself put users’ tokens at risk, as witnessed by the collapse of the Terra/Luna stablecoin in May 2022 [6], in which stablecoins worth 60 billion USD lost nearly all of their value.
-Our goal is to design a practical stablecoin as a central point of exchange for cross-chain commerce.
-Realizing this goal requires that two main challenges are addressed.
-The first challenge is to support a large and growing number of blockchains efficiently, including new and relatively low-use platforms.
-For established platforms, Automated Market Makers (AMMs, [17]) such as Uniswap can be used to exchange tokens for stablecoins within minutes, and the stablecoins can be transferred between blockchains using a wallet on a cryptocurrency exchange such as Coinbase or Binance.
-However, the latter step may take several hours before the withdrawal on the receiving chain is processed (e.g., 2-3 hours on Crypto.com [12]), is subject to the exchange’s transaction fees, and the exchange is a single point of failure.
-Furthermore, AMMs depend on depositors to deposit collateral for each supported pair of cryptocurrencies, and this may not be available for smaller platforms – similarly, centralized exchanges may not always support transfers on smaller platforms.
-Just as an AMM requires depositors willing to create liquidity pairs, Xclaim [58] requires the availability of a custodian who is willing to put up collateral on the receiving chain, which can be a challenge for small blockchains.
-Finally, secure bridges such as zkBridge [54] either require a direct bridge to transfer stablecoins between two small blockchains, which may again be unavailable, or require that they are sent via a larger blockchain in a two-step process, which incurs additional time and transaction fee costs.
-The second challenge in designing a cross-chain stablecoin is to ensure that it is resilient to black swan events (e.g., 51% attacks) that affect individual blockchain platforms, and to price fluctuations that affect multiple blockchains simultaneously.
-This is particularly relevant if the stablecoin’s peg is maintained through tokens held as collateral on different blockchains.
-If the total value of the collateral drops below the intended value of the issued stablecoins, then the stablecoin may lose its peg, forcing holders to sell their tokens at a loss – we will refer to such an event as system failure.
-In this work, we present CroCoDai, a stablecoin for cross-chain commerce that addresses the above challenges.
-The key insight is to back CroCoDai with collateral tokens from a multitude of blockchains, using the same principles as the existing Dai stablecoin system (DSS) [38] but without being restricted to the Ethereum ecosystem.
-To that end, CroCoDai manages communication between the different blockchains through a governance layer which can be implemented through a quorum of trusted nodes, or on top of an existing blockchain.
-By using the governance layer to create/destroy the same amount of stablecoins on different chains, CroCoDai supports cross-chain transfers by design.
-Crucially, CroCoDai allows tokens on smaller blockchains to be backed by collateral on larger chains, thus providing stablecoin support to emerging platforms and addressing the first challenge.
-Unlike Polkadot or Cosmos, CroCoDai achieves its goal without requiring intrusive changes to existing blockchains.
-To demonstrate how well CroCoDai addresses the second challenge, we present a technical analysis that shows that attacks on a limited number of blockchain platforms can be tolerated as long as the system is sufficiently overcollateralized.
-We also show that attacks against a limited number of price oracles (which determine the value of the collateral) can be tolerated – this is not a challenge specific to multi-chain settings, so this security property holds for the single-chain DSS as well.
-Next, we provide evidence for the greater resilience of a multi-chain stablecoin to price fluctuations through an empirical analysis based on the collected data from several periods, including the collapse of the FTX exchange which led to severe depreciation of the Solana cryptocurrency.
-Specifically, we compare (1) a portfolio that approximates the current state of the DSS’ collateral to a similar portfolio with tokens from multiple chains, and (2) a portfolio that is optimized for low failure risk using the tokens supported by the DSS to one with tokens from multiple chains.
-Finally, we provide a reference implementation for CroCoDai that uses a dedicated relay chain, and use experiments to show that the computational overhead of CroCoDai is small.
-The code/data used in this work have been temporarily redacted to preserve anonymity, but they will be open-sourced at a later date to facilitate reproducibility and reusability.
-Contributions
-Our work makes the following contributions:
-  • We present CroCoDai, a cross-chain stablecoin – we present two design variants: one uses a quorum of nodes as a trust anchor; the other uses an existing blockchain (Section 4).
-  • We demonstrate CroCoDai’s security through a technical analysis that demonstrates that CroCoDai can tolerate a limited number of compromised blockchains or price oracles (Section 5.1), and an empirical analysis in which we compare single-chain portfolios to portfolios that use collateral from multiple chains, and find that the latter has lower failure probabilities.
-Specifically, the failure probability is roughly a third lower for optimized pure-crypto portfolios (Sections 5.2 and 5.3).
-  • We present three implementations of CroCoDai subvariants that use the quorum-based approach and compare their gas costs and processing times, to demonstrate CroCoDai’s practicality and to make its generality/performance trade-offs explicit (Section 6).
+  3. (3)
+[…]
+In summary, CroCoDai is cheaper than a baseline with wrapped tokens, is less vulnerable to price fluctuations and black-swan events than a siloed stablecoin, and is not reliant on any trusted third parties unlike existing stablecoins such as Tether and Paypal USD.
+By facilitating stablecoin use on smaller chains and stablecoin transfers between chains, CroCoDai is a practical medium of exchange in cross-chain commerce (datta2024blockchain, ).
+To aid reproducibility, we have made the data and the source code for our experiments publicly available via https://github.com/ntublockchain/crocodai.
 The remainder of the paper is structured as follows.
-Section 2 provides background on DeFi and related work.
-Section 3 presents the system model.
+Section 2 provides background information on DeFi and related work.
 Section 3 presents the design of CroCoDai.
-Section 5 contains the analysis of CroCoDai’s resilience against black swan events and price fluctuations.
-Section 6 details the three subvariant implementations and corresponding experimental results.
+We analyze CroCoDai’s resilience against black swan events in Section 4 and price fluctuations in Section 5.
+Section 6 details the prototype implementation and corresponding experimental results.
 Section 7 discusses future work and concludes the paper.
-2 Background
-2.1 General Background
-Blockchains & Tokens.
-DeFi uses cryptographically secured, decentralized data structures to replace trusted intermediaries.
-Blockchains are by far the most prominent example of such data structures.
-A blockchain is an ordered list of transactions grouped into blocks.
-The principal function of transactions in DeFi is to transfer tokens, which represent ownership of assets.
-There are two common types of tokens: fungible tokens – also commonly called cryptocurrencies – and non-fungible tokens (NFTs) [13].
+2. Background
+2.1. General Background
+Tokens and DeFi.
+In DeFi, the principal function of blockchain transactions is to transfer tokens, which represent ownership of assets.
+There are two common types of tokens: fungible tokens – also commonly called cryptocurrencies – and non-fungible tokens (NFTs) (das2022understanding, ).
 Fungible tokens of the same type are identical, whereas each NFT represents a unique asset.
-In the first cryptocurrency platform, Bitcoin [45], each transaction transfers some amount of the native cryptocurrency (bitcoin, or BTC) between parties.
-In Ethereum [8], transactions were extended from transferring the native cryptocurrency (ETH) to the creation or execution of smart contracts, i.e., on-chain software programs.
-Smart contracts can store data, hold tokens, and transfer tokens according to the contract’s internal logic, e.g., to implement an auction.
-A major use case for smart contracts is to define new cryptocurrencies that are different from the blockchain’s native cryptocurrency.
-In Ethereum, this is particularly facilitated by the ERC20 token standard [51] which defines a common set of smart contract functions required by any new cryptocurrency.
+Smart contracts can be used to define cryptocurrencies that are different from the blockchain’s native cryptocurrency.
+In Ethereum, this is particularly facilitated by the ERC20 token standard (erc20, ) which defines a common set of smart contract functions required by any new cryptocurrency.
 Smart contracts can also support NFTs, e.g., via OpenSea or Mintable.
+A prominent use case for user-defined tokens is to create a version of one blockchain’s token on another blockchain – e.g., Wrapped Bitcoin (WBTC) on the Ethereum blockchain.
+In general, wrapped tokens (wbtc, ) are tokens maintained by custodians – e.g., for WBTC this includes the Kyber Network and BitGo – who are trusted to create wrapped tokens if and only if an equivalent amount of tokens under their control have been locked on the supporting blockchain.
+Xclaim (zamyatin2019xclaim, ) is an academic proposal that is similar to WBTC but which requires custodians to deposit additional collateral, thus adding a financial penalty as a disincentive for abuse (e.g., stealing the locked tokens).
 Stablecoins.
 A stablecoin is a cryptocurrency whose value is pegged to an underlying fiat currency, e.g., the US dollar.
-[^1] Cryptocurrencies whose value is pegged to a real asset such as gold, e.g., the PAXG token, are typically not considered stablecoins, e.g., they are not labelled as such by CoinMarketCap.
-There are multiple types of stablecoins [44, 32], of which the most common ones are:
-  1. 1.
+There are multiple types of stablecoins (moin2020sok, ; klages2020stablecoins, ), of which the most common ones are:
+  1. (1)
     Fiat-backed stablecoins, which are backed by fiat currency held largely in cash or cash equivalents.
 The essential guarantee that underpins the price stability of a fiat-backed stablecoin is that in case of system shutdown, each stablecoin can be used to redeem a corresponding amount of the pegged fiat currency.
-  2. 2.
+  2. (2)
     Crypto-collateral-backed stablecoins (or crypto-backed stablecoins for brevity), which are backed by other cryptocurrencies that are stored in one or more smart contracts.
-In this case, the guarantee is that in case of system collapse, each stablecoin can be redeemed for a basket of tokens whose value equals one unit of the pegged currency.
-  3. 3.
-    Algorithmic stablecoins, which do not use reserves or collateral and which attempt to maintain a peg by adjusting the supply of a paired token [44].
-A multitude of fiat-backed stablecoins exist, the most prominent (as of June 2023) being Tether (USDT), USD Coin (USDC), and Binance USD (BUSD).
-The largest crypto-backed stablecoin is the Dai stablecoin (DAI), which is maintained by the MakerDAO foundation and which we discuss in more detail in Section 2.2.
-During our observation period (as discussed in Section 5), most fiat- and crypto-backed stablecoins did an overall sound job in maintaining their peg with the US dollar: the observed price ranges of DAI, USDT, USDC, and BUSD were [0.9921, 1.0062], [0.9834, 1.0005], [0.9991, 1.0073] and [0.9988, 1.0126] respectively.
-The analysis in [37] suggests that fiat-backed stablecoins have a trade-off between price stability and risk of system failure in the sense that better arbitrage improves the former at the cost of increasing the latter.
+In this case, the guarantee is that in case of system shutdown, each stablecoin can be redeemed for a basket of tokens whose value equals one unit of the pegged currency.
+  3. (3)
+    Algorithmic stablecoins, which do not use reserves or collateral and which attempt to maintain a peg by adjusting the supply of a paired token (moin2020sok, ).
+A multitude of fiat-backed stablecoins exist, the most prominent (as of September 2023) being Tether (USDT), USD Coin (USDC), Binance USD (BUSD), and TrueUSD (TUSD).
+The most prominent crypto-backed stablecoin is DAI on the Ethereum blockchain, which is maintained by the MakerDAO foundation and which we discuss in more detail in Section 2.2.
+Other examples of crypto-backed stablecoins include Liquity USD (LUSD) and Reflexer’s RAI, which are backed entirely by ETH.
+Other recent stablecoin proposals include FRAX (kazemian2022frax, ), which is an algorithmic stablecoin with a cryto-backed component, and Celo (clabs2019celo, ), which is backed by a mixture of fiat and crypto assets.
+During our observation period (see Section 5), most fiat- and crypto-backed stablecoins did an overall sound job in maintaining their peg with the US dollar: the observed price ranges of DAI, USDT, USDC, BUSD, and TUSD were [0.9921, 1.0062], [0.9834, 1.0005], [0.9991, 1.0073], [0.9988, 1.0126], and [0.9937, 1.0056], respectively.
+The analysis in (ma2023stablecoin, ) suggests that fiat-backed stablecoins have a trade-off between price stability and risk of system failure in the sense that better arbitrage improves the former at the cost of increasing the latter.
+Recently, established financial services have launched their own fiat-backed stablecoins, e.g., PayPal’s PYUSD, which is available on the Ethereum and Solana blockchains (paypal, ).
+Finally, there have been proposals by central banks to issue their own digital tokens, which would have the same functionality as stablecoins – e.g., the digital euro, which recently completed its 2-year investigation phase and is now moving to the preparation phase (euro, ).
 Fiat-backed stablecoins introduce a trust assumption, namely on the agents who handle and store the fiat currency.
-In Tether, this trust assumption has been widely questioned – accusations include Tether using fiat collateral to buy high-risk assets [20] and creating unbacked stablecoins to prop up the price of bitcoins [25].
-Furthermore, algorithmic coins have nothing to underpin them in case of a death spiral [44, 6], as witnessed by the loss of over 60 billion USD when the Terra/Luna stablecoin system failed.
-As such, in the rest of this paper we will focus on crypto-backed stablecoins such as Dai.
-2.2 The Dai Stablecoin System (DSS)
-In this section, we describe the core features of the existing single-chain DSS.
-Although there are documents and code repositories that describe the DSS in its various stages of evolution, e.g., [38, 39, 40, 44, 3, 10], we focus on the GitHub code [40] and the description of the multi-collateral DSS [39].
+In Tether, this trust assumption has been widely questioned – accusations include Tether using fiat collateral to buy high-risk assets (tetherallegation, ) and creating unbacked stablecoins to prop up the price of bitcoins (griffin2020bitcoin, ).
+Furthermore, algorithmic coins have nothing to underpin them in case of a death spiral (moin2020sok, ; briola2023anatomy, ), as witnessed by the loss of over 60 billion USD when the Terra/Luna stablecoin system failed.
+The main threat to crypto-backed stablecoins is that the system may fail if the value of the collateral decreases too rapidly – however, we show in this paper that this risk can be made small with sufficient overcollateralization.
+In the following, we focus particularly on Dai, which is the most prominent crypto-backed stablecoin.
+2.2. The Dai Stablecoin System (DSS)
+In this section, we describe the existing single-chain DSS.
+Although there are documents and code repositories that describe the DSS in its various stages of evolution, e.g., (daiwhitepaper, ; multicollateraldai, ; daigithub, ; moin2020sok, ; bhat2021simulating, ; chen2022understanding, ), we focus on the GitHub code (daigithub, ) and the description of the multi-collateral DSS (multicollateraldai, ).
 Dai Tokens.
@@
 The procedure for a user, Alice, who wants to mint new Sai tokens is as follows.
-First, Alice must send a transaction to a Dai vault to create a Collateralized Debt Position (CDP) [38].
-Next, Alice must send ETH tokens as collateral to the contract to fund the CDP.
-If the CDP is funded, then Alice can use it to create Dai stablecoins and withdraw them from the CDP to her wallet, incurring debt.
+First, she must send a transaction to a Dai vault to create a Collateralized Debt Position (CDP) (daiwhitepaper, ).
+Next, she must send ETH tokens as collateral to the contract to fund the CDP.
+Alice can use a funded CDP to create Dai stablecoins and withdraw them to her wallet, incurring debt.
 At any later time, she can send Dai tokens back to the CDP to repay her debt.
-The above can be generalized to a multi-collateral (but single-chain) DSS by associating each CDP with a token type.
-Debt increases exponentially over time, and the rate at which it increases is a system parameter called the stability fee.
+The above can be generalized to a multi-collateral (but single-chain) DSS by associating each CDP with a token type: in the DSS, this could theoretically include any ERC20 token, including wrapped tokens such as WBTC.
+Debt increases exponentially over time, and the rate at which it increases is determined by the stability fee.
 Liquidations.
-At all times, Dai CDPs must be overcollateralized, i.e., the ratio of the total value of the Alice’s collateral to her debt must exceed some threshold called the liquidation ratio [3, 38].
-By default, the liquidation ratio in Dai is set to $1.5$ – i.e., Alice can withdraw Dai stablecoins worth at most $\frac{2}{3}$ of the total value of the collateral in their CDP [44].
+At all times, Dai CDPs must be overcollateralized, i.e., the ratio of the total value of the Alice’s collateral to her debt must exceed some threshold called the liquidation ratio (bhat2021simulating, ; daiwhitepaper, ).
+By default, the liquidation ratio in Dai is set to $1.5$ – i.e., Alice can withdraw Dai stablecoins worth at most $\frac{2}{3}$ of the total value of the collateral in their CDP (moin2020sok, ).
 Similarly, Alice can return Dai and withdraw collateral from her CDP as long as the above condition holds.
@@
 CDPs can go underwater either because the value of the collateral has dropped due to price fluctuations, or because the stability fee has caused the debt to become too high.
-To avoid liquidation, Alice must either deposit additional collateral, or raise Dai to clear the position when it is at risk – if this fails, a collateral auction can be started by other users.
-During a collateral auction, other users can bid to buy (portions of) Alice’s collateral using Dai stablecoins.
-Once enough Dai has been raised to repay the CDP’s total debt, the remaining collateral (if any) is sent back to Alice, minus a penalty.
+To avoid liquidation, Alice must either deposit additional collateral, or raise Dai to clear an at-risk position – if this fails, a collateral auction can be started by other users.
+During a collateral auction, other users can bid to buy (portions of) Alice’s collateral using Dai tokens.
+Once the CDP’s debt has been cleared, Alice receives the remaining collateral (if any) minus a penalty.
 Price Feeds.
-To determine the value of deposited collateral, MakerDAO uses information from price feeds supplied by oracles [35].
-The number of price feeds used by the DSS has increased from 14 between 2017 and November 2019 [26] to 20 currently [43].
-Smart contracts that maintain CDPs – vaults – periodically update token prices by setting each price equal to the median of a set of recent prices for the associated token.
-Dai Savings Rate.
-Due to the stability fee, the total amount of debt in the protocol may grow to exceed the total number of stablecoins in circulation.
-For example, if Alice’s debt of 100 Dai has grown to 105 Dai, then clearing her debt and closing the CDP would (in principle) leave 5 Dai in the vault.
-These surplus Dai are part of the Dai Savings Rate (DSR) mechanism, which allows users who deposit stablecoins in a DSR contract to periodically withdraw surplus Dai tokens from associated vaults.
-The DSR helps MakerDAO to keep the value of Dai soft-pegged to the dollar by reducing or decreasing the interest rate: if Dai becomes less valuable than the dollar, then the interest rate is increased to incentivize users to buy Dai to deposit in the DSR contract (and decreased for higher Dai values).
+To determine the value of deposited collateral, MakerDAO uses information from price feeds supplied by oracles – e.g., Chainlink or Uniswap contracts (liu2021first, ).
+The number of price feeds used by the DSS has increased from 14 between 2017 and November 2019 (gu2020empirical, ) to 20 in late 2023 (mips, ).
+Smart contracts that maintain CDPs – i.e., the vaults – periodically update token prices by setting each price equal to the median of a set of recent prices for the associated token.
 Governance.
@@
 Such decisions include 1) verifying and adding smart contracts that implement any of the functionalities above, 2) changing the Dai savings rate, and 3) occasionally changing other system-level parameters.
-The following are the main system-level parameters (see page 15 of [39]): 1.
-the debt ceiling for each token type, i.e., the maximum amount of total debt that can be created using a given collateral type, 2.
-the stability fee, 3.
-the liquidation ratio, 4.
-the liquidation penalty, 5.
-the duration of collateral auctions, 6.
-the duration of auction bids, and 7.
-the minimum auction step size (which prevents bids that are only marginally above the previous bid, which would waste transaction fees).
+The following are the main system-level parameters (see page 15 of (multicollateraldai, )): (1) the debt ceiling for each token type, i.e., the maximum amount of total debt that can be created using a given collateral type, (2) the stability fee, (3) the liquidation ratio, (4) the liquidation penalty, (5) the duration of collateral auctions, (6) the duration of auction bids, and (7) the minimum auction step size (to prevent bids with a minor increment).
 Out-of-Scope Concepts.
-Some of the concepts in the DSS are beyond the scope of this paper.
-In particular, these are 1) debt auctions, which occur when a collateral auction fails to raise enough Dai to cover the CDP’s debt, 2) surplus auctions, which occur when the amount of Dai left in vaults due to auctions and stability fees exceeds some threshold, and 3) (emergency) shutdowns, which can be triggered after a catastrophic event (e.g., system failure) and during which the total remaining collateral is distributed to Dai holders.
-2.3 Existing Cross-Chain Approaches
-Cross-chain deals [27] provide a modeling framework to describe sequences of cross-chain token transfers.
-A cross-chain deal consists of five phases: clearing, escrow, transfer, validation, and commitment.
-The goal of cross-chain deals is to guarantee atomicity, i.e., that either all transfers occur or none do.
-In [27], two frameworks for cross-chain deals are proven to satisfy technical safety (all honest users end up with an acceptable payoff) and liveness (no assets are escrowed forever, and if all users are honest then all transfers eventually happen): 1.
-hashed time lock contracts, 2.
-dedicated blockchains for commit messages .
-In cross-chain deals, the cost of a participant abandoning the protocol before completion (sore-loser attacks [56]) can be made explicit – this form of misbehavior can be disincentivized through a premium [18] paid by the parties that abandon the protocol before termination.
-Cosmos [7] and Polkadot [33] are blockchain platforms that were specifically developed with blockchain interoperability in mind, but which do not trivially interoperate with existing blockchains outside their ecosystems.
-Furthermore, Xclaim [58] and zkBridge [54] claim to provide secure cross-chain transfers, but have limitations when transferring to/from smaller blockchain platforms as discussed in the introduction.
-3 System Model
-3.1 Entities
-Our system model has the following types of entities. In the following, we use the notation $[n]$ for $\{1,\ldots,n\}$. A list of symbols can be found in Table 4 in Appendix A.
-  • Token: a type of cryptocurrency or an NFT (although typically the former due to the better availability of price information for cryptocurrencies).
-We denote the number of different tokens in the system by $M\geq 2$.
-  • Governance Token: this is a type of cryptocurrency that gives holders the right to vote on decisions that affect the functioning of the crypto-backed stablecoin.
-The governance token cannot be used as collateral for the stablecoin.
-  • Blockchain (or chain): an immutable, distributed ledger that tracks the ownership of its tokens.
-We assume that each chain either supports smart contracts or that a wrapped version of its native tokens exist on a chain that does (e.g., WBTC [53]).
-We denote the number of supported chains by $B\geq 2$, and the chain of token $m\in[M]$ by $b_{m}\in[B]$.
-  • Consensus Node: is responsible for maintaining the blockchain, e.g., through the creation and/or validation of new blocks.
-The nature of the consensus nodes depends on the blockchain’s consensus protocol – examples include variations of proof-of-work, proof-of-stake, or Byzantine Fault Tolerance (BFT) protocols.
-Each consensus node has some amount of consensus power – e.g., processing power in proof-of-work, staked tokens in proof-of-stake, or a private key in a BFT protocol.
-  • User: owns blockchain tokens and seeks to exchange them for tokens on another chain. We denote the number of users in our system by $U\geq 1$.
-  • Smart contract: a software program on a blockchain that may hold tokens which can be sent to other contracts or wallets depending on the program’s internal logic.
-  • Vault: a smart contract or address that holds tokens as collateral, and to/from which tokens can be sent/received leading to the possible creation/destruction of stablecoins managed by another smart contract.
-  • Price Oracle: a type of user who provides price information to the system [35], e.g., Chainlink [5].
-We denote the number of price oracles in our system by $O\geq 1$.
-Each oracle $o\in[O]$ can provide price information for any selection of tokens.
-Let $p_{m\,o\,t}$ be the price of token $m\in[M]$ reported by oracle $o\in[O]$ at time $t\geq 0$.
-A price feed is a sequence of prices for a given token $m$ reported by an oracle $o$.
-3.2 Token Price Model
-[…]
-To model the evolution of token prices over time, we use Geometric Brownian Motion (GBM).
-This is a common choice in the financial literature on asset prices [47] and related work on stablecoins, e.g., by Salehi et al.
-[48], particularly due to the GBM’s role in the Black-Scholes formula for option prices.
-A stochastic process $S_{t}$ that follows GBM is defined using a stochastic differential equation:
-$dS_{t}=\mu S_{t}dt+\sigma S_{t}dW_{t},$ (1)
-where $\mu$ determines the drift of $S_{t}$ (i.e., its expected change per unit of time), $\sigma$ represents the volatility of $S_{t}$ (i.e., the variance of the change), and $W_{t}$ is the Wiener process that represents standard Brownian motion [47].
-In this model, each token price process $p_{m\,t}$, $t\geq 0$, can be represented using a process defined as in Equation 1.
-Essentially, this means that the logarithm of the relative difference of token prices, $\log\left(p_{m\,t+t^{\prime}}/p_{m\,t}\right)$, is normally distributed for all $t^{\prime}>0$ with mean and variance dependent on $t^{\prime}$.
-Unlike [48], we consider a setting with multiple assets; i.e., we have a stochastic process $S_{m\,t}$ for each token $m$ with its own mean $\mu_{m}$ and variance $\sigma^{2}_{m}$.
-In this setting, we must also consider the covariance between the token prices [47], as cryptocurrency prices are often highly correlated – we will discuss this in more detail in Section 5.
-More details about how the standard GBM can be used to simulate asset prices can be found in Appendix B.
-The standard GBM as decribed above is based on the normal distribution, i.e., $\log(S_{t+t^{\prime}}/S_{t})$ has a normal distribution.
-However, it is known from the academic literature that the probability distribution of stock price fluctuations is usually more heavy-tailed and that Student’s $t$-distribution is often a better fit [29, 34].
-This is also visible in the data for price changes in our dataset – see Figure 5 in Appendix C.
-As such, we will focus on the $t$-distribution in our analysis.
-Although the $t$-distribution does not allow for an analytical solution for the asset price change after each individual time slot (e.g., as per (5) in Appendix B), we can use Monte Carlo simulation to analyze its behavior, as we discuss in more detail in Section 5.
-3.3 DSS Execution Model
-Let $\mathcal{J}_{t}$ denote the set of all CDPs at time $t$, and $c_{j\,t}$ the total number of collateral tokens held in CDP $j\in\mathcal{J}_{t}$ at time $t$.
-Let $D_{j\,t}$ be the outstanding debt resulting from Dai withdrawals/deposits from/to CDP $j\in\mathcal{J}_{t}$.
-Each CDP is associated with a specific token type – let $y_{j}$ be the token type of CDP $j$.
-Debt is subject to fees, i.e., $D_{j\,t+1}=(D_{j\,t})^{1+\delta}$, where $\delta$ represents the stability fee.
-We denote the number of stablecoins held by user $u\in[U]$ at time $t$ by $h_{u\,t}$.
-We note that, in principle, $\sum_{u\in[U]}h_{u}\leq\sum_{j\in\mathcal{J}}D_{j\,t}$, as stablecoins are created and transferred to a user account either through withdrawal from a CDP (in which case the created stablecoins and debt are equal) or from a vault through the Dai savings rate (in which case the created stablecoins cannot exceed the excess debt).
-However, we will consider attacks in Section 3.4 that can cause this assumption to be violated.
-We formalize the liquidation criterion as follows: for each CDP $j\in\mathcal{J}_{t}$ and $t\geq 0$, it must hold that $p_{y_{j}\,t}c_{j\,t}/y_{j\,t}>\gamma,$ where $\gamma$ is the liquidation ratio – if this condition does not hold, then CDP $j$ can be liquidated through a collateral auction.
-We assume that under standard conditions (i.e., the absence of an attack or sudden price collapse), collateral auctions are concluded successfully, i.e., a sufficient amount of stablecoins is raised to close the position.
-No user can create debt if the result would violate the liquidation criterion.
-We also note that price fluctuations do not affect the liquidation criterion, nor does the Dai savings rate as the proportional effect on all CDP debts is the same.
-We denote the debt ceiling for each token type $m\in[M]$ by a fraction $\zeta_{m}\in[0,1]$, i.e., the total debt in CDPs of type $m$ relative to the system’s total debt cannot exceed $\zeta_{m}$.
-(Debt ceilings are specified in absolute terms in DSS contracts, but we assume that the difference is negligible for short periods, e.g., for escrow.) The debt ceilings are used to enforce $\sum_{\begin{subarray}{c}j\in\mathcal{J}_{t}\\ y_{j}=m\end{subarray}}D_{j\,t}<\zeta_{m}\sum_{j\in\mathcal{J}_{t}}D_{j\,t}$ for all $m\in[M]$.
-Let $D_{t}^{*}=\sum_{j\in\mathcal{J}_{t}}D_{j\,t}$ and $C^{*}_{t}=\sum_{j\in\mathcal{J}_{t}}p_{y_{j}\,t}c_{j\,t}$ denote the total debt and the total collateral value across the system at time $t$, respectively.
-3.4 Threat Model
-We make the following assumptions about the system entities.
-We assume that a majority of the governance token holders are honest.
-For each underlying blockchain, we assume that it is either compromised or not.
-If it is not compromised, we assume that it guarantees safety and liveness for all of its transactions.
-However, both safety and liveness can be violated in a compromised blockchain – in particular, transactions in which collateral is sent to/from a CDP can be overturned.
-For the price oracles, we assume that a majority is online and honest at all times, but that dishonest oracles can report arbitrary prices.
-3.5 System Requirements
-The central requirement that we focus on in this paper is the following:
-  1. (*)
-    Full Backing. At any time $t\geq 0$, the ratio of the value of the collateral tokens to the total number of stablecoins in the system exceeds a safety threshold $\theta$, i.e.,
-    $C^{*}_{t}/D^{*}_{t}>\theta.$ (2)
-We assume that $\theta>1$, i.e., system failure occurs when the ratio of the collateral value to the debt gets sufficiently close enough to $1$.
-The motivation behind this requirement is as follows: the fundamental guarantee that underpins crypto-backed stablecoins is that in the event of a system shutdown, there is enough collateral in the vaults to refund each user who holds $h$ stablecoins with a basket of tokens whose value equals $h$ units of the pegged asset.
-If price fluctuations or an attack would cause this requirement to be violated, then we can reasonably expect the stablecoin to lose its peg.
-This would be detrimental to the user experience as their tokens held in escrow would depreciate, causing financial losses.
-Due to random price fluctuations, we cannot guarantee with absolute certainty that (2) will hold in all circumstances.
-However, we use our analysis to show that for reasonable choices of $\theta$ and the liquidation ratio $\gamma$, the probability that (2) is violated after a medium-length period (i.e., one day) is very small for historical data.
-We do this both for portfolios that approximate the current state of the DSS (as of late April 2023), and portfolios that are optimized for minimal risk given a selection of possible single-chain or cross-chain tokens.
-In Section 5, we also investigate under which conditions (2) continues to hold in the event of a black swan event such as a sudden token price collapse or a 51% attack on a blockchain.
-Furthermore, we require the following properties under the threat and token price model.
-  1. 1.
-    Safety. Users do not lose tokens during stablecoin transfers.
-  2. 2.
-    Liveness. Stablecoin transfers eventually happen.
-  3. 3.
-    Safety of Collateral (informal). The probability that a minority of corrupted oracles cause a CDP that is not underwater to be inadvertently liquidated is small.
-More formally, for the last property we consider a minority of corrupted price oracles whose goal is to convince the protocol to accept a price $p^{*}$ instead of a true base price $p$ – safety of collateral then holds if their probability of success vanishes exponentially as $|p-p^{*}|$ increases (see also Theorem 5.3 in Section 5.1).
-As we will see in Section 5, the former two properties largely follow from results in the existing scientific literature.
-The latter is a result that we prove for the multi-chain setting in Section 5, but which is also applicable in the single-chain setting.
-4 CroCoDai: A Cross-Chain Stablecoin
-Figure 1: Overview of the single-chain DSS.
+The following DSS concepts are beyond the scope of this paper: (1) the Dai Savings Rate mechanism, which allows users to deposit Dai to collect excess stablecoins created from the stability fee, (2) debt auctions, which occur when a collateral auction fails to raise enough Dai to cover the CDP’s debt, (3) surplus auctions, which occur when the amount of Dai left in vaults due to auctions and stability fees exceeds some threshold, and (4) emergency shutdowns, which can be triggered after a catastrophic event (e.g., system failure) and during which the total remaining collateral is distributed to Dai holders.
+2.3. Related Work
+The DSS can be seen as an extension of red-black coins (salehi2021red, ), which feature two token types of which one has reduced price volatility (e.g., Dai tokens vs.
+CDP debt).
+In (salehi2021red, ), single-asset price simulations are used to estimate the price of the volatile tokens – the main difference with the analysis in Section 5 is that we estimate failure risks, that we use multiple correlated assets, and that we use a heavy-tailed distribution to better represent token price fluctuations.
+Price simulations of DSS collateral can also be found in (bhat2021simulating, ), and (klagesmundt2022while, ) presents a stochastic model for a single-asset stablecoin.
+In (grobys2021stability, ), the relationship between stablecoin and Bitcoin price fluctuations is investigated.
+In (ma2023stablecoin, ), fiat-backed stablecoins are found to have a trade-off between price stability and risk of system failure.
+For secure cross-chain transfers, cross-chain deals (herlihy2021cross, ) provide a modeling framework to describe sequences of cross-chain token transfers.
+In cross-chain deals, the cost of a participant abandoning the protocol before completion (sore-loser attacks (xue2021hedging, )) can be made explicit – this form of misbehavior can be disincentivized through a premium (engel2021failure, ) paid by the parties that abandon the protocol before termination.
+Cosmos (Polkadot20, ) and Polkadot (cosmos20, ) are blockchain platforms that were specifically developed with blockchain interoperability in mind, but which do not trivially interoperate with existing blockchains outside their ecosystems.
+However, a version of CroCoDai could be implemented as a native stablecoin on these platforms with collateral on multiple subchains.
+Alternatively, collateral on multiple chains can be integrated into CroCoDai through trusted relays or zero-knowledge proofs as discussed in Section 3.
+Finally, Xclaim (zamyatin2019xclaim, ) claims to provide secure cross-chain transfers, but has limitations when transferring to/from smaller blockchain platforms that lack sufficient liquidity because the custodian must submit additional collateral.
+3. CroCoDai: A Cross-Chain Stablecoin
+Figure 2. Overview of the existing single-chain DSS.
 In this section, we present CroCoDai, which extends the functionality of the single-chain DSS (depicted in Figure 2) to a multi-chain setting (depicted in Figure 3).
-Our main design goal is to have stablecoin tokens that are backed by collateral on multiple blockchains, and which can be moved securely and efficiently between the different chains.
-This is non-trivial as transactions between different blockchains are not atomic by default, so our design incorporates a secure cross-chain protocol.
-In the following, we will refer to blockchains that store tokens as coin chains (as per [27]).
-The communication between the different blockchains is managed by a relay chain, which may be a coin chain or a separate chain.
-The main difference between the single-chain DSS and CroCoDai is the inclusion of relay nodes – depicted with double circles in Figure 2 – which is a non-intrusive change as the other DSS contracts remain unchanged.
-In the following, we first discuss in detail which components as discussed in Section 2.2 remain local to each chain and which require interaction with the relay chain.
-We then discuss the essential components of the smart contracts on the coin chains and the relay chain.
-Finally, we discuss two ways to implement the relay chain: one that relies on a quorum of trusted nodes, and one that uses an existing blockchain.
-We will present three implementations of the former in Section 6.
-4.1 Overview
+Our main design goal is a stablecoin that is backed by collateral on multiple blockchains, and whose tokens can be moved securely and efficiently between chains.
+In the following, we will refer to blockchains that store collateral and stablecoin tokens as coin chains (as per (herlihy2021cross, )).
+The communication between the coin chains is managed by a relay chain, which may be an existing (coin) chain.
+The main difference between the single-chain DSS and CroCoDai is the inclusion of relay nodes – depicted with double circles in Figure 3.
+In CroCoDai, all vaults are local to the coin chains, and cross-chain price oracles are supported to complement the oracles on smaller coin chains.
+This allows CroCoDai to have lower costs for local operations such as creation and transfer of stablecoins on smaller blockchain platforms (we will discuss this in more detail in Section 6.)
+In the following, we first discuss in more detail which DSS operations remain local and which require interaction with the relay chain.
+We then discuss the main components of the coin and relay chains.
+Next, we discuss two design variants for the relay chain: the first relies on a quorum of trusted nodes who perform most operations off-chain, and the second uses an existing blockchain.
+We conclude by listing the security requirements of a cross-chain stablecoin, which will be the subject of the Sections 4 and 5.
+3.1. Overview
 Local Operations.
-In CroCoDai, each coin chain keeps track of the balance of stablecoins in accounts on that chain.
-The stablecoins can be exchanged locally between accounts.
-Furthermore, each coin chain has its own vaults that allow users to create CDPs using tokens on that blockchain.
-Since CDPs are local, liquidation auctions and savings rate mechanisms are also processed locally.
-Finally, price feeds are still predominantly local, although price feeds from the relay chain can add an additional source of price information.
+CroCoDai supports the following local operations.
+CDP Management: Each coin chain has its own vaults that allow users to create CDPs using tokens on that blockchain.
+As such, stablecoins can be created locally on minor chains, i.e., without the need for transactions on a more expensive chain, as long as the debt limit on the minor chain is not violated.
+Since CDPs are local, liquidation auctions and other mechanisms such as savings rates are also processed locally.
+Local Stablecoin Transfers: Each coin chain keeps track of the balance of stablecoins held by its accounts, so the stablecoins can be exchanged locally between accounts.
+Local Price Information: Finally, local oracles may provide the vaults with price information.
 Cross-Chain Operations.
-The relay chain has four main functions in CroCoDai.
-The first is to enable cross-chain transactions.
-In a cross-chain transaction, stablecoins are burned on one coin chain and an equivalent number is minted on another chain.
-The second function is to add/remove addresses that are authorized to call security-critical functions such as mint or burn on another contract – we call such addresses wards as per DSS terminology.
-Periodically, smart contracts may be added or deprecated – e.g., 1.
-to add support for an emerging blockchain platform, 2.
-because a critical bug has been found in an existing smart contract, 3.
-because a chain has undergone a hard fork that interferes with the functioning of the contract, or 4.
-to implement a more efficient version of a component (e.g., upgrading the collateral auction mechanism from Cat.sol to Dog.sol in the DSS).
-The third function is to update the system parameters discussed in Section 2.2.
-The final function is to provide an additional source of token price information to vaults through relay chain oracles – this may particularly benefit smaller blockchains that lack a mature DeFi ecosystem.
-The mechanism for cross-chain transfers is as follows: a node that holds stablecoins on a coin chain must create a transaction that stores the tokens to be transferred in escrow – furthermore, the transaction takes as input two strings to represent the target blockchain and the address on the target blockchain.
-This is then added to a list where it awaits commitment.
-The nodes in the relay network monitor the coin chains, and once they detect a cross-chain transfer request, they will sign a commit message.
-When the coin chain receives the commit message, it will commit the transaction and burn the tokens held in escrow.
-A commit message is also sent to the target blockchain, upon which it will mint an equivalent number of stablecoins.
+The following operations require relay chain involvement.
+Cross-chain Transactions: In a cross-chain transaction, stablecoins are burned on one coin chain and an equivalent number is minted on another coin chain, as we discuss in more detail below.
+Ward Management: Governance layer approval is needed to add/remove addresses that are authorized to call security-critical functions such as mint or burn on another contract – we call such addresses wards as per DSS terminology.
+Periodically, the governance layer may grant or revoke ward status to/from smart contracts for a variety of reasons – e.g., (1) to add support for an emerging blockchain platform, (2) because a critical bug has been found in an existing smart contract, (3) because a chain has undergone a hard fork that interferes with the functioning of the contract, or (4) to implement a more efficient version of a component.
+System Parameter Updates: Governance layer approval is also needed to update the system parameters discussed in Section 2.2, e.g., a chain’s debt ceiling or a token’s overcollateralization rate on a chain.
+Global Price Information: The final function is to provide an additional source of token price information to vaults through relay chain oracles – this particularly benefits smaller blockchains that lack a mature DeFi ecosystem.
+The exact mechanism for cross-chain transfers is as follows: a user who holds stablecoins on a coin chain first submits a cross-chain transfer request to a relay node and transfers the tokens for escrow.
+The request is added to a hashmap where it awaits commitment.
+The nodes on the relay chain monitor the coin chains, and once they detect a cross-chain transfer request, they sign a commit message.
+When the coin chain receives the commit message, it commits the transaction and burn the tokens held in escrow.
+A commit message is also sent to the relay node on the target blockchain, which mints an equivalent number of stablecoins.
 If the transaction is aborted for any reason, then the stablecoins in escrow are returned to the sender.
-Figure 2: Overviews of CroCoDai, a multi-chain stablecoin system based on the DSS. The relays (double circle) are components that are not shared with the single-chain DSS.
-4.2 Coin Chain Components
+Figure 3. Overviews of CroCoDai, a multi-chain stablecoin system based on the DSS. The relays (double circle) are not shared with the single-chain DSS.
+3.2. Core Components
+Coin Chain Components.
 Each coin chain has the following components.
-Coin: implements a contract that is similar to Dai.sol in the DSS, and processes transfers of stablecoins between users on the coin chain.
+Coin: implements a contract that is similar to Dai in the DSS, and processes transfers of stablecoins between users on the coin chain.
 There is only one coin contract per chain.
-Vault: implements contracts that are similar to Vat.sol and Join.sol in the DSS, which allow users to create CDPs and withdraw/deposit collateral and stablecoins.
+Vault: implements contracts that are similar to Vat, GemJoin and DaiJoin in the DSS, which allow users to create CDPs and withdraw/deposit collateral and stablecoins.
 There can be multiple vaults per chain.
-Auction: implements auction contracts that liquidate the collateral of underwater CDPs, with any appropriate auction logic (not necessarily limited to Cat.sol or Dog.sol) in the DSS.
+Auction: implements auction contracts that liquidate the collateral of underwater CDPs, with any appropriate auction logic – e.g., Cat or Dog in the DSS.
 There can be multiple auction contracts per chain.
-Savings: implements a savings rate contracts that are similar to Pot.sol and Jug.sol in the DSS.
-Bottom-Level Oracle: implements a price feed using a contract similar to Spot.sol in the DSS, combined with a Medianizer.
+Bottom-Level Oracle: implements a price feed using a contract similar to Spot in the DSS, combined with a Medianizer.
 Bottom-Level Relay: validates messages from the relay chain.
-It processes 1.
-adding wards to each of the aforementioned smart contracts, 2.
-incoming cross-chain transfers, 3.
-commitments of outgoing cross-chain transfers, and 4.
-updates of system parameters.
-4.3 Relay Chain Components
+It processes (1) the addition of wards to CroCoDai contracts, (2) incoming cross-chain transfers, (3) commitments of outgoing cross-chain transfers, and (4) updates of system parameters.
+Relay Chain Components.
 The relay chain has the following components.
-Governance: allows users to transfer governance tokens.
-Users who hold governance tokens can vote on system-level actions such as adding wards, processing cross-chain transfers, and updating system-level parameters.
+Governance: tracks the ownership of governance tokens, and allows users to transfer them.
+As in the DSS, users who hold governance tokens can vote on system-level actions such as adding wards and updating system-level parameters.
+Our security assumption regarding governance token holders is that a majority is honest and online at any point during the voting window for system-level decisions.
 Top-Level Oracle: The relay chain may implement system-level price feeds to complement the price feeds on the coin chains.
-Top-Level Relay: To process cross-chain transfers, the relay network must be able to read the state of the coin chains, in particular the list of cross-chain transfer requests.
-4.4 Relay Chain Designs
-In this section, we discuss two possible implementation designs of the relay chain.
+Top-Level Relay: To process cross-chain transfers, the relay chain must be able to process information from the coin chains, particularly the transfer requests – we discuss this in more detail below.
+3.3. Relay Chain Designs
+In this section, we discuss two possible design variants of the relay chain.
 The first relies on a quorum of trusted nodes with known public keys – each trusted node runs an off-chain (full or light) blockchain client for each of the underlying chains to check the validity of cross-chain transfers, and the nodes maintain a distributed ledger for governance decisions.
-The second uses an existing blockchain, and uses smart contracts that implement the consensus logic of the underlying chains on-chain to verify cross-chain operations.
+The second uses smart contracts on an existing chain to implement the consensus logic of the underlying chains (potentially using zero-knowledge cryptography) to verify cross-chain operations on-chain.
 Off-Chain Blockchain Clients.
 In this approach, the stablecoin system is maintained by $n$ relay nodes with known public keys.
-The relay nodes use their (light) clients to detect and validate cross-chain transactions, and generate signatures using their private/public key pair to commit cross-chain transfers and approve governance decisions.
+The relay nodes use their (light) clients to detect and validate cross-chain transactions, and generate signatures using their private/public key pair to commit cross-chain transfers and process system-level actions approved by governance token holders.
 In this setting, the only functionality of the bottom-chain relay contracts is to verify digital signatures.
-If different chains use different built-in signature schemes, or the same scheme with a different parametrization, then the relay nodes may need to generate multiple signatures for a single cross-chain transfer (this is safe if the threat model of Section 3.4 holds when each signature is generated).
-We discuss three different implementations of this approach in Section 6.
-To ensure security, we note that it is vital that the relay nodes are controlled by independent entities: e.g., in the attack against the Ronin bridge, 5 out of 9 validators were successfully attacked [30], of which 4 belonged to the same entity [36].
+We discuss three different subvariants of this approach in Section 6.
+As is common in Byzantine fault tolerant systems (castro1999practical, ; lamport1982byzantine, ), we assume that at most $f$ nodes are faulty or malicious such that $n\geq 3f+1$.
+We note that it is vital that the relay nodes are controlled by independent entities: e.g., in the attack against the Ronin bridge, 5 out of 9 validators were successfully attacked (roninhack, ), of which 4 belonged to the same entity (roninhack2, ).
 On-Chain Consensus Logic.
@@
 Essentially, the functionality of a light client – i.e., validating block headers and proving the state of a contract using the header – is implemented inside the smart contracts.
-This is the approach taken in, e.g., Xclaim [58].
-For example, in Bitcoin’s Nakamoto Consensus [45] this means the ability to compute header hashes and compare them to the proof-of-work difficulty, whereas in Ethereum’s proof-of-stake it means validating that nodes who control a sufficient amount of stake have approved a block.
-Techniques to increase the efficiency of light clients have been proposed in the literature for both proof-of-work (e.g., Bitcoin [14]) and proof-of-stake blockchains (e.g., Algorand [42] and Mina [4]).
-Another variation of this approach is to encode the consensus logic of the blockchains in a zero-knowledge proof, e.g., using the zkBridge protocol [54].
-Finally, the data on the different coin chains can be integrated using TrustBoost [52].
-Although this design avoids a trust assumption on a quorum of relay nodes, current iterations of such designs have a large computational overhead, typically requiring several kilobytes of data sent to relay contracts and more than 10 seconds to generate signatures [42, 54].
-As such, we focus on the off-chain design in Section 6.
-5 Analysis
-In this section, we analyze whether our system satisfies the requirements of Section 3.5.
-In Section 5.1, we focus on the technical properties of safety and liveness, full backing under the threat model, and safety of collateral.
-We then focus on full backing under the price fluctuation model: we present our approach in Section 5.2, and the evaluation using empirical data in Section 5.3.
-The proofs of the theorems in this section can be found in Appendix D.
-Our analysis shows that CroCoDai is resilient to serious failures of multiple coin chains, as long as the over-collateralization of the CDPs, $\gamma$, is high compared to the total debt ceiling of the affected chains (which we will denote by $\zeta^{\prime}$).
-We emphasize that CroCoDai is therefore fundamentally more resilient than a single-chain stablecoin such as the DSS, which cannot tolerate the failure of its single coin chain.
-Our analysis also shows that a broader selection of collateral tokens leads to a lower probability of system failure than a single-chain system such as the DSS.
-Although the findings of our price analysis strongly depend on the choice of data, we emphasize that we included a period of notoriously high volatility (i.e., the FTX collapse) and that the risks would appear smaller if we used a less volatile period.
-5.1 Safety Analysis
-Technical Safety and Liveness.
-For local operations, the consensus protocol of the coin chain on which the local transfer occurs ensures the safety and liveness of operations on this chain – see, e.g., [22].
-This includes local stablecoin transfers, operations on collateral, auctions, and interest withdrawals.
+This is the approach taken in, e.g., Xclaim (zamyatin2019xclaim, ).
+For example, in Bitcoin’s Nakamoto Consensus (btc_origin, ) this means that the smart contract computes header hashes and compares their integer value to the proof-of-work difficulty, [^4] This may be computationally efficient if blockchains share built-in hash functions, otherwise this can be expensive.
+whereas in Proof-of-Stake Ethereum the smart contract verifies that the total stake of the nodes who have approved a block is sufficient.
+In this setting, the main role of the relay nodes is to generate and submit validity proofs.
+Techniques to increase the efficiency of light clients have been proposed in the literature for both proof-of-work (e.g., Bitcoin (daveas2020gas, )) and proof-of-stake blockchains (e.g., Algorand (micali2021compact, ) and Mina (bonneau2020mina, )) – in some cases using zero-knowledge proofs that may incorporate the blockchain’s consensus logic (xie2022zkbridge, ).
+Zero-knowledge proofs that are verifiable on multiple chains, e.g., ZKM’s entangled rollups (zkm2024entangled, ), can also be used to bridge the chains that implement collateral to a layer-2 network that implements the vaults.
+Finally, the data on the different coin chains can be integrated using TrustBoost (wang2022trustboost, ).
+Although the second design variant avoids a trust assumption on a quorum of relay nodes, current iterations of such designs have a large computational overhead, typically requiring several kilobytes of data sent to relay contracts and more than 10 seconds to generate signatures (micali2021compact, ; xie2022zkbridge, ).
+[^5] Improving the performance of such zero-knowledge schemes, e.g., using folding schemes (kothapalli2022nova, ), is an active research area.
+As such, we focus on the first variant in Section 6 and leave the second as future work.
+3.4. System Requirements
+The central requirement that we focus on in this paper is the following:
+  1. (*)
+    Sufficient Overcollateralization. At any time $t\geq 0$, the ratio of the value of the collateral tokens to the total number of stablecoins in the system exceeds a safety threshold $\theta>1$.
+The motivation behind this requirement is as follows: if the value of the collateral becomes too low, then the system is no longer able to reimburse all stablecoin holders with an equivalently valued basket of tokens in case of system shutdown.
+As such, we can expect users to start selling their stablecoin tokens at a loss to redeem as much value as possible, causing the stablecoin to lose its peg.
+We assume that this sell-off may occur before the critical point is reached at which the value of issued stablecoins exceeds that of the collateral, because nervous investors may preempt the loss of the peg.
+As such, we choose a value of $\theta$ that is greater than $1$: a value of strictly $1$ corresponds to a stablecoin system that fails as soon as the aforementioned critical point is reached, whereas $\theta>1$ also covers scenarios in which system failure occurs earlier, in anticipation of the depegging.
+Due to random price fluctuations, we cannot guarantee with absolute certainty that full backing will hold in all circumstances.
+However, we use our analysis to show that for reasonable choices of $\theta$ and the liquidation ratio, the probability that full backing is violated after a medium-length period (i.e., one day) is very small for historical data.
+In Section 4, we also investigate under which conditions full backing continues to hold in the event of a black swan event such as a sudden token price collapse or a 51% attack on a blockchain.
+We also require the following properties under the threat and token price model.
+  1. (1)
+    Safety. Users do not lose tokens during stablecoin transfers and CDP operations.
+  2. (2)
+    Liveness. Stablecoin transfers and CDP operations eventually happen.
+  3. (3)
+    Safety of Collateral (informal). The probability that a minority of corrupted oracles cause a CDP that is not underwater to be inadvertently liquidated is small.
+More formally, for the last property we consider a minority of corrupted price oracles whose goal is to convince the protocol to accept a price $p^{*}$ instead of a true base price $p$ – safety of collateral then holds if their probability of success vanishes exponentially as $|p-p^{*}|$ increases (see also Theorem 4.3 in Section 4).
+As we will see in Section 4, the former two properties largely follow from results in the existing scientific literature.
+The latter is a result that we prove for the multi-chain setting in Section 4, but which is also applicable in the single-chain setting.
+3.5. Illustrative Application
+We consider an example of cross-chain commerce: a setting in which an asset – e.g., a basket of cryptocurrencies or an NFT – is auctioned on one chain, and for which there are interested bidders who own tokens on multiple other chains.
+For the duration of the auction, price stability is important because changes in the value of bids could lead to unnecessary resubmissions or withdrawals of bids, causing a diminished user experience and/or opportunity costs (xue2021hedging, ).
+The auction example allows us both to illustrate the end-to-end processes of creating and transferring stablecoins in CroCoDai, as well as to compare CroCoDai to existing approaches in terms of performance and user experience.
+For this example, we will refer to the chain on which the asset is sold as the ‘auction chain’ and to the chains on which bidders hold tokens as ‘coin chains’.
+To achieve an auction with price stability, we consider two main approaches: 1.
+bidders use a centralized exchange (CEX) to swap their tokens into stablecoins on the auction chain, 2.
+bidders using CroCoDai’s built-in mechanism for cross-chain transfers.
+The first represents the most straightforward way to achieve a cross-chain auction in the current ecosystem by leveraging established exchanges such as Coinbase or Binance, whereas the second is our proposal.
+We also discuss other approaches at the end of this section.
+Figure 4. Steps of a cross-chain exchange in the CEX baseline and CroCoDai. The edge labeling corresponds to the step numbers in Section 3.5.
+Comparison.
+In the first approach, we assume the existence of a centralized exchange that can be used to buy and sell stablecoins on the coin chain and auction chain.
+In particular, the bidder would then complete the following steps:
+  1. (1)
+    Create an account with the CEX if she doesn’t have one already.
+  2. (2)
+    Transfer her tokens to the CEX to add it to her wallet on the CEX.
+  3. (3)
+    Sell the tokens in her account and buy stablecoins, or swap the tokens for stablecoins directly if supported by the CEX, and withdraw the stablecoins to her account on the auction chain.
+  4. (4)
+    Place a bid by transferring the stablecoins to the auction contract.
+In CroCoDai, the bidder would complete the following steps:
+  1. (1)
+    Deposit her tokens into a CroCoDai vault.
+  2. (2)
+    Create new CroCoDai stablecoins and withdraw them to her account.
+  3. (3)
+    Issue request to the relay chain for a cross-chain transfer of her CroCoDai tokens on the coin chain to an account owned by her on the auction chain.
+  4. (4)
+    Place a bid by transferring the stablecoins to the auction contract.
+The fundamental difference between the two approaches is the reliance on a trusted third party in the first approach.
+By contrast, the second approach relies on a decentralized relay chain.
+Furthermore, the CEX baseline is more likely to suffer from a lack of liquidity: whereas steps 1 and 2 in the first baseline always rely on the CEX as a counterparty to swap tokens, in CroCoDai the bidder can always create new stablecoins and transfer them to the auction chain via the relay chain.
+In terms of fees and transaction delays, there are no reasons to expect major differences between the two approaches, since the cost of sending and processing cross-chain messages is typically small compared to the cost of monitoring the supported coin chains through a blockchain node, and the latter is a requirement for all cross-chain systems.
+Additional Intermediaries.
+In both approaches discussed above, either the CEX or CroCoDai’s relay chain acts an an intermediary to facilitate the cross-chain token transfer.
+Additional steps that involve intermediaries may be needed if 1) the CEX does not support sales of the bidder’s tokens, 2) CroCoDai does not accept the bidder’s tokens as collateral (e.g., because they belong to an unsupported type, or because the token type’s debt limit would be violated), 3) the auction contract does not accept any token sold by the CEX, or 4) the auction contract does not accept CroCoDai stablecoins.
+In those cases, a DEX may exists that allows for on-chain swaps between unsupported and supported tokens.
+The additional steps are involving DEX swaps are depicted in Figure 4 through step $2^{\prime}$, which replaces step 2, and step $4^{\prime}$, which replaces step $4$.
+Other Approaches.
+A cross-chain auction can also be implemented using the framework of cross-chain deals (herlihy2021cross, ).
+However, this requires that the auctioneer creates smart contracts on each of the supported coin chains, instead of the just the auction chain.
+Finally, the bidder could create a wrapped version of her tokens on the coin chain, but this would typically require another trusted intermediary and is therefore similar to the CEX baseline.
+3.6. Deployment
+Real-world use of CroCoDai would require the deployment of the relay chain, the smart contracts (e.g., coin and vault contracts) on the supported smaller chains, and the development of tools that allow users to interact with the system.
+The costs of developing and deploying the relay chain are similar to those of existing bridges such as Celer [^6] https://cbridge.celer.network/bridge/ethereum-polygon/eth/ and Relay [^7] https://relay.link/bridge/ethereum.
+The deployment of smart contracts has roughly the same cost as in a fully siloed system, and are small over long run compared to fees (e.g., at a gas price of 20GWei and an ETH price of \$3000, deploying a 1m gas contract would cost \$60).
+The throughput and scalability of the relay chain are highly dependent on the design choices and implementation, e.g., Celer’s cBridge claims a latency of at most 20 minutes during peak demand periods (cbridge, ), whereas the zkBridge prototype has a latency of under 2 minutes (xie2022zkbridge, ) but at a high computational cost, while the Hephaestus protocol (belchior2023hephaestus, ) claims to be able to process 600 cross-chain transactions in under 6 seconds.
+In practice, we expect the user interface design to be very similar to existing designs for the DSS: web apps such as summer.fi888 https://summer.fi/ allow users to interact with vaults and to process local and cross-chain transfers.
+The incentives for developers are the same as for Dai, i.e., a combination of fees, publicity (an app can draw attention to other services), and in some cases volunteer work.
+The active DAI ecosystem, including summer.fi, daistats.com (daistats, ), DeFiExplore, [^9] https://maker.defiexplore.com demonstrates that such incentives are sufficient in practice.
+Finally, the system parameters in CroCoDai can be set to similar values as in the DSS.
+The overcollateralization rates, i.e., $\gamma$, will likely be in the order of $30-50\%$ for the supported token types, as they are in existing Dai vaults (daistats, ).
+Similarly, the debt limits for smaller blockchains are expected to be similar to those for ERC20 tokens such as Aave and Gnosis – as in the DSS, these are proposed and voted on by governance token holders.
+4. Security Analysis
+In this section, we analyze whether our system satisfies the following technical properties from Section 3.4: safety and liveness, full backing under the threat model, and safety of collateral.
+We first discuss our threat model and then present our results.
+In particular, we prove that if tokens whose total debt limit [^10] We note that price crashes in different tokens and platforms may occur simultaneously, e.g., the collapse of FTX also had consequences for other platforms such as Solana.
+is $\zeta^{\prime}$ suffer a price crash (Section 4.3) or platform-level attack (Section 4.4), then full backing continues to hold if $\theta<\gamma(1-\zeta^{\prime})$, where $\gamma$ denotes the liquidation ratio (see Section 2.2).
+In practice, this allows stablecoin holders to judge whether they find the risk of system failure low enough given their estimations of $\theta$ and $\zeta^{\prime}$.
+We conclude by comparing the security of CroCoDai to the two baselines from the introduction.
+The technical proofs of the main theorems can be found in Appendix A.
+[…]
+4.1. Security Model
+In the following, we first discuss the threat assumptions that are specific to CroCoDai, and then discuss the threats and countermeasures shared with generic stablecoin and cross-chain systems.
+In CroCoDai’s security model, a majority of the governance token holders are honest and online, and the relay chain satisfies safety and liveness.
+Each underlying blockchain is either compromised or not.
+If it is not compromised, then we assume that it guarantees safety and liveness for all of its transactions.
+However, both safety and liveness can be violated in a compromised blockchain – in particular, transactions in which collateral is sent to/from a CDP can be rolled back.
+We assume that the underlying communication networks satisfy partial synchrony, i.e., bounded delivery times, to ensure that nodes do not wait indefinitely for information to be transmitted between chains.
+It has shown that liveness cannot be guaranteed in the presence of even a single faulty node in a fully asynchronous model (fischer1985impossibility, ).
+We do note that CroCoDai itself provides synchrony to the overall cross-chain process by requiring that a transfer is completed on the sending chain before processing it on the receiving chain, as depicted in Figure 4b.
+For the price oracles, a majority is online and honest at all times, in which case they report a measurement error that is random and normally distributed.
+Furthermore, the output of each oracle is independent of the output of other oracles.
+Dishonest oracles can report arbitrary prices.
+We emphasize that these assumptions are common in the scientific literature, e.g., Byzantine fault tolerant protocols also require honest and online (super)majorities and partial synchrony (lamport1982byzantine, ).
+As a cross-chain stablecoin system, CroCoDai also inherits threats that affect generic blockchains, cross-chain infrastructures, and stablecoins.
+Recent surveys that investigate the security of such systems are available for individual blockchains (homoliak2020security, ; zhang2019security, ), cross-chain systems (belchior2021survey, ; zamyatin2021sok, ; herlihy2021cross, ), and stablecoins (mell2023understanding, ).
+Threats that affect individual blockchains include the following categories:
+  • Network-layer threats, which may prevent users from connecting with honest blockchain nodes, potentially violating safety (if the remaining nodes are malicious) or liveness.
+Examples include eclipse attacks on blockchain peers, DNS/routing attacks, and Denial-of-Service (DoS) attacks.
+Countermeasures include generic network security proposals such as DNSSEC and BGPsec, or blockchain-specific defences such as SABRE (apostolaki2019sabre, ).
+    Consensus protocols that announce a single block proposer beforehand – e.g., many Proof-of-Stake protocols (brown2019formal, ) – may be vulnerable to targeted DoS attacks against the block proposer, violating liveness.
+CroCoDai itself does not introduce such a vulnerability because cross-chain transfers can be initiated by any of the relay nodes.
+  • Consensus-layer threats, which may cause established transactions to be reverted or the network to accept incorrectly formed blocks.
+Examples include majority (51%) attacks or selfish mining.
+Countermeasures include protocol-level defences, e.g., finality/checkpointing protocols, and robust incentive mechanism design.
+    Finality protocols are also helpful to CroCoDai implementations, because CroCoDai needs a logic for each supported chain to determine when a cross-chain transfer has been processed (i.e., highly unlikely to be reverted).
+Finality protocols provide such a logic by default.
+  • Contract-layer threats, which may allow attackers or malicious owners to subvert the intended functionality of a smart contract.
+Examples include smart contract bugs or deliberate backdoors.
+Countermeasures include static code analysis or formal verification, potentially run by third parties in the case of malicious owners or insider threats.
+CroCoDai’s similarity to the DSS suggests that we can leverage a larger market for smart contract validation services.
+  • Application-layer threats, e.g., bugs or backdoors in applications that support the stablecoin system, such as tools that allow users to interact with CDPs.
+As with contract-layer threats, countermeasures include static code analysis and formal verification.
+The main threats to stablecoin systems as enumerated by researchers at NIST (mell2023understanding, ) are categorized as security, stability, and trust “issues”.
+The following security “issues” are mentioned: (1) Unauthorized or Arbitrary Minting of Stablecoins, (2) Collateral Theft, (3) Malicious Smart Contract Update and Hijack, (4) Data Oracles, (5) Exploiting the Underlying Blockchain, and (6) Writing Secure Software and Vulnerabilities.
+Of these threats, (1), (2), and (5) are covered by the analysis in Section 4.4, as all scenarios effectively result in threats to the system’s overcollaterization.
+Finally, (4) is covered by the analysis in Section 4.5, whereas (3) and (6) are generic blockchain threats as discussed previously, which can be addressed using the countermeasures mentioned in (homoliak2020security, ; zhang2019security, ).
+Security threats to generic cross-chain systems are an active area of research (belchior2021survey, ; zamyatin2021sok, ), although some threats such as cross-chain replay attacks and data availability violations have been studied in the scientific literature.
+We discuss the fundamental properties of safety and liveness in CroCoDai in more detail below.
+4.2. Technical Safety and Liveness
+For local operations, the consensus protocol of the coin chain on which the local transfer occurs ensures the safety and liveness of operations on this chain – see, e.g., (garay2015bitcoin, ) – under the assumption of partial synchrony.
+This includes local stablecoin transfers, CDP operations, auctions, and interest withdrawals.
+As such, tokens cannot be stolen during local transfers or CDP operations or locked forever in a CDP unless the coin chain is compromised.
 For cross-chain operations, safety is violated if tokens in a bottom-level relay are burnt without a corresponding amount being minted on the receiving chain, and liveness is violated if tokens are held in escrow forever.
 However, the relay chain satisfies safety and liveness in our threat model, i.e., the transfer is eventually committed and coins are burned and minted on the outgoing and incoming chains, respectively.
-An alternative way to reach this conclusion is to view cross-chain transfers in CroCoDai as simple cross-chain deals [27] with the sender as the only party who votes, and the relay chain as a certified blockchain – depending on the design (as discussed in Section 4.4), either $12f+$ 1 nodes or the smart contracts on the relay chain sign the commit vote.
+An alternative way to reach this conclusion is to view cross-chain transfers in CroCoDai as simple cross-chain deals (herlihy2021cross, ) with the sender as the only party who votes, and the relay chain as a certified blockchain – depending on the design (as discussed in Section 3.3), either $12f+$ 1 nodes or the smart contracts on the relay chain sign the commit vote.
 To ensure safety and liveness, we emphasize that it is essential that the relay network is not compromised.
@@
 Although the security of the relay chain is therefore essential to CroCoDai, we investigate the impact of compromised coin chains on full backing below.
-Sudden Crash of Tokens.
+4.3. Sudden Crash of Tokens
 Occasionally, the price of a cryptocurrency token may suddenly collapse, e.g., because of the discovery of a serious bug in one of its smart contracts, or because of off-chain events such as a bankruptcy or a legal inquiry into the organization responsible for its maintenance.
@@
 We show that CroCoDai remains fully backed at time $t^{\prime\prime}$ if it is sufficiently over-collateralized and if the total debt created using tokens on the affected chains is limited.
-Formally, let $M^{\prime}\subset[M]$ denote the set of affected tokens, and $\zeta^{\prime}=\sum_{m\in M^{\prime}}\zeta_{m}$ the total debt ceiling of these tokens. Let
+Formally, let $\mathcal{J}_{t}$ denote the set of all CDPs at time $t$, $c_{j\,t}$ the total number of collateral tokens held in CDP $j\in\mathcal{J}_{t}$ at time $t$, and $D_{j\,t}$ the outstanding debt resulting from Dai withdrawals/deposits from/to CDP $j\in J_{t}$ at time $t$.
+We denote the token type associated with CDP $j$ by $y_{j}\in M$.
+Let $C_{t}^{*}=\sum_{j\in J}p_{y_{j}\,t}c_{j\,t}$ denote the total value of the system’s collateral at time $t$ and $D_{t}^{*}=\sum_{j\in J}D_{j\,t}$ the total debt.
+Let $M^{\prime}\subset M$ denote the set of affected tokens and $\zeta^{\prime}=\sum_{m\in M^{\prime}}\zeta_{m}$ the total debt ceiling of the affected tokens.
+Let
 $C^{-}_{t}=\sum_{\begin{subarray}{c}j\in\mathcal{J}_{t}\\ y_{j}\in M^{\prime}\end{subarray}}p_{y_{j}\,t}c_{j\,t},\;\;\;\;\;C^{+}_{t}=\sum_{\begin{subarray}{c}j\in\mathcal{J}_{t}\\ y_{j}\notin M^{\prime}\end{subarray}}p_{y_{j}\,t}c_{j\,t}$
@@
 We consider an attack that occurs at time $1$, so that $C^{-}_{1}=0$, $C^{+}_{1}=C^{+}_{0}$, $D^{-}_{1}=D^{-}_{0}$, and $D^{+}_{1}=D^{+}_{0}$.
-We can then prove the following theorem (see Appendix D):
+We can then prove the following theorem (see Appendix A):
 Theorem 4.1.
 […]
-Individual Compromised Coin Chain.
-If a majority of consensus power on a coin chain $b^{\prime}\in[B]$ is compromised then the consensus nodes could perform a 51% attack in the following way: send some amount of stablecoins to another chain, wait for the transaction to be committed by the relay network, then fork the compromised coin chain and undo the outgoing transfer request and the burning of the blockchain tokens.
+4.4. Individual Compromised Coin Chain
+If a majority of consensus power on a coin chain is compromised then the consensus nodes could perform a 51% attack in the following way: send some amount of stablecoins to another chain, wait for the transaction to be committed by the relay network, then fork the compromised coin chain and undo the outgoing transfer request and the burning of the blockchain tokens.
 Similarly, they could undo the funding of a CDP.
 As a consequence, stablecoins are created on another chain without backing collateral.
-In the worst case, this attack would be performed simultaneously on all of the token types supported by the compromised chain $b^{\prime}$.
-Let $M^{\prime}=\{m\in[M]:b_{m}=b^{\prime}\}$ be the set of affected tokens.
+In the worst case, this attack would be performed simultaneously on all of the token types supported by the compromised chain.
+Again, let $M^{\prime}\subset M$ denote the set of affected tokens.
 We again consider a worst-case attack that has been completed at $t=1$.
-Let $h^{\prime}$ be the total amount of stablecoins created in the attack, so that $D^{-}_{1}=D^{-}_{0}+h^{\prime}.$ We can then prove the following theorem (see Appendix D):
+Let $h^{\prime}$ be the total amount of stablecoins created in the attack, so that $D^{-}_{1}=D^{-}_{0}+h^{\prime}.$ We can then prove the following theorem (see Appendix A):
 Theorem 4.2.
 […]
-Compromised Price Feeds.
+4.5. Compromised Price Feeds
 Finally, we consider an attack in which a dishonest minority of price oracles tries to corrupt the price maintained by a vault.
 This could cause CDPs to be liquidated even if they are not underwater, thus violating safety of collateral.
-Let $p$ be the base price for token $m\in[M]$ at time $t\geq 0$ (we omit $m$ and $t$ from the notation in this section as they are not relevant to the analysis).
-Let $p_{o}$ be the price returned by oracle $o\in[O]$, so that $p_{o}=p+\Delta_{o}$ as discussed in Section 3.2.
-In the following, we assume some subset $O^{\prime}\subset[O]$ is controlled by malicious parties who aim to convince a vault in CroCoDai that of a price $p^{\prime}$ that is different from the base price $p$.
+We note that the DSS already has implemented safety mechanisms for corrupted price oracles, e.g., its Oracle Security Module (OSM).
+The OSM [^11] https://docs.makerdao.com/smart-contract-modules/oracle-module/oracle-security-module-osm-detailed-documentation increases the available time to detect corrupted price feeds by adding an artificial delay before price information is accepted by the system.
+Such safety mechanisms are complementary to our analysis in the sense that they make individual price oracles harder to compromise, whereas our analysis shows that attacks against the system as a whole have negligible probability of success if only a minority of oracles is compromised.
+Furthermore, CroCoDai’s inclusion of global price oracles on minor chains provides another complementary safety mechanism.
+After all, if price oracles from established global sources are included, then it is harder for attackers to compromise a majority among the full set of oracles than if this set only includes local sources from the minor chain’s less developed DeFi ecosystem.
+Let $p$ be the base price for token $m\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}$ at time $t\geq 0$ (we omit $m$ and $t$ from the notation in this section as they are not relevant to the analysis).
+Let $O$ denote the set of all oracles and $p_{o}$ the price returned by oracle $o\in O$, so that $p_{o}=p+\Delta_{o}$ where $\Delta_{o}$ is a normally distributed random variable as per Section 4.1.
+In the following, we assume some subset $O^{\prime}\subset O$ is controlled by malicious parties who aim to convince a CroCoDai vault to accept a price $p^{\prime}$ that is different from the base price $p$.
 Let $p^{*}$ be the actual median price among all price feeds, including those that have been distorted.
-We can then prove the following theorem (see Appendix D):
+We can then prove the following theorem (see Appendix A):
 Theorem 4.3.
 If $|O^{\prime}|<M/2$, then the probability that the difference between the base price $p$ and the median price $p^{*}$ exceeds a constant $c>0$ vanishes exponentially as $c$ increases.
-5.2 Financial Analysis: Approach
+4.6. Comparison to Baselines
+In a fully siloed baseline, each coin chain has its own DSS-like stablecoin.
+If the chain is not compromised, then local and cross-chain transfers are secure (i.e., tokens cannot be stolen or locked permanently), whereas they are not if the chain is compromised.
+In this sense, the baseline is similar to CroCoDai.
+The fully siloed system is unaffected by price crashes or platform failures on other blockchains – however, it is less resilient to price crashes and platform failures on its own chain than a system with more diversified collateral: e.g., in CroCoDai, a stablecoin holder on a minor chain that experiences a price collapse would still be able to transfer stablecoins to another chain and exchange them for other tokens.
+Finally, because CroCoDai supports additional price oracles, it is more difficult for an adversary to compromise a majority of oracles than in the baseline.
+The baseline that consists of the DSS with a wider set of wrapped tokens has the same security guarantees as CroCoDai, as it is similarly able to diversify its collateral.
+However, we will see in Section 6 that this baseline has lower efficiency.
+5. Financial Analysis
+The main system requirement from Section 3.4, full backing, can be violated not just due to critical attacks and platform failures, but also due to natural price fluctuations in token prices.
+This vulnerability to price fluctuations is compounded by the strong correlation between the prices of different cryptocurrencies (see Figure 6 in the appendix).
+In this section, we use a financial risk model to estimate the probability that the system violates full backing over a 1-day period.
+As is common in financial analysis, we will refer to a basket of collateral tokens as a portfolio, and to the relative increase in value of a token type over one time slot as a return.
+For our risk analysis, we compare a collateral portfolio that approximates the collateral composition of the existing DSS (without other stablecoins) to a portfolio made up of a wider range of blockchain tokens.
+In the following, we first present our model for token prices, and then present our risk analysis for the DSS-based portfolios.
+5.1. Token Price Model
+As mentioned in Section 4, we denote the base price of token $m\in M$ at time $t>0$ by $p_{m\,t}$.
+To model the evolution of token prices over time, we use Geometric Brownian Motion (GBM).
+This is a common choice in the financial literature on asset prices (ross2011elementary, ) and related work on stablecoins, e.g., by Salehi et al.
+(salehi2021red, ), particularly due to the GBM’s role in the Black-Scholes formula for option prices.
+A stochastic process $S_{t}$ that follows GBM is defined using a stochastic differential equation:
+(1) $dS_{t}=\mu S_{t}dt+\sigma S_{t}dW_{t},$
+where $\mu$ determines the drift of $S_{t}$ (i.e., its expected change per unit of time), $\sigma$ represents the volatility of $S_{t}$ (i.e., the variance of the change), and $W_{t}$ is the Wiener process that represents standard Brownian motion (ross2011elementary, ).
+In this model, each token price process $p_{m\,t}$, $t\geq 0$, can be represented using a process defined as in Equation 1.
+Essentially, this means that the logarithm of the relative difference of token prices, $\log\left(p_{m\,t+t^{\prime}}/p_{m\,t}\right)$, is normally distributed for all $t^{\prime}>0$ with mean and variance dependent on $t^{\prime}$.
+Unlike (salehi2021red, ), we consider a setting with multiple assets; i.e., we have a stochastic process $S_{m\,t}$ for each token $m$ with its own mean $\mu_{m}$ and variance $\sigma^{2}_{m}$.
+In this setting, we must also consider the covariance between the token prices (ross2011elementary, ), as cryptocurrency prices are often highly correlated – we will discuss this in more detail below.
+The standard GBM as decribed above is based on the normal distribution, i.e., $\log(S_{t+t^{\prime}}/S_{t})$ has a normal distribution.
+However, it is known from the academic literature that the probability distribution of stock price fluctuations is usually more heavy-tailed and that Student’s $t$-distribution is often a better fit (kelly2014tail, ; li2020review, ).
+This is also noticeable in the data for price changes in our dataset.
+As such, we will focus on the $t$-distribution in our analysis.
+Although the $t$-distribution does not allow for an analytical solution for the asset price change after each individual time slot, we can use Monte Carlo simulation to analyze its behavior, as we discuss in more detail below.
+5.2. Risk Analysis: Approach
 In this section, we investigate the probability that CroCoDai remains fully backed even in the presence of standard price fluctuations.
@@
 The second probability corresponds to the safety of the single-chain DSS, whereas the third option corresponds to the safety of CroCoDai.
-The purpose of this section is to demonstrate that the third option has considerably lower risk than the second, using multi-chain collateral baskets that are similar to those observed for the DSS in practice (see Table 8 in Appendix E).
+The purpose of this section is to demonstrate that the third option has considerably lower risk than the second, using multi-chain collateral baskets that are similar to those observed for the DSS in practice (see Table 5 in Appendix B).
 We discuss our approach in Section 5.2 and present our empirical results in Section 5.3.
-As is common in financial analysis, we will refer to a basket of collateral tokens as a portfolio, and to the relative increase in value of a token type over one time slot as a return.
-We will also discuss a method based on convex optimization that allows protocol engineers to set the system parameters such that the risk of system failure is minimized.
 Problem Statement.
@@
 In our model, failure of the stablecoin system occurs if, for any $t$ we have
-$\sum_{m\in[M]}\alpha_{m}p_{m\,t}\leq\theta D^{*},$ (3)
+(2) $\sum_{m\in M}\alpha_{m}p_{m\,t}\leq\theta D_{t}^{*},$
 […]
 In our simulation, we represent the portfolio as relative proportions of different assets.
-We define a vector $\vec{v}$ that denotes the value of individual assets in the portfolio as $\textbf{v}=(v_{1},v_{2},\ldots,v_{M})^{T},$ where $v_{i}$ is value of asset $i$.
-Similarly, we define a vector of random variables for log returns of individual assets $\textbf{X}=(X_{1},X_{2},\ldots,X_{M})^{\intercal},$ where $X_{m}$ is the random variable for log returns of asset $m$.
+We define a vector $\vec{v}$ that denotes the value of individual assets in the portfolio as $\textbf{v}=(v_{1},v_{2},\ldots,v_{|M|})^{T},$ where $v_{i}$ is value of asset $i$.
+Similarly, we define a vector of random variables for log returns of individual assets $\textbf{X}=(X_{1},X_{2},\ldots,X_{|M|})^{\intercal},$ where $X_{m}$ is the random variable for log returns of asset $m\in M$.
 We then simulate the price of each asset as an individual GBM with Student’s $t$-distribution for the return after each time slot, as discussed in Section 5.1.
@@
 A matrix of random samples $\mathbf{R}$ can then be generated as $\mathbf{R}=\mathbf{LT}$, where the matrix $\mathbf{T}$ contains a $t$-distributed random sample for each asset type and simulated time slot.
-In Appendix B we discuss alternative simulation methods based purely on historical data and the standard GBM.
-Choosing the System Parameters.
-In CroCoDai, the values $\alpha_{m}$, $m\in[M]$, which represent the fraction of each asset $m$ in the total portfolio, are not set by the protocol engineers – instead, they emerge naturally as users create CDPs and withdraw stablecoins using the currencies of their choosing.
-However, protocol engineers can influence the range of potential portfolios by adjusting the debt ceiling $\zeta_{m}$ for each token $m$.
-When optimizing the portfolio of cryptocurrencies for a stablecoin, the protocol engineers’ main goal should be to minimize the portfolio’s volatility.
-In principle, a positive mean drift in the returns is beneficial, as that would imply that the over-collateralization grows naturally over time.
-However, the mean return over our observation period was close to zero, and for risk analysis, with cross-chain commerce as our driving use case, we focus mainly on short-term price changes.
-In this context, the mean return is not a meaningful optimization criterion, so we focus on minimizing the volatility instead.
-Assuming a normal or $t$-distribution of returns, we can calculate the covariance matrix $\text{Cov}(\mathbf{X})$ as discussed previously.
-We can then use convex optimization to find a portfolio that minimizes total volatility.
-We can write this optimization problem as
-$\begin{array}[]{rcl}\displaystyle\min_{\textbf{v}}&\textbf{v}^{T}\text{Cov}(\mathbf{X})\textbf{v},&\\[5.69046pt] \text{s.t.}&\displaystyle\sum_{i=1}^{M}v_{i}=1,&\\[5.69046pt] &\displaystyle 0\leq v_{i}\leq\lambda_{i}&\forall i\in[M],\end{array}$ (4)
-where we consider $M$ different cryptocurrencies and $v_{i}$ denotes the fraction of the portfolio invested in asset $i\in[M]$.
-The bounds $0\leq\lambda_{i}\leq 1$ enforce maximum contributions for the different individual asset types, to avoid portfolios with a small number of low-volatility tokens (such portfolios would be vulnerable to black-swan events as discussed in Section 5.1).
-Given a solution $v_{1},\ldots,v_{M}$ to (4), we can enforce the optimal portfolio up to some margin by setting $\zeta_{m}=(1+\beta)v_{m}$ where $\beta>0$ is the flexibility margin: a higher margin encourages more users to create CDPs, but also allows for larger deviations from the optimal portfolio.
-5.3 Financial Analysis: Evaluation
+5.3. Risk Analysis: Evaluation
 In this section, we analyze the risk of system failure in both single-chain and multi-chain settings using the approach described above.
@@
 We have collected data using the CoinMarketCap API over six different periods that span a total of 1264 hours of data, separated into 5-minute slots.
-The duration of these periods can be found in Table 9 in Appendix F.
+The duration of these periods can be found in Table 6 in Appendix C.
 The final period includes the collapse of the FTX exchange on November 11, 2022.
+Our main motivation was to choose a diverse range of market conditions, including a period of turmoil, to ensure that the estimated system failure probabilities represent a worst-case scenario (i.e., in more typical market conditions, CroCoDai would be safer).
 In each time slot, we collected price information of 100 tokens, namely those that were the 100 largest in terms of market capitalization during the first time slot.
-A complete list of the included token symbols, and the observed correlation between them, can be found in Figure 6 in Appendix F.
+A complete list of the included token symbols, and the observed correlation between them, can be found in Figure 6 in Appendix C.
 DSS-Based Portfolios.
@@
 We use the following multi-asset portfolios:
-  1. 1.
-    D-All, which approximates the collateral portfolio of the DSS as of early 2023 (see Table 8 in Appendix E).
+  1. (1)
+    D-All, which approximates the collateral portfolio of the DSS as of early 2023 (see Table 5 in Appendix B).
 (We replace GUSD, which was not in our dataset, with another fiat-backed stablecoin (USDT), and Maker’s real-world assets with gold-backed PAXG.)
-  2. 2.
+  2. (2)
     D-ERC, which approximates the DSS portfolio after removing all fiat-backed assets.
-  3. 3.
+  3. (3)
     D-Mix1 and D-Mix2, which use the same tokens as D-ERC but which increase the proportion of the minor ERC-20 tokens.
-  4. 4.
+  4. (4)
     C-Mix1 and C-Mix2, which replace the ERC-20 tokens in D-Mix1 and D-Max2 with tokens from other popular blockchain platforms: ADA, DOT, TRX, and SOL.
-The exact compositions of these portfolios can be found in Table 10 in Appendix G.
+We use these portfolios to show how the system performs with a mix of different types of collateral: other (fiat-backed) stablecoins, major cryptocurrencies (e.g., ETH/WBTC), and minor ERC-20 tokens.
+Tokens of the first type typically have much lower volatility, but introduce a dependence on trusted third parties.
+We will discuss this trade-off in more detail later in this section.
+The exact compositions of the portfolios can be found in Table 7 in Appendix D.
 In our experiments, the expected over-collateralization to avert system failure is $\theta=1.1$.
@@
 We then compute the risk of failure as the fraction of runs that resulted in failure among the total number of runs.
-The results are displayed in Table 1Appendix B includes similar tables with different simulation methods and confidence interval widths.
-Table 1: Chance of failure for single asset portfolios and initial over-collateralization.
+The results are displayed in Table 1.
+Since the 95% confidence intervals widths are very small for the results in this table, we have omitted them for brevity.
+Table 1. Chance of failure for single-asset portfolios and initial over-collateralization $\gamma^{\prime}$.
 single-asset multi-asset $\gamma^{\prime}$     BTC     ETH     ADA     SOL     DOT     TRX     AVAX     ALGO     EOS     D-All     D-ERC     D-Mix1     D-Mix2     C-Mix1     C-Mix2 1.2 0.021 0.047 0.078 0.228 0.063 0.018 0.126 0.166 0.097 0.001 0.051 0.040 0.050 0.030 0.029 1.3 0.003 0.008 0.014 0.093 0.012 0.003 0.029 0.051 0.016 0.000 0.007 0.005 0.006 0.003 0.003 1.4 0.001 0.003 0.005 0.059 0.004 0.001 0.009 0.019 0.004 0.000 0.002 0.002 0.002 0.001 0.001 1.5 0.001 0.001 0.002 0.041 0.002 0.001 0.004 0.007 0.002 0.000 0.001 0.001 0.001 0.001 0.001
+Table 2.
+Chance of failure for single-asset portfolios and initial over-collateralization $\gamma^{\prime}$.
+In the table, ‘0.0 $\pm$ 0.0’ indicates that system failure events were observed during the 1,000,000 simulation runs, but that the estimated probability has been rounded down to 00, whereas ‘—’ indicates that no failure events were observed.
+single-asset C-Mix1, stablecoin contribution $\delta$ $\gamma^{\prime}$ USDT USDC BUSD TUSD DAI $\delta=0.5$ $\delta=0.25$ $\delta=0.1$ $\delta=0.05$ $\delta=0$ vol.
+4.85e-5 9.70e-5 4.06e-4 3.48e-4 2.84e-4 7.86e-4 1.17e-3 1.39e-3 1.47e-3 1.55e-3 1.2 0.00001 $\pm$2.2e-6 0.00009 $\pm$9.2e-6 0.00177 $\pm$4.2e-5 0.00103 $\pm$3.2e-5 0.00078 $\pm$2.8e-5 0.00314 $\pm$5.6e-5 0.01198 $\pm$1.1e-4 0.02153 $\pm$1.5e-4 0.02548 $\pm$1.6e-4 0.02976 $\pm$1.7e-4 1.3 0.0 $\pm$0.0 0.00002 $\pm$4.7e-6 0.00041 $\pm$2.0e-5 0.00027 $\pm$1.6e-5 0.00023 $\pm$1.5e-5 0.00047 $\pm$2.2e-5 0.00137 $\pm$3.7e-5 0.00243 $\pm$4.9e-5 0.00304 $\pm$5.5e-5 0.00345 $\pm$5.9e-5 1.4 0.00001 $\pm$3.0e-6 0.00021 $\pm$1.4e-5 0.00012 $\pm$1.6e-5 0.00009 $\pm$9.3e-6 0.00018 $\pm$1.3e-5 0.00052 $\pm$2.3e-5 0.00098 $\pm$3.1e-5 0.00105 $\pm$3.2e-5 0.00119 $\pm$3.5e-5 1.5 0.0 $\pm$0.0 0.00001 $\pm$2.8e-6 0.00012 $\pm$1.1e-5 0.00001 $\pm$8.3e-6 0.00001 $\pm$7.7e-6 0.00009 $\pm$9.6e-6 0.00029 $\pm$1.7e-5 0.00046 $\pm$2.1e-5 0.00033 $\pm$1.8e-5 0.00065 $\pm$2.5e-5
 From Table 1, we observe a considerable difference between the failure probabilities of the single-token portfolios.
-The most high-risk portfolio was Solana (SOL), which was severely affected by the collapse of FTX, since it was one of FTX’s most prominent backers [11].
+The most high-risk portfolio was Solana (SOL), which was severely affected by the collapse of FTX, since it was one of FTX’s most prominent backers (business_insider, ).
 Even for an initial collateralization $\gamma^{\prime}=1.5$, a pure SOL portfolio would have a failure probability of more than $4\%$ after a day, given the data from our observation period.
 In contrast, BTC and TRX were the most stable cryptocurrencies over this period.
-Figure 3: Correlation between the different token types.
+(a) Different blockchains
+(b) ERC20 tokens
+Figure 5. Correlation between the different token types.
 Furthermore, we observe from Table 1 that the DSS’s approximate portfolio has a very low probability of failure due to $80\%$ of its collateral being stablecoins and other pegged assets.
@@
 However, when comparing the C-Mix1 and C-Mix2 portfolios we do note that C-Mix2 has a lower failure probability despite the inclusion of less low-volatility WBTC.
-The reason is the inclusion of TRX, which is not only similarly low-risk as WBTC, but which also has a relatively low correlation with the other asset types – this can be seen in Figure 2(a).
+The reason is the inclusion of TRX, which is not only similarly low-risk as WBTC, but which also has a relatively low correlation with the other asset types – this can be seen in Figure 5a.
 We also note that the relative differences become starker when $\gamma^{\prime}$ increases – e.g., for $\gamma^{\prime}=1.2$, the risk of D-Mix1 is roughly a third higher than C-Mix2 (4.0% vs.
@@
 0.3%).
-Optimized Portfolios.
-The portfolios in the previous section approximated the existing DSS portfolio, which is aligned with the willingness of real-world users to deposit collateral of the various types.
-Although we did observe a slightly lower risk probability for the best multi-chain portfolio (C-Mix2) than for the best single-chain portfolio (D-Mix1), the underlying fundamental issue is how they compare if we were to use the convex optimization method based on (4) to determine low-volatility portfolios.
-A first question in understanding this is whether to include stablecoins in the portfolio.
-If we use convex optimization across the entire dataset of 100 token types, we obtain the portfolio from Table 12 in Appendix G.
-With the exception of BTT (the BitTorrent token), this portfolio only includes USD- or gold-pegged assets.
-A crypto-backed stablecoin system that only uses fiat-backed stablecoins as collateral has trivially low price volatility, yet would share most of the drawbacks of a fiat-backed stablecoin as discussed in Section 2 (i.e., higher vulnerability to black swan events that involve off-chain entities).
-As such, we will also consider a stablecoin that purely uses unpegged cryptocurrency assets as collateral.
-This represents a worst-case scenario in terms of volatility risk: in practice, an appropriate mixture of stablecoins and unpegged cryptocurrencies can be chosen.
-If we restrict the portfolios to the same tokens as the portfolios of Table 10, we obtain the optimal portfolios D-Opt and C-Opt of Table 11 (both in Appendix G).
-We also include the portfolio A-Opt which may contain any of the Top 100 tokens except those already pegged to a real-world asset.
-Since we observe that the optimal portfolios tend to restrict themselves to a handful of tokens (e.g., only WBTC and GNO/TRX for D-Opt and C-Opt), we also set a threshold of $\lambda=0.2$ for D-Opt and C-Opt and $0.1$ for A-Opt to avoid high exposure to events on a single blockchain.
-Table 2: Chance of failure for given portfolio and initial over-collateralization.
-D-Opt C-Opt A-Opt risk vs. D-Opt
-$\gamma^{\prime}$ $\lambda=0.2$ $\lambda=1$ $\lambda=0.2$ $\lambda=1$ $\lambda=0.1$ $\lambda=0.2$ $\lambda=1$ C-Opt A-Opt
-1.2 0.03419 0.01423 0.02103 0.00934 0.00895 0.00624 0.00617 -38.49% -81.75%
-1.3 0.00397 0.00133 0.00232 0.00131 0.00076 0.00055 0.00081 -41.56% -86.14%
-1.4 0.00115 0.00058 0.00087 0.00056 0.00030 0.00020 0.00024 -24.34% -82.61%
-1.5 0.00057 0.00027 0.00037 0.00024 0.00014 0.00008 0.00012 -35.08% -85.97%
-The results of our optimal portfolio simulations are displayed in Table 2.
-The final (bold) column displays the decrease in terms of failure risk estimates for C-Opt and A-Opt versus D-Opt for $\lambda=0.2$.
-In general, we find that the optimized portfolios have slightly lower failure probabilities than the DSS-based ones from the previous section (e.g., for $\gamma^{\prime}=1.1$ and $\lambda=0.2$, we obtain 0.034 for D-Opt vs.
-0.040 for D-Mix1, and 0.021 for C-Opt vs.
-0.029 for C-Mix2).
-The C-Opt portfolios again outperform D-Opt portfolios by roughly a third, confirming our expectation that a multi-chain stablecoin is less vulnerable to price fluctuations than a single-chain one.
-Furthermore, we find even lower failure probabilities for A-Opt, although this portfolio contains many tokens on blockchains that are not as high-use as the tokens in the other portfolios – e.g., BSV or KLAY vs.
-BTC and ETH in other portfolios.
-Less-used tokens may be more likely to experience black swan events as there are fewer incentives to scrutinize smart contract code or the identity of the consensus nodes.
-In any case, we find that even for pure crypto-backed portfolios (C-Opt) and sufficient over-collateralization ($\gamma^{\prime}=1.4$, which is similar to $\gamma=1.5$ in the DSS and $\theta=1.1$) CroCoDai’s estimated system failure (i.e., de-peg) risk after a day is less than one in a thousand, which may be acceptable for low-to-medium-value commerce (e.g., a ticket auction).
-6 Implementation
-In this section, we discuss the technical details of three prototype implementations of our stablecoin system.
-We compare them in terms of the on-chain processing cost (gas) and the off-chain computation time.
-Both are essential to system performance and thus usability: if the on-chain cost is too high then each transfer will incur large fees, whereas if the computation time is too high then there is a large delay for cross-chain transactions.
-We focus on the implementation of the relay nodes, because all other components depicted in Figure 2 (the coin, vault, auction, savings, and oracle contracts) can be re-used from the DSS without modification – we only add the relay contracts as wards for the coin contract.
-As such, we cannot provide a straightforward comparison with the DSS, as the performance of CroCoDai contracts is either exactly the same or there is no DSS equivalent.
-However, we will discuss the comparison with alternatives from the scientific literature, e.g., zkBridge.
-Each of our three designs uses a quorum of trusted nodes, but has a different tradeoff in terms of scalability and generality.
-Each implementation consists of two smart contracts written in Ethereum’s Solidity language: Relay.sol and CCTransfer.sol, although only Relay.sol differs between implementations.
-The Relay and CCTransfer contracts interact with an ERC-20 compliant contract for token transfers, Coin.sol, which re-uses the code from the Dai.sol contract in the DSS [41].
-Our experiments were run on a Lenovo IdeaPad with an Intel(R) Core(TM) i5-1035G1 CPU (1.00GHz) and 8GB RAM (although the gas costs are machine-independent).
-The scripts that performed off-chain computations were written in Python and used the Web3 package for most cryptographic operations, while the underlying blockchain was simulated using Truffle’s Ganache tool [21].
-For communication about cross-chain transfers, the relay nodes can use a private blockchain, e.g., a private Go-Ethereum blockchain, or a bespoke cross-chain communication framework such as [46].
-6.1 Smart Contracts
+Comparison to Existing Stablecoins.
+From Table 1, we observe that D-All has a considerably lower failure risk than the other portfolios.
+D-All represents the collateral portfolio of the DSS and is the only multi-asset portfolio from Table 1 that includes other stable assets (around 75% as per Table 5).
+All stablecoins except Dai in our dataset are fiat-backed stablecoins, which exhibit much smaller price volatility than regular tokens, but rely on trusted third parties which expose the system to external black swan events.
+We conduct experiments to further understand the price volatility.
+In particular, we estimate the failure risks for single-stablecoin portfolios and for a range of DSS-based portfolios with varying degrees of included stablecoins.
+For the latter, we consider modifications of the C-Mix1 portfolio, which represents a portfolio similar to Dai’s collateral portfolio but which includes a minor fraction (20%) of tokens from other blockchains.
+We modify this portfolio by assigning a fraction $\delta$ of fiat-backed stablecoins to it and multiplying the contribution of the other token types by $1-\delta$, for varying values of $\delta$.
+The low volatility of other stablecoins means that the failure probabilities of single-stablecoin portfolios are very low.
+Therefore, we conduct 1,000,000 runs for each table entry and display the 95% confidence interval half-width below each estimate in Table 2.
+We also display the variance, which is strongly related to the failure risks, of each portfolio in the table (under vol.).
+From Table 2, we observe that the failure risks of all single-stablecoin portfolios are much lower than those of the single-asset portfolios in Table 1, although there are still considerable differences between individual stablecoins: e.g., BUSD’s volaility is 10 times higher than USDT’s, and, for $\gamma^{\prime}$, BUSD’s estimated 1-day probability of failure is 177 times higher than that of USDT.
+For the modified C-Mix1 portfolios, we have used USDC as a representative stablecoin because of its high trade volume (there have been over $20$bn USDC in circulation since July 2021) and because its volatility is neither exceptionally high nor low among the included stablecoins.
+It can be seen that, as expected, the inclusion of other stablecoins leads to low failure risks, although we observe a small decrease for higher values of $\gamma^{\prime}$: in particular, the relative difference between $\delta=0$ and $\delta=0.25$ is roughly 2.5$\times$ (0.02976 vs.
+0.01198), but this decreases to 2.24$\times$ for $\gamma^{\prime}=1.5$.
+For $\delta=0$ and $\delta=0.5$, the relative difference decreases from 9.48$\times$ for $\gamma^{\prime}=1.1$ to 7.22$\times$ for $\gamma^{\prime}=1.5$.
+Overall, we observe that the inclusion of stablecoin fraction of $25\%$ can reduce the system’s failure risk by more than $60\%$.
+We note that a stablecoin system that is overcollateralized through a mixture of fiat-backed stablecoins is more resilient to black-swan events than the individual stablecoins.
+The reason is that the collapse of any single trusted third party would have a limited impact.
+However, black swan events may affect multiple fiat-backed stablecoins simultaneously, e.g., due to new legislation affecting stablecoin collateral.
+Thus, it would still be advisable to ensure that the total contribution of fiat-backed stablecoins is limited.
+We also note, that, in theory, a principal crypto-backed stablecoin could include other crypto-backed stablecoins as collateral.
+In this setting, the portfolio of the supporting stablecoins should be included in the portfolio of the principal stablecoin.
+Finally, protocol designers with substantial fiat assets (e.g., Dai) also have the option to steer collateral portfolios towards those with minimal risk (hajek2024collateral, ), which would reduce the risks associated with the crypto collateral even further.
+6. Implementation
+In this section, we present a prototype implementation of CroCoDai and compare its performance in terms of both cost and processing time to a baseline that combines the existing DSS with wrapped tokens.
+We present three different subvariants of the relay chain design with a quorum of nodes (i.e., the offchain variant of Section 3.3).
+Each subvariant has a different tradeoff in terms of scalability and generality, and consists of two smart contracts written in Ethereum’s Solidity language: Relay and CCTransfer, with only Relay different between the three subvariants.
+The Relay and CCTransfer contracts interact with an ERC-20 compliant contract for token transfers, Coin, which re-uses code from the Dai contract in the DSS (daisol, ).
+We discuss the smart contracts in more detail in Section 6.1.
+The other components depicted in Figure 3 (i.e., the vault, auction, and oracle contracts) are also re-used from the DSS.
+We perform two sets of experiments: microbenchmarks for the relay subvariants with varying quorum sizes, and end-to-end experiments for CroCoDai’s main operations: creating new tokens, and cross-chain transfers.
+6.1. Smart Contracts
 CCTransfer.
-This contract stores cross-chain transfer requests as a hashmap of Request instances, which are implemented as a Solidity struct.
+This contract stores cross-chain transfer requests as a hashmap of Request instances, which are implemented using a Solidity struct.
 Each CCTransfer contract is linked to a specific Coin and Relay contract pair, and must be made a Coin contract ward to allow it to mint or burn stablecoins upon the commitment of cross-chain transfers.
 Relay.
-In the first design, a hashmap of relay node addresses in Solidity’s address format is stored in the contract.
-For each cross-chain transfer, at least $12f+$ 1 relay nodes must call the Relay node’s approve function using their stored address.
+In the first design (named $n$/$n$), a hashmap of relay node addresses in Solidity’s address format is stored in the contract.
+For each cross-chain transfer, at least $12f+$ 1 relay nodes must call the Relay contract’s approve function using their stored address.
 This design exploits the EVM’s built-in signature verification, and generalizes trivially to other blockchain platforms.
 However, the downside is that each relay node must have an account and spend gas on each supported blockchain for every cross-chain transfer.
-In the second design, a hashmap of relay node public keys is stored in the contract, and a batch of $12f+$ 1 32-byte signatures is passed in the form of a byte array as a function parameter for each transaction.
+In the second design ($n$/$1$), a hashmap of relay node public keys is stored in the contract, and a batch of $12f+$ 1 32-byte signatures is passed in the form of a byte array as a function parameter for each transaction.
 These signatures are then validated one by one by repeatedly calling the built-in ecrecover function.
 Although this is not as generic as the previous approach (e.g., each chain with its different cryptographic primitives would require its own off-chain signature generation script) it is more efficient and does not require each individual relay node to maintain an account.
-In the third design, we use a $t$-out-of-$n$ threshold signature scheme (where in our case $t=2f+1$) to produce a single 32-byte signature for the relay node quorum, e.g., using the procedure of [23].
+In the third design ($1$/$1$), we use a $t$-out-of-$n$ threshold signature scheme (where in our case $t=2f+1$) to produce a single 32-byte signature for the relay node quorum, e.g., using the procedure of (gennaro2018fast, ).
 The resulting signature is a standard ECDSA signature that can be verified using ecrecover.
-For our implementation we have used the multi-party ECDSA library by ZenGo X [59].
-Table 3: Performance comparison of the three implementation designs.
+For our implementation we have used the multi-party ECDSA library by ZenGo X (zengox, ).
+Table 3. Performance comparison of the three implementation subvariants.
 gas cost $T_{\textnormal{proc}}$ (s)
@@
 (3) $1$/$1$ 53112 53117 53110 1.101 7.780 34.513
-6.2 Experimental Results
+6.2. Relay Microbenchmarks
 Table 3 compares the different designs in terms of the gas cost and the processing time at the relay nodes, $T_{\textnormal{proc}}$.
-We observe that the first design is the worst in terms of performance, which is the cost of its generality and ease of implementation.
-The second design is more efficient in terms of gas cost for large relay node quora, and only has little processing time at the relay nodes.
-The third design is the most efficient in terms of gas cost, but at the cost of considerable processing time at the relay nodes, especially for large quora.
-For the first ($n/n$) design, we could not distinguish $T_{\textnormal{proc}}$ from the transaction processing time in Ganache as both happen in a single step in Python’s Web3 package – however, these are assumed to be similar to signing in the $n/1$ design, and are indicated with ‘—’ in the table.
-By comparison, zkBridge [54] reports a gas cost of $\approx$230k gas for on-chain verification, but with a processing cost of $\approx$18 seconds to generate a proof with 32 high-end Amazon Web Services machines (c5.24xlarge).
-This lower performance is to be expected, as the additional complexity of zkBridge removes the need for a trusted quorum of relay nodes.
-7 Conclusions and Future Work
-In this work we have presented CroCoDai, a stablecoin system for cross-chain commerce.
-CroCoDai facilitates cross-chain transfers by design, and is more resilient than a single-chain system to black swan events and price fluctuations.
-Our empirical analysis shows that with sufficient overcollateralization, the probability of system failure can be made small even under a model whose price data includes the highly volatile period after the collapse of the FTX exchange.
-We have presented three prototype implementations that rely on a quorum of relay nodes for cross-chain transfers and governance decisions.
-For future work, a first interesting direction is to implement a version of CroCoDai that performs all relay operations on-chain, as per the second design of Section 4.4.
-Promising directions to achieve such a design would be to use trustless bridges as proposed in zkBridge, or succinct proofs of state such as those proposed for Algorand or Mina.
-This would remove the need for a trusted quorum of relay nodes, while maintaining CroCoDai’s advantages over the single-chain DSS, i.e., higher resilience to black swan events and price fluctuations.
-A second interesting direction is to further investigate alternative collateral liquidation mechanisms – for example, Salehi et al.
-[48] propose to mitigate liquidations from the DSS or remove them completely.
-For example, collateral of the same token type in a vault could be pooled together and liquidated only when the entire pool is underwater – however, this would remove an incentive from users to overcollateralize.
-A third and final interesting research direction is system recovery after a 51% attack on a coin chain.
-As we discussed in Section 5, such an attack could create stablecoins without matching collateral on a different chain.
-This is similar to the situation after a failed collateral auction in the DSS, so it may be possible to resolve it similarly – i.e., with the governance token holders absorbing the losses.
-Appendix A List of Symbols
-A list of symbols used in this paper can be found in Table 4.
-Table 4: List of symbols.
-symbol description
-$p_{m\,t}$ base price of token $m$ at time $t$
-$p_{m\,o\,t}$ price of token $m$ at time $t$, as reported by oracle $o$
-$\mu_{m}$ mean drift of the stochastic process $p_{m\,t}$
-$\sigma^{2}_{m}$ variance of the stochastic process $p_{m\,t}$
-$\mu_{m\,o}$ mean of the random variable $p_{m\,o\,t}-p_{m\,t}$
-$\sigma^{2}_{m\,o}$ variance of the random variable $p_{m\,o\,t}-p_{m\,t}$
-$y_{j}$ collateral token type of CDP $j$
-$c_{j\,t}$ number of collateral tokens held in CDP $j$ at time $t$
-$D_{j\,t}$ debt of CDP $j$ at time $t$
-$\zeta_{m}$ debt ceiling of token type $m$
-$\zeta^{\prime}$ total debt ceiling of the failed/compromised asset types
-$\gamma$ liquidation ratio
-$\lambda$ maximum contribution of each individual asset type during convex optimization
-$\theta$ system failure threshold
-$\mathcal{J}_{t}$ set of all CDPs at time $t$
-$M$ number of tokens
-$O$ number of price feeds
-$U$ number of users
-Appendix B Alternative Simulation Methods
-In the following, we discuss two alternative methods for estimating failure risks through simulation: using pure historical data (without an underlying price model), and with a standard GBM as discussed in Section 3.2.
-For the first method, we assume for each time slot that our portfolio had a given overcollateralization at the start of the slot and check if system failure occurs within a period of interest (e.g., one day).
-We repeat this process for all time steps in our dataset.
-The downside of this approach is that our dataset captures a period with many sharp price drops, which results in a somewhat pessimistic outlook on the failure risk – on the other hand, it captures realistically the limitations of stablecoins that rely on crypto-assets.
-For the second method, we generate correlated, normally distributed samples that have the same covariance matrix as the original prices.
-To achieve this, let $\mathbf{L}$ be computed as $\textnormal{Cov}(\mathbf{X})=\mathbf{L}\mathbf{L}^{T}$ using Cholesky decomposition.
-Random samples can then be generated as $\mathbf{LW}=\mathbf{R}$, where the matrix $\mathbf{W}$ contains random values sampled from the standard normal distribution.
-We can then calculate the returns of the assets in portfolio as
-$S_{t}=S_{0}\cdot\left\{\left(\mu-\frac{\sigma^{2}}{2}\right)+\sigma W_{t}\right\}.$ (5)
-We emphasize that the method used in Section 5.2 is the same as the second method, with the exception being that we use the multivariate $t$-distribution to represent the returns instead of the multivariate normal distribution.
+Both are essential to system performance: if the on-chain cost is too high then each transfer will incur large fees, whereas if the computation time is too high then cross-chain transactions incur a large delay.
+For the experiments, we used a Lenovo IdeaPad with an Intel(R) Core(TM) i5-1035G1 CPU (1.00GHz) and 8GB RAM – however, the gas costs are machine-independent.
+The scripts that performed off-chain computations use Python’s Web3 package for cryptographic operations, and the underlying blockchain was simulated using Truffle’s Ganache tool (ganache, ).
+To execute our experiments, we ran the Python scripts in different terminals that each interacted with an active Ganache instance on a single machine.
+From Table 3, we observe that the first subvariant is the worst in terms of performance, which is the cost of its generality and ease of implementation.
+The second subvariant is more efficient in terms of gas costs for large quora, and only has little processing time at the relay nodes.
+The third subvariant is the most efficient in terms of gas costs, but at the cost of considerable processing time at the relay nodes, especially for large quora.
+For the first subvariant (i.e., $n/n$), we were unable to distinguish $T_{\textnormal{proc}}$ from the transaction processing time in Ganache as both happen in a single step in Python’s Web3 package – however, these are assumed to be similar to signing in the $n/1$ design, and are indicated by ‘—’ in the table.
+6.3. End-to-End Experiments
+As discussed in Section 4.6, a baseline that uses the existing DSS with a greater variety of wrapped tokens has similar resilience to black swan events and price fluctuations as CroCoDai.
+However, we provide experimental results in this section that demonstrate that the DSS baseline has worse end-to-end performance than CroCoDai.
+We focus on the two main stablecoin actions: (1) creating new stablecoins on a coin chain and (2) transferring stablecoins from one coin chain (Chain A) to another (Chain B).
+To ensure that gas costs are comparable, we implement Chain A, Chain B, and the relay chain as private Ethereum chains using geth.
+For collateral, we implement a custom ERC20 token contract on the underlying chain (GemCoin) – for the DSS baseline, we also require a bridge for the custom token on Chain A (GemBridge), and another (DaiBridge) for Dai on the relay chain.
+For the collateral vault, we use instances of Vat, DaiJoin, and GemJoin where the latter is linked to the GemCoin contract – in CroCoDai the vault contracts exist on Chain A, whereas in the DSS baseline they exist on the relay chain.
+Finally, the baseline requires a WrGemCoin contract on the relay chain to implement the wrapped collateral token as an ERC20 token, and a WrDai contract for wrapped Dai tokens on each coin chain.
+Creating Stablecoins.
+For CroCoDai, the procedure for creating new stablecoins is as follows: the user first transfers the collateral to the vault and creates a CDP using GemCoin and GemJoin.
+Next, she uses the CDP to create stablecoins and transfers them to her account using Vat and DaiJoin.
+For the DSS baseline, the procedure is as follows.
+The user first sends GemCoin tokens to the bridge using GemCoin and BaseBridge.
+Next, a relay detects the transfer and creates an equivalent amount of wrapped GemCoin tokens for the user using WrGemCoin.
+The user then transfers the collateral to the vault, creates a CDP, creates stablecoins, and transfers them to her account using WrGemCoin, GemJoin, Vat and DaiJoin.
+Next, the user sends the Dai tokens to the bridge using Dai and DaiBridge.
+Finally, a relay detects the transfer and mints an equivalent amount of wrapped Dai tokens for the user using WrDai.
+Cross-Chain Transfers.
+For CroCoDai, the procedure for transferring stablecoins from Chain A to B is as follows.
+The user first transfers the tokens to CCTransfer and creates a transfer request.
+Next, a relay detects the request, sends a signed approval to the relay chain and Relay on Chain A, which calls Coin to burn the stablecoins.
+Finally, the relay sends the signed message to Relay on Chain B, which calls Coin to mint the stablecoins.
+For the DSS baseline, the procedure is as follows: the user first burns her wrapped Dai tokens using WrDai.
+Upon detection by a relay, the burned Dai is refunded to the user on the relay chain.
+The user then transfers the Dai to the DaiBridge for Chain B.
+Finally, a relay detects the transfer and mints an equivalent amount of wrapped Dai tokens for the user using WrDai on Chain B.
+Table 4. Computation costs (units of gas) of the transactions in CroCoDai and the DSS baseline, grouped by block. Bold: relay chain computations.
+CroCoDai DSS baseline
+creation transfer creation transfer
+Block 1 (Chain A) 466537 229325 244036 27762
+Block 2 (Relay) 133280 770724 294016
+Block 3 (Chain B) 185503 71498
+Block 3’ (Chain A) 71498
+Total (coin chains) 466537 414828 315534 99260
+Total (relay chain) 0 133280 770724 294016
+Results.
+Table 4 displays the cost of the processes discussed above, with transactions that can occur within the same block grouped together.
+For the DSS baseline, we assume that a single trusted custodian acts as a relay – to keep the comparison with CroCoDai as fair as possible, we hence use the first prototype subvariant ($n/n$) but with a single signer.
+For our implementation, we assume that the recipient data for outgoing cross-chain transfers (i.e., the ID of the receiving chain and the recipient’s account data) is not written to the blockchain; even in the fully on-chain design (see Section 3.3), this data can be exchanged off-chain and incorporated in the commit message or zero-knowledge proof.
 […]
-Table 5: Estimated portfolio failure probabilities - historical method.
-single-asset multi-asset $\gamma^{\prime}$     BTC     ETH     ADA     SOL     DOT     TRX     AVAX     ALGO     EOS     D-All     D-ERC     D-Mix1     D-Mix2     C-Mix1     C-Mix2 1.2 0.052 $\pm$0.0057 0.072 $\pm$0.0067 0.060 $\pm$0.0062 0.198 $\pm$0.0103 0.048 $\pm$0.0056 0.040 $\pm$0.0051 0.134 $\pm$0.0088 0.135 $\pm$0.0089 0.089 $\pm$0.0074 $0$$\pm$0.0000 0.067 $\pm$0.0065 0.060 $\pm$0.0062 0.060 $\pm$0.0062 0.058 $\pm$0.0061 0.052 $\pm$0.0058 1.3 0.002 $\pm$0.0012 0.024 $\pm$0.0040 0.003 $\pm$0.0013 0.054 $\pm$0.0059 0.006 $\pm$0.0020 0.002 $\pm$0.0012 0.020 $\pm$0.0036 0.037 $\pm$0.0049 0.022 $\pm$0.0038 $0$$\pm$0.0000 0.024 $\pm$0.0039 0.022 $\pm$0.0038 0.024 $\pm$0.0039 0.019 $\pm$0.0035 0.006 $\pm$0.0020 1.4 $0$$\pm$0.0000 0.004 $\pm$0.0016 $0$$\pm$0.0000 0.037 $\pm$0.0049 $0$$\pm$0.0000 $0$$\pm$0.0000 0.003 $\pm$0.0015 0.019 $\pm$0.0035 $0$$\pm$0.0000 $0$$\pm$0.0000 0.001 $\pm$0.0010 0.001 $\pm$0.0010 0.001 $\pm$0.0010 $0$$\pm$0.0000 $0$$\pm$0.0000 1.5 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 0.030 $\pm$0.0044 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 0.003 $\pm$0.0015 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000
-Table 6: Estimated portfolio failure probabilities - standard GBM.
-single-asset multi-asset $\gamma^{\prime}$     BTC     ETH     ADA     SOL     DOT     TRX     AVAX     ALGO     EOS     D-All     D-ERC     D-Mix1     D-Mix2     C-Mix1     C-Mix2 1.2 0.024 $\pm$0.0005 0.046 $\pm$0.0007 0.043 $\pm$0.0006 0.147 $\pm$0.0011 0.036 $\pm$0.0006 0.024 $\pm$0.0005 0.073 $\pm$0.0008 0.124 $\pm$0.0010 0.069 $\pm$0.0008 $0$$\pm$0.0000 0.045 $\pm$0.0007 0.037 $\pm$0.0006 0.042 $\pm$0.0006 0.034 $\pm$0.0006 0.031 $\pm$0.0005 1.3 0.001 $\pm$0.0001 0.005 $\pm$0.0002 0.003 $\pm$0.0002 0.070 $\pm$0.0008 0.002 $\pm$0.0001 0.001 $\pm$0.0001 0.010 $\pm$0.0003 0.025 $\pm$0.0005 0.007 $\pm$0.0003 $0$$\pm$0.0000 0.004 $\pm$0.0002 0.002 $\pm$0.0001 0.003 $\pm$0.0002 0.002 $\pm$0.0001 0.001 $\pm$0.0001 1.4 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 0.038 $\pm$0.0006 $0$$\pm$0.0000 $0$$\pm$0.0000 0.001 $\pm$0.0001 0.004 $\pm$0.0002 $0$$\pm$0.0001 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 1.5 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 0.019 $\pm$0.0004 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0001 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000
-Table 7: Estimated portfolio failure probabilities - Student’s $t$ GBM.
-single-asset multi-asset $\gamma^{\prime}$     BTC     ETH     ADA     SOL     DOT     TRX     AVAX     ALGO     EOS     D-All     D-ERC     D-Mix1     D-Mix2     C-Mix1     C-Mix2 1.2 0.022 $\pm$0.0005 0.047 $\pm$0.0007 0.077 $\pm$0.0008 0.227 $\pm$0.0013 0.062 $\pm$0.0008 0.019 $\pm$0.0004 0.126 $\pm$0.0010 0.164 $\pm$0.0012 0.099 $\pm$0.0009 0.001 $\pm$0.0001 0.050 $\pm$0.0007 0.040 $\pm$0.0006 0.048 $\pm$0.0007 0.028 $\pm$0.0005 0.029 $\pm$0.0005 1.3 0.003 $\pm$0.0002 0.008 $\pm$0.0003 0.014 $\pm$0.0004 0.092 $\pm$0.0009 0.012 $\pm$0.0003 0.003 $\pm$0.0002 0.029 $\pm$0.0005 0.051 $\pm$0.0007 0.015 $\pm$0.0004 $0$$\pm$0.0000 0.007 $\pm$0.0003 0.005 $\pm$0.0002 0.007 $\pm$0.0003 0.003 $\pm$0.0002 0.003 $\pm$0.0002 1.4 0.001 $\pm$0.0001 0.003 $\pm$0.0002 0.005 $\pm$0.0002 0.059 $\pm$0.0007 0.004 $\pm$0.0002 0.001 $\pm$0.0001 0.009 $\pm$0.0003 0.019 $\pm$0.0004 0.005 $\pm$0.0002 $0$$\pm$0.0000 0.002 $\pm$0.0001 0.001 $\pm$0.0001 0.002 $\pm$0.0001 0.001 $\pm$0.0001 0.001 $\pm$0.0001 1.5 0.001 $\pm$0.0001 0.002 $\pm$0.0001 0.002 $\pm$0.0001 0.042 $\pm$0.0006 0.002 $\pm$0.0001 0.001 $\pm$0.0001 0.004 $\pm$0.0002 0.008 $\pm$0.0003 0.002 $\pm$0.0002 $0$$\pm$0.0000 0.001 $\pm$0.0001 0.001 $\pm$0.0001 0.001 $\pm$0.0001 0.001 $\pm$0.0001 0.001 $\pm$0.0001
-We visualize the difference between the standard and $t$-distribution GBM for 10 separate simulation runs in Figure 4(a) and Figure 4(b), respectively.
-The $t$-distribution GBM runs are visibly more likely to experience sharp jumps than the standard GBM runs.
-Figure 4: For 10 separate asset price simulation runs, the total value of the portfolio 50% BTC and 50% ETH) over a 1 day period, using the standard GBM (a) and the GBM with the $t$-distribution (b).
-The dashed line represents the system failure threshold for $\gamma$ = 1.2 and $\theta$ = 1.1.
-Appendix C Distribution Fitting Examples
-The GBM is usually assumed to follow a normal distribution.
-In previous simulations we used normal distributions to simulate the returns.
-As mentioned by many papers, distribution of stock returns is usually more heavy-tailed and Student’s t-distribution is better fit (see, e.g., [34, 29]).
-When comparing residual sum of squares for normal and t-distribution on returns of ETH, we get a residual sum of squares (RSS) that is approximately equal to 31900 and 14000 for the normal and $t$- distribution, respectively.
-Moreover, the fitted $t$-distribution has a lower RSS for all other assets as well.
-In Figure 5, we have displayed a comparison of the normal and $t$-distribution fitted on Ethereum and Solona returns.
-As such, we prefer to use the $t$-distribution for the risk analysis in this paper.
-Figure 5: Comparison between t-distribution and normal distribution fitted on historical returns.
-Appendix D Proofs of Theorems 5.1 – 5.3
-Before we prove the main theorems, we need the following technical lemmas. We need the first to prove that if all CDPs are over-collateralized, then so is system as a whole.
-Lemma D.1.
+7. Conclusions and Future Work
+We have presented CroCoDai, a stablecoin for cross-chain commerce.
+CroCoDai facilitates cross-chain transfers by design, and is more resilient than a siloed baseline to black swan events and price fluctuations.
+We have presented a prototype implementation that relies on a quorum of relay nodes for cross-chain transfers and governance decisions, and which has better performance than a DSS-based baseline in terms of (expensive) transactions on the relay chain.
+For future work, an interesting direction is to implement a version of CroCoDai that performs all relay operations on-chain, as per the second variant of Section 3.3 – e.g., using the approach of zkBridge, or succinct proofs of state as in Algorand or Mina.
+Another interesting direction is to further investigate alternative collateral liquidation mechanisms – for example, Salehi et al.
+(salehi2021red, ) propose to mitigate liquidations from the DSS or remove them completely.
+A final interesting research direction is system recovery after a 51% attack on a coin chain.
+8. Acknowledgements
+This work was supported by Ministry of Education (MOE) Singapore’s Tier 2 Grant Award No. MOE-T2EP20120-0003.
+Appendix A Proofs of Security Theorems
+Before we prove the main theorems, we need the following technical lemmas. We need the first to prove that if all CDPs are overcollateralized, then so is system as a whole.
+Lemma A.1.
 If $a_{i}/b_{i}>c$ for $i=1,\ldots,n$, then $\frac{\sum_{i=1}^{n}a_{i}}{\sum_{i=1}^{n}b_{i}}>c$.
-Proof D.2.
+Proof.
 Note that $\frac{\sum_{i=1}^{n}a_{i}}{\sum_{i=1}^{n}b_{i}}=\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}\frac{b_{i}}{\sum_{j=1}^{n}b_{j}}$.
@@
 As such, it cannot be smaller than $\min_{i=1,\ldots,n}a_{i}/b_{i}$, which is greater than $c$ by the statement of the lemma.
+∎
 The second lemma is a commonly-known result, but we include the proof here for completeness.
-Lemma D.3.
+Lemma A.2.
 If $X$ has a normal distribution with mean 00 and variance $1$, then
 $R(x)=\mathbb{P}(X>x)\leq\frac{e^{-x^{2}/2}}{x\sqrt{2\pi}}.$
-Proof D.4.
+Proof.
 Since for $y\geq x\geq 0$ we have that $1\leq\frac{y}{x}$, we have that
@@
 With these lemmas, we are ready to prove the three theorems from Section 4.
-Proof D.5 (Proof of Theorem 5.1).
-If $p_{y_{j}\,t}c_{j\,t}/y_{j\,t}>\gamma$ holds for all $j\in\mathcal{J}$, then by Lemma D.1 it must hold that
-$\frac{C^{+}_{0}}{D^{+}_{0}}>\gamma.$ (6)
+Proof of Theorem 4.1.
+If $p_{y_{j}\,t}c_{j\,t}/y_{j\,t}>\gamma$ holds for all $j\in\mathcal{J}$, then by Lemma A.1 it must hold that
+(3) $\frac{C^{+}_{0}}{D^{+}_{0}}>\gamma.$
 Furthermore, we know from the definition of the debt ceiling that $D^{-}_{0}<\zeta^{\prime}(D^{-}_{0}+D^{+}_{0})$, where $\zeta^{\prime}$ (as defined in Section 4) is the total debt ceiling of the affected tokens.
 The above can be rewritten to
-$D^{-}_{0}<\zeta^{\prime}/(1-\zeta^{\prime})D^{+}_{0}.$ (7)
+(4) $D^{-}_{0}<\zeta^{\prime}/(1-\zeta^{\prime})D^{+}_{0}.$
 We therefore find that
 $\frac{C^{*}_{1}}{D^{*}_{1}}=\frac{C^{+}_{0}}{D^{-}_{0}+D^{+}_{0}}>\frac{C^{+}_{0}}{\left(1+\frac{\zeta^{\prime}}{1-\zeta^{\prime}}\right)D^{+}_{0}}>\gamma(1-\zeta^{\prime})$
-where the first inequality holds because of (7) and the second because of (LABEL:eq:unaffected_over-collateralized).
-Proof D.6 (Proof of Theorem 5.2).
+where the first inequality holds because of (4) and the second because of (3). ∎
+Proof of Theorem 4.2.
 Because of the debt ceilings on the tokens, the total debt incurred for each token type $m$ must still satisfy $\sum_{\begin{subarray}{c}j\in\mathcal{J}_{t}\\ y_{j}=m\end{subarray}}D_{j\,t}<\zeta_{m}\sum_{j\in\mathcal{J}_{t}}D_{j\,t}.$ – as such, even after the attack, we must have $D^{-}_{1}=D^{-}_{0}+h^{\prime}<\zeta^{\prime}(D^{+}_{0}+D^{-}_{0}+h^{\prime})$ or
-$D^{-}_{0}+h^{\prime}<\zeta^{\prime}/(1-\zeta^{\prime})D^{+}_{0}$ (8)
+(5) $D^{-}_{0}+h^{\prime}<\zeta^{\prime}/(1-\zeta^{\prime})D^{+}_{0}$
 where $D^{-}_{0}$, $D^{+}_{0}$, and $\zeta^{-}$ are as defined in the previous section. We have that
@@
 where the first inequality holds because of (5) and the second because $C^{-}_{0}>0$ and because of (3).
-As such, we again find that the debt ceilings ensure that CroCoDai remains fully backed if it is sufficiently over-collateralized.
-Proof D.7 (Proof of Theorem 5.3).
-In the following, we first assume that $O$ is odd and that some subset $O^{\prime}\subset[O]$ is controlled by malicious parties who aim to convince a vault in CroCoDai that the price is at most than some $p^{*}>p$.
+As such, we again find that the debt ceilings ensure that CroCoDai remains fully backed if it is sufficiently overcollateralized.
+∎
+Proof of Theorem 4.3.
+In the following, we first assume that $O$ is odd and that some subset $O^{\prime}\subset O$ is controlled by malicious parties who aim to convince a vault in CroCoDai that the price is at most than some $p^{*}>p$.
 In this case, the median equals the $\frac{1}{2}(M+1)+|O^{\prime}|$ th order statistic from a sample of $M-|O|$ normally distributed random variables with mean $p$ and standard deviation $\sigma_{o}^{2}$.
-This order statistic is greater than the maximum of the $M-O^{\prime}$ random variables, which we denote by $Y$.
-Let $\sigma_{*}=\min_{o\in[O]}\sigma_{o}$.
+This order statistic is greater than the maximum of the $|M|-|O^{\prime}|$ random variables, which we denote by $Y$.
+Let $\sigma_{*}=\min_{o\in O}\sigma_{o}$.
 $Y$’s cumulative distribution function equals
-$\begin{split}\mathbb{P}(Y\geq p^{*})&=\mathbb{P}\left(\bigwedge_{o\in[O]\backslash M^{\prime}}\Delta_{o}>p^{*}-p\right)=\prod_{o\notin O^{\prime}}R\left(\frac{p^{*}-p}{\sigma_{o}}\right)\\ &\leq R^{M-|O|}\left(\frac{p^{*}-p}{\sigma_{*}}\right)\leq\frac{e^{-(M-|O^{\prime}|)(p^{*}-p)^{2}/2}}{(p^{*}-p)\sigma\sqrt{2\pi}},\end{split}$
-where the first inequality holds because $R$ is a monotonically decreasing function, and the second inequality follows from Lemma D.3.
+$\begin{split}\mathbb{P}(Y\geq p^{*})&=\mathbb{P}\left(\bigwedge_{o\in O\backslash M^{\prime}}\Delta_{o}>p^{*}-p\right)=\prod_{o\notin O^{\prime}}R\left(\frac{p^{*}-p}{\sigma_{o}}\right)\\ &\leq R^{|M|-|O|}\left(\frac{p^{*}-p}{\sigma_{*}}\right)\leq\frac{e^{-(|M|-|O^{\prime}|)(p^{*}-p)^{2}/2}}{(p^{*}-p)\sigma\sqrt{2\pi}},\end{split}$
+where the first inequality holds because $R$ is a monotonically decreasing function, and the second inequality follows from Lemma A.2.
 As such, the probability $\mathbb{P}(Y\geq p^{*})$ vanishes exponentially as $p^{*}-p$ increases.
 The proof for $p^{*}<p$ is very similar and uses the symmetry of the normal distribution.
-Appendix E DSS Parametrization
+∎
+Appendix B DSS Parametrization
 The parameters of the various different vault types in the DSS can be found here.
 A sample of parameter values as of 25 April 2023 (Ethereum block 17121982) can be found in Table 5.
-For each of the supported token type $j$, we indicate the token name, total value of the collateral of CDPs of this type $\sum_{j\in\mathcal{J}}p_{j\,t}c_{j\,t}$, total debt of the collateral of CDPs of this type $\sum_{j\in\mathcal{J}}D_{j\,t}$, collateralization ratio, and debt ceiling $\zeta_{j}$.
+For each of the supported token type $j$, we indicate the token name, total value of the collateral of CDPs of this type $\sum_{j\in\mathcal{J}_{t}}p_{j\,t}c_{j\,t}$, total debt of the collateral of CDPs of this type $\sum_{j\in\mathcal{J}_{t}}D_{j\,t}$, collateralization ratio, and debt ceiling $\zeta_{j}$.
 Although the exact values vary throughout the day, this table is meant to illustrate the typical composition of the collateral type in the DSS.
@@
 The following liquidity pool tokens were deprecated: AAVE/ETH, LINK/ETH, UNI/ETH, wBTC/Dai, and wBTC/ETH.
-Table 8: Information about the various collateral types used by the DSS (as of Ethereum block 17121982).
+Table 5. Information about the various collateral types used by the DSS (as of Ethereum block 17121982).
 Name Token/Asset Platform debt coll. value overcoll. debt ceil. stab. fee liq. rate contrib.
-Standard ERC-20 Tokens
-ETH-A Ether 233580929 421276897 180% 15000000000 1.50% 145% 4.69%
+Standard ERC-20 Tokens ETH-A Ether 233580929 421276897 180% 15000000000 1.50% 145% 4.69%
 ETH-B 45480335 67797760 149% 250000000 3.00% 130% 0.91%
@@
 Total (ERC-20 tokens): 994893882 1407641598 20.00%
-LP Tokens
-UNIV2USDCETH-A USDC/ETH 2139902 0 0% 1.50% 120% 0.04%
+LP Tokens UNIV2USDCETH-A USDC/ETH 2139902 0 0% 1.50% 120% 0.04%
 UNIV2DAIUSDC-A Dai/USDC 23983390 0 0% 0.02% 102% 0.48%
@@
 Total (Liquidity Pool ERC-20 tokens): 246749074 66542992 4.96%
-Real-World Assets
-RWA001-A 6s Capital 14348036 15000000 105% 3.00% 0.29%
+Real-World Assets RWA001-A 6s Capital 14348036 15000000 105% 3.00% 0.29%
 RWA002-A Centrifuge 6429638 20000000 311% 3.50% 0.13%
@@
 Total (Real-World Assets): 713241736 1589000000 14.34%
-PSMs
-PSM-USDC-A USD Coin 2118503205 2608700392 123% 10000000000 42.58%
+PSMs PSM-USDC-A USD Coin 2118503205 2608700392 123% 10000000000 42.58%
 PSM-USDP-A Pax Dollar 499997224 500000000 100% 500000000 10.05%
@@
 Total (Peg Stability Modules): 3006039027 3546238990 60.42%
-Appendix F Price Data Table
+Appendix C Price Data Table
 Table 6 contains a specification of the time periods duirng which we collected price data for the empirical analysis of Section 5.
 Figure 6 visualizes the correlations between the 100 tokens in our dataset.
-Table 9: Data periods (times are in Greenwich Mean Time).
+Table 6. Data periods (times are in Greenwich Mean Time).
 start end number of slots
@@
 Period 6 04-11-2022 03:50 23-11-2022 05:15 5490
-Figure 6: The correlations between all 100 tokens in our dataset.
-Appendix G Portfolio Compositions
+Figure 6. The correlations between all 100 tokens in our dataset.
+Appendix D Portfolio Compositions
 Table 7 displays the composition of the DSS-based simulation portfolios.
@@
 Table 9 displays the optimal portfolio can be constructed using the 100 tokens in our dataset, if other stablecoins can be included.
-Table 10: DSS-based portfolio compositions.
+Table 7. DSS-based portfolio compositions.
 portfolio name token distribution
@@
 15% DOT, 15% TRX, 15% AVAX
-Table 11: Optimal portfolio compositions.
+Table 8. Optimal portfolio compositions.
 name optimal token distribution
@@
 TRX 10%, TWT 8.61%, WBTC 10%, XLM 6.951%
-Table 12: Optimal simulation portfolio if other stablecoins are included.
+Table 9. Optimal simulation portfolio if other stablecoins are included.
 BTT 0.34% BUSD 0.1504% PAXG 8.72%